A Complete Bibliography of Publications in Biometrika for the decade 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.04

Title word cross-reference

2 × 2 [Arm55, Lat53, Les55, Pat48, Pat59, Yat55a]. 2 × n [Hal55b]. 5 [Dav52e]. 99.9 [Lew53]. B [Tuk57]. b_{2} [GW47, GW50]. β [KL59]. β_{1} [PM51]. β_{2} [PM51]. χ [Tho53]. χ^{2} [AA54a, Bar51, Bar52b, Bat50, Ber57, Blo54, Blo56, Cad53, Dav50a, Daw54, Fra50, Hal55b, Hal55c, HP50b, Joh55c, Joh59b, Lan49b, Lan50b, Lan50c, Lew53, Nas59, Pea59b, Tuk57, Wat57b, Wis56b, Joh55c]. χ_{r}^{2} [VN59]. F [Aro50, DJ51, Dav52e, Hac58, Har50a, Hor53, PH51]. G [Kab58]. Γ [Joh55c]. k [AA54b, BO58, Coo51a, Coo51b, Jon54, Kap52, KR57, Wis52b]. n [Stu50]. n = 200 [HP57]. S [Sil50]. t [Ano54c, DS54, DS55, Gro53, Har57b, HP50a, Hyr50, Lor47, Lor50, Lor52, MP58, Moo57b, Rus50, Rus52, Sri58, Wel58]. T^{2} [PS59]. τ [Sun53b]. u [Lor50]. X [Bar56d]. z [Dav59i, Fin50].

-Approximation [Tho53]. -Distribution [Wel58, PH51, DS54, DS55, HP50a, Sil50]. -Distributions [Aro50].

-Function [Daw54, Joh55c, KL59]. -Funktsii [Joh55c]. -Integral [HP50b]. -Probability [Joh55c]. -Ratio [Hac58, Har50a, Dav52e]. -sample [Jon54]. -statistics [AA54b, Coo51a, Coo51b, Kap52, Wis52b]. -Test [Ano54c, Bar56d, Lor52, VN59, DJ51, Hor53, Dav59i, Fin50, Lor47, Lor50, Moo57b, Rus50, Rus52, Sri58]. -tests [Gro53].

1 [Con59]. 1950 [Stu52]. 1977 [Efr79].

Absenteeism [AS54, Nas59]. Absolute [Kam53a]. abstract [Tuk58]. Abundances [Gru51]. Academy [Ano57l]. Accident [Gre50, Joh57d, MPW53, Tan58]. Accidents [Bar53c, MPW52, Nas59]. According [Dar53]. Accuracy [Jam56]. Acheson [Cox53b]. Ackoff [Doi59]. Action [Ash58]. Adams [Dav56b]. Addendum [Ano50a, Bar53b]. Additional [Bra64, Bra54]. Additivity [Cox58a]. adequacy [Pat57]. Adjusted [HG58]. Admitting [Huz55]. Advanced [Pea59a, Dav54a]. against [Jon54]. Age [Wau55]. Age-Dependent [Wau55]. Agricultural [Fin58a]. Alan [Pea59a]. Algebra [Arc58, Mal59a]. Allen [Dav57c]. Allied [BDM58, Dav56d]. allo [Smi59a]. Allometry [KH50]. Allowing [Hea56]. Alternate [Bat50]. Alternative [Dav50a, Fin58b, Wel51]. Alternatives [Bar59a, Bar59b, BDM58, BD58a, Jon54]. Amalgamation [Mil59]. Amendments [Saw79]. American [Bow59]. among [Gur57]. Amongst [ZSP58]. Analogue [Som59, DS55]. analysi [Olk53]. Analysis [AS57, Ano56d, Ans58, AS54, Ash58, Bai53b, Bak50, Bar50b, Bar51, Bar58a, Ber57, Box55, BT52, Bra54, Bra64, Bul58, Cox53a, DJ51, Dav51b, DO51, Fos55a, Fuc52, HG58, Ham53, Har50c, Hea53, Jam54, Jen55, Jen58, Joh52b, Jow57, JW59, Law59, Les57, MPW53, Mal59d, Mar52b, McM55, Moo58b, Mos53, Nel51, PH51, Pil56, PB59, Pre53, Ray56, RM56, Rud52, Sch53, Sch69, Wat55, WH56, Wil55a, Wil57, Wil52b, Wil52c, Yat55b, Bra55, KR57, Rao59, Sri59, Joh51, Ano57c, Ans58, Dav52b, Dav59b, Dur59a, Edw59, Lew57, Lew58, Moo56b, Pow58, Wis52a, Jen59]. Analytic [Dav59a]. Anderson [Ans58, Dav54d, Mos59]. Andrew [Fos56, Vaj59b]. Angular [Har56, Har57a, Jen54a, Ste53]. Animal [Gru51, Mor51b]. Any [Gay50a, Gay51, Sil51]. Apodemus [HP52]. Apparent [Wal50]. Applicability [Ske58]. Application [BH54, Dav56e, Dur59c, GW50, HM59, Lan50c, Pea56, Som59, Stu51, Yat55a, GW47]. Applications [BO59, BO61, Cox58b, Dav51b, Dav55c, Kab58, Kam53a, LCC53, Mor51a, Tho55a, Coo51b, Dav55a, Dav58a, Dur55, Moo53a, Vaj59a]. Applied [Mal59d, Ray56, Pea52a]. Appraisal [Bar58a]. Approach [Wel51, Dav57c]. Approximate [AA54a, Bar53d, Bar53e, Bar55a, Bul57, Cox53c, Dan56, Dro58, KL59, Que49]. Approximating [Cad53, Law56a]. Approximation [Bar51, Bha59, Bro56b, Cad52a, McF56, MP58, Pea59b, Tho53]. Approximations [DS55, Joh52a, Pea52b, SK57, Tuk55, Tuk57]. Area [Ano55m]. Area-Ordinate [Ano55m]. Argument [Ans57, PM51]. Arising [Irw54, IK54, Lan49a, Lan50a]. Arithmetic [Pla58]. Arnoff [Doi59]. Arranged [Dar57a]. Arrangements [Cox54a]. Art [Stu56b]. Ascertaining [Gib50]. Aspects [Ans58, Ban51, Mal59d]. Assay [AB54, Hea53, Hea53]. Assigning [Hui55]. Associated [MT57, Mac58, Cox53a]. Association [Stu53, Wil52c, Wis54b]. Assumptions [Hoo58]. Astronomy [Dav57b]. Asymptotic [Der56, Fin58b, VN59]. Asymptotische [Dav52d]. Aurelia [Les57]. Author [Ano55a]. auto [Mur51]. auto-regressive [Mur51]. Autocorrelated [Hut56, Wal50]. Autocorrelation [Ken54a, Wis55]. Autocorrelations [MP54]. Autoregression [Pat58]. Autoregressive [Ghu50, Jen54b, Jen56, Wis56a]. Average [Dur59b, Wis56a]. Axis [LD53].

B [Bar56d, Bar57g, Dav58d, Joh51, Joh58f, Kal57, Ken55]. Back [Ano50b, Ano50c, Ano51a, Ano51b, Ano52a, Ano52b, Ano53a, Ano53b, Ano54a, Ano54b, Ano55b, Ano55c, Ano55d, Ano56a, Ano56b, Ano57a, Ano57b, Ano58a, Ano58b, Ano59a, Ano59b]. Bacteria [Pow55]. Bacterial [Mac53]. Bailey [Fos55c, Moo58c, Whi55]. Balanced [Ban50, Pat52, Sil57, VN59]. Balking [Hai57]. Bancroft [Smi55b, Dav54d]. Bartlett [Whi51, Dav55a]. Based [BCT56, Bro56a, Cha55, Dav53b, DO51, Hab55, Kam54a, Lon51b, Pea50a, Pea51c, SK57, Sun53a, Lor50, Moo57b]. Basic [Dav56b]. Basis [Pea56]. Bayes [BB58]. be [Ban50, Mau58b]. Beginnings [Ken56, Pen59]. Behavioral [Dav57a]. Behaviour [Dar56]. Behrens [Bar50a]. Bell [Joh54b]. Bennett [Box55]. Berkeley [Dav50b, Dav52c, Dav57b]. Bers [Mal59d]. Bessel [Bar55b]. Best [Nag50, ZSP58, Rao53]. Beta [Aro50, Cad52a, FR57, Fos57, Fos58, HF51, Wis50]. Beta-Function [HF51]. Between [Bar53c, Ber57, Bos56, Dar53, Dar57a, Gay50b, GR54, Gro51, Han55a, Has57, Her58a, Her58b, Joh59b, MPW52, Pat48, Pat59, Sil51, Tan53, Tat55, Wal50, Wis54c, Har57b, Lie54, Pal57, Sun54]. Beverton [Dav59g]. Bi [Coo51a]. Bi-variate [Coo51a]. Bias [Bar55a, Dur59c, Ken54a, MP54, Que56, Tuk58]. Bibliography [Ano58e, Doi57, Men58]. Bickley [Bar55b]. Binary [Cox58b]. Binomial [Ans50, AP54, Ans56, Arm58, Bai56c, Bla56, Blo54, Blo56, Bra58, CR58, HF51, Joh57a, Joh58d, Lin57a, Moo55c, Ram59, Sam55, Ste57, Stu57, BO58, Dar59b]. Binomials [Bai53b]. Bio [Hea53]. Bio-Assay [Hea53]. Bioassay [AC55]. Biological [Bar57e, Cra53a, Hea53, Les58]. Biology [Dav57b, Dav59h, Smi55b]. Biomathematics [Ken55]. Biometric [Dav54a]. Biometrika [Ano51c, Ano55i]. Biometry [Smi55a]. Birth [Dar53, Dar56, Wau55]. Bits [Bar52c]. Bivariate [Bro57, Cad51, Hyr52, Lan57, Coo51b, DS54]. Bizley [Bar58b]. Blackwell [Mor56]. Bliss [Hea53, Smi55a, Dav59h]. Block [Ano56d, Ban50, BT52, Bra54, Bra64, GF57, Sil57, VN59, Wil55a, Bra55]. Blocks [Mil59, Pea57a]. Book [Ano57e, Ano57d, Ano57f, Ano57c, Ans58, Arc58, Arm54, BS57, Bai58, Ban58, Bar53a, Bar56a, Bar57b, Bar55b, Bar56c, Bar56d, Bar57f, Bar57g, Bar58a, Bar58d, Bar58b, Bar58c, Bar59f, Bar59c, Bar59e, Bar59d, Bea50, Bea51, Boo57, Bow59, Box55, Bro56c, Bur59, Car59, Cor58, Cox53b, Cox53a, Dav50c, Dav50b, Dav51a, Dav52a, Dav52b, Dav52d, Dav52c, Dav54d, Dav54a, Dav54e, Dav54b, Dav54c, Dav55c, Dav55a, Dav55b, Dav56a, Dav56b, Dav56c, Dav57d, Dav57c, Dav57b, Dav57a, Dav58a, Dav58c, Dav58d, Dav58b, Dav59e, Dav59b, Dav59c, Dav59a, Dav59f, Dav59g, Dav59h, Dav59d, Doi59, Dur53, Dur55, Dur59a, Edw59, Ela59, Est59, Fin58a, Fos55a, Fos55b, Fos56, Gum58, Hea53, Jen55, Jen58, Jen59, Joh51, Joh54b, Joh54a, Joh54c, Joh55a, Joh55b, Joh55c, Joh56, Joh58a, Joh58b, Joh58c]. Book [Joh59a, Jon56, Kal57, Ken55, Ken59, Kes57, Kes58, Lew57, Lew58, Mal59b, Mal59c, Mal59a, Mal59d, McM55, McM57, Moo53a, Moo55b, Moo55c, Moo55a, Moo56a, Moo56b, Moo57a, Moo58b, Moo58a, Moo58c, Mor51a, Mor56, Mor57, Mos53, Mos59, Owe58, Owe59, Pea56, Pea51a, Pea51b, Pea52a, Pea59a, Pen59, Pow58, Smi51, Smi55a, Smi55b, Smi56, Smi59a, Stu52, Stu54, Stu56b, Stu56a, Stu58, Vaj59b, Vaj59a, Wat57a, Wel54, Wel58, Whi51, Wil54, Wil58a, Wis52a, Wis54b, Wis59b, Wis59a]. Books [Ano55n, Ano56h, Ano56i, Ano57i, Ano57j, Ano58f, Ano58g, Ano59f, Ano59g]. Boole [Dav59c]. Booth [Fos56]. Bootstrap [Efr79]. Both [Mal56]. Bound [RG58]. Bounds [Chu55b, RG57, RG61, Stu56c, Tho55b]. Brambilla [Dav58c]. Brand [Pow58]. Bross [Joh54a]. Buckland [Gum58]. Bureau [Ano55o, Ano56j, Ano57k, Har53, Pea54]. Burington [Dav59d]. Buros [Stu52]. Bush [Jon56]. Business [Stu56a, Vaj59b, Dav56b]. Busy [Con59].

C [Arc58, Ban58, Bar56a, Bar59c, Bar59e, Box55, Bro56c, Dav54a, Dav56b, Dav56c, Dav59e, Dav59a, Dav59f, Dav59h, Dav59d, Doi59, Hea53, Mal59d, Moo58b, Pea52a, Smi55a, Wel54, Wil58a]. Calculating [TW54, Les55]. Calculation [GW47, GW50, Hal55b, Joh55c]. Calculus [BS57, Ken56, RK50, Dav59c]. Calhoun [Smi55a]. Cambridge [Wel54]. can [Mau58b]. Canonical [Law59, Mar52b]. Capture [Gil56, Ham53, LC51, Les52, LCC53, Mor52]. Capture-Mark-Recapture [Mor52]. Capture-Recapture [Gil56, Ham53, LC51, Les52]. Cardpack [Dav52a, Dav56a]. Cards [Ken57b]. Carlo [Mor57]. Carolina [DO51]. Carried [Les57]. Case [AS57, Cle59, Sun53b]. cases [DS54]. categories [KR57]. Caudatum [Les57]. Caused [Cha55]. Cedric [Ken55]. Censored [Coh55, Coh57, DJ54, Gru52, Gup52, MH58, Moo56c, Saw58a, Saw58b, Saw59, Saw79]. Census [Dar59a, Dar58]. Censuses [Smi51, Stu54]. Central [AA54a, MP58, PH51, Pea59b, San59, Tuk57, Wel58]. Certain [Ano53c, Ano55g, Ano55m, Bar52a, Cox54a, Dav54f, DO51, Dra52, Gan56, Kha59, Lan49b, Lan50b, Les58, Mar52a, Nai50, Ray56, Sil51, DS54, DS55, Olk53]. Chain [Ano56c, Bai53b, Bai56c, Dar59c, Gan55b]. Chain-Binomial [Bai56c]. Chain-Binomials [Bai53b]. Chains [Bar52b, Der56, Gan56, Goo55, Goo57, Goo58, Hoe54]. Chance [Bai53b, Bai56c, Dav56a, Dav52a]. Chances [BB58]. Change [Cha55, Pag55b, Pag57]. Changes [Stu56a]. Chapters [BS57]. Characteristic [Her58a, Lei58a]. Charles [BS57]. Charlier [BD52, She51].

Chart [HF51]. Charts [Pag55a, PH51]. Chayes [Bar58a]. Chemical [Box55, McM57]. Chemistry [Box55]. Chessboard [Dav52a, Dav56a]. Chi [San59]. Chi-Square [San59]. Ching [Smi56]. Choice [Dav52a, ZSP58, Dav56a]. Chun [Smi56]. Churchman [Doi59]. Circle [Dan52, WW56]. Circulant [Goo50]. Circular [Dan52, Wat56b, Wat56a]. Class [Bar56e, Rao53, RS56, Sil57, Sim55]. Classical [BS57]. Classification [RB58, Stu55b]. Classifications [Joh58f]. Clew [Joh59a]. Climatology [Dav59h]. closed [Dar58]. Cochran [Joh51, Joh58a, Dav54c]. Coefficient [Ano55h, BC58, Ber58, Cox54b, DK58a, Gay51, Har54, Har56, Har57a, Hil54, Jen54a, LD53, San58, Stu51, Sun53b, Har57b]. Coefficients [CM51, Dan56, FHP57, Hoo58, Que57, Rei56, Wat56b, Wis52b, Moo55c]. Collecting [Cox57a]. Colonies [Mac53]. Combination [Lan49a, Lan50a, Pea50a, Pea51c, Tan58, Yat55a]. Combinations [Hui55, MM59]. Combinatorial [Dav50d, Dav59b]. Come [Dav50d]. Comment [Bar57c]. Commerce [Ano55o, Ano56j, Ano57k, Har53, Pea54, Dav58b]. Common [BDS54, BO58]. Community [RM55]. compare [Fin50]. Comparing [Bar56d, Bir53, Suk58, Dar59b]. Comparison [Bos56, Coc50, Dar59c, Jam51, Joh50, Joh52b, Pea52b, Stu53, Tay50, TW54, Wel51, Wil57, Ber57]. Comparisons [BT52, Bra54, Bra64, Dav59j, HG58, Mau58b, vdH52, Bra55]. Competing [LG58, LG59]. Competitive [Bar57e]. Compilation [Joh55b]. Complete [Mil59]. Complex [Bar56c, Est59, Woo56]. Components [Bul57, Nel54, Pre53]. Composite [Rao53]. Composition [Cha55]. Compound [Gur57, Hyr52, Hyr50]. Computation [GW47, GW50, Lor54b]. Computer [Moo58a]. Concentration [AC55]. Concept [Lin57a]. Concepts [Dav56b]. Concerning [Cha50, Eva53]. Concise [Bar59c]. Concomitant [Cox57a, Cox57b]. Concordance [Bos56, Stu51]. Concurrence [Toc52b]. Concurrent [Wil53]. Conditioned [Wau58, Wau59]. Conditions [Ano56c, Ash58, Gan55b, Gan56, Wis56a, BD52]. Confidence [Bar53d, Bar53e, Bar55a, BH54, Bul57, Bul58, Cle59, Cro56, Cro58a, CG59, Hui55, Joh57c, RG57, RG58, RG61, Ste54, Tuk58]. Confusum [CS57]. Connexion [Joh59b]. Consecutive [Dar57a]. Considerations [Joh58f]. Considered [Lan57]. Consistency [Cha54, Pat57]. Constants [HP51]. Construction [Pat52, WW56]. Contagious [Bat50, Eva53, Tho54]. Container [CS57]. Contingencies [Joh55a, Joh58b]. Contingency [Daw54, FH51, Lan57, Nas59, Som59, Ste51a, Stu53, Wil52c, Arm55, Lat53, Les55]. Continued [She54]. Continuous [Bar50b, BM55, HM59, Mor50a, Pag54, Sil51, Tat55]. Contour [Wis50]. Contrasts [Sch53, Sch69]. Contribuito [Smi59a]. Contributions [RG57, RG58, RG61, Pea51a]. Control [Cox53b, FH54, Lan52, Lon51b, Moo58d, Moo59, Pag55a, Moo57a]. Controls [Hea56]. Convergent [Ano55e, Wis54a]. Converging [Wis50]. Cook [Joh55a, Joh58b]. Coombs [Bar56a]. Cornell [Dav57d]. Corporation [Moo55a]. Corrected [May52]. Correction [Bar55a, Har50b, vdH52]. Corrections [DW51a, Kup52, Saw79, Sch69]. Correlated [But56, Lei58b, Wil52d]. Correlation [Ano57e, Ber58, Cha50, CM51, Dan56, DK58a, Dav50c, DKS51, DW50, DW51a, DW51b, Dur57, FHP57, Gay51, Gib50, Han55a, Han55b, Han57, Har54, Har56, Har57a, Hoo58, Jen54a, KH50, Law56b, LD53, Lei58a, Moh55, Mor51c, Que57, Rei56, San58, Som59, Stu56c, Sun53b, Tat55, Wal50, Wat55, Wat56b, WH56, Har57b, Que49]. Correlational [Dur59a]. Corrigenda [Ait59, Ano53c, Ano53d, Ano54d, Ano54c, Ano55g, Ano55h, Ano55e, Ano55f, Ano56d, Ano56c, Ano60, BD57a, Blo56, BO61, Cox57a, Cro58a, GW50, Goo57, Joh58d, Kam58, Lan50a, Lan50b, LG59, Lor52, May58, Moo59, Pat59, Pea51c, RG61, Rub65, Saw58a, She59a, She61, Smi59b, Wau59]. Corrigendum [Bra64]. Costs [Moo58b]. Counting [Lan52]. Counts [Mac53]. Course [Ans58, Pea51b, Pen59]. Covariance [HG58, Law56b]. Coverage [GT56]. Covering [Dan52]. Cowden [Moo57a]. Cox [Joh51, Joh58a, Mal59c]. Criteria [Coo57, Law56a]. Criterion [BCT56]. Critical [Hui58, Rao53, TWJ56]. Cronbach [Ban58]. Crops [Owe59]. Cross [Has57, Joh58f]. Cross-Classifications [Joh58f]. Cross-Infection [Has57]. Crossing [Tan51]. Cubes [MT57, Mac58]. Cubic [Bar59d]. Cumulants [Jam55, Kup52, Smi59b, Smi59c, Coo51a, Kap52, Sil50, Stu50]. Cumulative [Ano55e, HP50b, HF51, Wis54a]. Curtiss [Lew58]. Curve [Pat58, PL59]. Curves [Hut56, PM51, BD52]. Cut [Har50a].

D [Arc58, Arm54, Dav59d, Fos56, Hea53, Joh54b, Joh54a, Mal59c, Moo57a, Moo58a, Mor56, Pea56, Smi55a, Stu58, Wel54, Whi51]. D. [Bar57c]. Dams [Gan55a]. Data [AS54, Ash58, Bai51, Bro56a, DJ54, Gue53, HM54, Har51, Hut56, Lan49a, Lan50a, LC51, Les52, LCC53, Les57, MPW53, MH58, Roy56, Dav58d]. Dates [Ano55i]. David [Wil54]. Davies [McM55, McM57]. Davis [Bar56a]. Death [Dar59a, Dar53, Dar56, LC51, Mor52, Wau55]. Death-Rate [LC51]. Death-Rates [Mor52]. Decision [Bar53a, Bar54a, BDS54, Dar59b, Joh54a, Bar56a]. Decisions [Ban58, Mor56]. definite [BD52]. Degrees [Tho53, Wis56b]. Delay [Tan51]. Demand [Fos55a]. Demografia [Car59]. Demography [Joh56]. Density [Ker58, Wis55, Pat57]. Denumerable [Der56]. Department [Ano55o, Ano56j, Ano57k, Har53, Pea54]. Departure [Bar57a]. Dependence [Ans57]. Dependent [Gab59, Wau55]. Dept [Dav58b]. Derivation [Lan49b, Lan50b, Ske58]. Derived [GW47, GW50, PH51, Ste51b]. Deriving [Rao53]. Description [Ber57]. Design [AC55, Ano56d, BH54, BL59, Cox53a, Cox54a, Cox57a, Cox57b, GF57, Joh51, Joh54c, MM59, Mos53, Pea56, Toc52a, VN59, Wil55a, Joh54a, McM55]. Designs [Ban50, Ban51, Big59, Bos56, Box52, BT52, Bra54, Bra64, But56, Cox51, Dav58b, Joh59a, Pat52, Ray56, Sil57, Wil57, Wil52d, Bra55, Joh51, Joh58a].

Determination [Les59]. Determined [Joh57c]. Deterministic [RM55]. Deviate [Dav56f, Nai52, PT59, Pil59]. Deviates [Moo55a]. Deviation [Cad54, Cro58b, DHP54, Gup52, Joh57a, Joh58d, Joh58e, Kam54a, Kam54b, Lor52, Ram58, Tho55b, Lor47]. Deviations [Cha50, LD53]. Dichotomized [Tat55]. Dicing [Dav55d]. Dictionary [Her58b, Gum58]. Difference [Dar57a, Gay50b, GR54, Gro51, Kam53b, Kam53c, Kam58, Pat48, Pat59, Que53, Ram58]. Differences [Dar53, Kam54a, Ram59, SK57, Wil58a, Mur51, Dav59c]. Differential [Mal59d, Phi59, Ste51b]. Digital [Moo58a]. Digits [Moo55a]. Dimensional [Hei55]. Dimensions [Har57c, Dav59a]. Direct [Ano55e, Wis54a]. Discontinuous [DJ50, Ker58, Pea50a, Pea51c, Ram58, Ste50, Toc50]. Discovery [Ken59]. Discrete [Lan49a, Lan49b, Lan50a, Lan50b, Mar52a]. Discriminant [Wil55b]. Discrimination [Rud52]. Discussion [TW54]. Diseases [Fre53, Irw54]. Disorientation [MT57, Mac58]. Dispersal [Ske51]. Dispersion [Bat50, Cad53, CS55, Dar57b, Kat53, Nai50, SK57]. Distance [Bul58]. Distributed [Gru51, MH58, Dav59i]. Distribution [AA54a, Ano53c, Ano55g, BP59, Bar57a, Bat50, Ber57, Bha59, Bra58, Cad52b, Cle59, CR58, CS57, Cro58b, Dan52, Dan56, DKS51, Dav54f, DHP54, Der56, Dra52, Fin52, FR57, Fos57, Fos58, Gab59, Gay50a, Gay51, Hac58, HP52, HS56, HP57, HP50b, HF51, HP51, Her58b, Hil54, Hyr50, Irw54, Joh52a, Joh57a, Joh58d, Kab58, Kam53a, Kam54a, Kam56, Ken57a, Lan57, Law56a, Lei58a, Lew53, MP58, Moo52, Moo56c, Pal57, PH51, Pea52b, Pea59a, Pea59b, Pil52, Pil56, PT59, PB59, Sam55, San59, She59b, Sim55, Ske58, Ste50, Ste51b, Stu51, Suk58, Tay56, Tuk57, Wal56, Wat56b, Wat56a, Win57, Wis54c, Woo56, Coo51a, DS54, DS55, HP50a, Jon54, She51, Sil50, Wel58]. Distribution-Free [Kam56, Jon54]. Distributions [Ano53c, Ans50, Aro50, Bar56e, Blo54, Blo56, Cad53, Chu55a, Coh53, Coh55, Dra52, Dro58, Eva53, Gru52, Gur57, Hab55, Har57c, Huz55, Hyr52, IK54, Jen54b, Jen56, Joh59b, Kha59, Lan49a, Lan49b, Lan50a, Lan50b, Les59, Mal56, Mar52a, MH58, Pat54, Pea50a, Pea51c, Pre53, Que59, Ram58, Ram59, SK57, Ste57, Stu55b, Tho51b, Tho54, Wat57b, BO58, Har57b, Mul51]. Diversity [Her58a]. Dixon [Ano57c, Wis52a]. Dlya [Joh55c]. Doctrine [BB58]. Doob [Dav54e]. Doolittle [Gue53]. Doses [AC55]. Double [Cox52]. Drawn [Gay50a, Gay51, GW47, GW50]. DuBois [Dur59a]. Duncan [Cox53b]. duration [Tha57]. Dwight [Dav58d]. Dynamics [Dav59g, Mal59d].

Each [Ste52]. Early [Lie54, Wil56a, Tha57]. Ecology [Eva53, Ske52, Tho55a]. Econometrics [Dav57b, Fos55a, Dur53]. Edgeworth [BD52, Pea59c]. Editorial [Ano52c]. Educational [Dav57d]. Edwin [Mal59d]. Effect [Bar51, Coo57, DJ51, Hor53, Mac53, Que57, Sri58, Sri59, Fin50]. Effects [Cox58a]. Efficiencies [Fin58b, Nai50]. Efficiency [BM55, Har59, She50, She51, Sun54, VN59, Wil58a]. Efficient [Dur59b].

Einführung [Bar57f]. Elasticity [Mal59d]. Electrical [Joh54b]. Elementary [Bar58a, Dav56e, Wel54, Arm54, Boo57, Dav56b, Mal59a]. Ellinger [Stu56b]. Elliptic [Win57]. Empirical [Hac58]. Engineering [Moo53a, Joh54b]. Entry [Ste51a]. Epidemic [Bai50, Bai53a, Has54, RM55, Whi55, Fos55c]. Epidemics [Bai53b, Moo58c]. Equality [Bar58e]. Equally [Gue54]. Equation [Fin58b]. Equations [BH54, Cha54, Coh57, Dur57, LC51, Mal59d, Phi59, Ste51b, Wil58b, Bar59d]. Equidistant [Har51]. Equivalence [Bar58e]. Errata [DK58b]. Error [Cur57, Ehr50, GF57, Med56]. Essay [BB58]. Essentials [Dav57d]. Estimate [BC58, Ber57, HS56, Nag50]. Estimates [Ban50, Kam54a]. Estimating [Ans56, Bai51, Bai56a, Bai56b, Ber57, CM51, Coh53, Coh57, Cra53a, Cra53b, Ghu50, GR54, LD53, LC51, Nai50, Sun54]. Estimation [AB54, Arm58, Big59, Cha55, CR58, Cox52, Dav57e, Dow53, DK51b, Dur59b, Dur59c, Ehr50, Fin52, Goo53, Gul55, Gup52, Gut58, HM54, Ken54a, LC51, Les52, LCC53, Lin53, Llo52, Llo59, Lon51a, MP54, MH58, Moo52, Moo56c, Mor52, Olk58, Phi59, Que56, Saw59, Stu53, Stu55a, Stu57, Bra55, Dar58, Dar59a]. Estimator [Ano56c, Bar56e, Dar56, Gan55b, Pat50]. Estimators [Fin58b, Joh50, She58, She59a, She59b, She61]. Evaluation [Dav53a]. Events [Bar56b, Bar58e, CS53]. Evidence [Eva53, Dav51a]. Evolutionary [Whi51]. Exact [FH51, Han55a, Han55b, Huz55, Kup52, Lan50c, Tay56, Wil52b, Les55, Mul51]. Examination [Bro56a]. Example [LCC53]. examples [Lor54a]. Exercises [Bai58, Dav54b]. Existence [Bla56]. Expansion [Ano55e, Wis54a]. Expectation [CG59]. Expectations [Lan50c, Nas59]. Expected [Gru51]. Experiment [Cox53a, Cox54a, Dav58b]. Experimental [Big59, BH54, But56, Cox51, Cox57a, Cox57b, Eva53, Pea56, Ray56, Wil52d, Joh51, Joh58a, Joh59a, Pea56]. Experimentation [Wil58b]. Experimenting [Pea57a]. Experiments [BL59, Cur57, Lan52, Les57, MM59, Pat52, Tay50, Wil52a, Yat55b, Joh51, Joh54c, Mal59c, McM55, Mos53]. Exploited [Dav59g]. Exponential [AP54, Bar57a, Har59, Pat58, PL59]. Exponentially [MH58]. Expressed [PM51]. Expression [Daw54, Has54]. Extended [May52]. Extension [Joh59b, Lah53, Sil57, Toc50, Lat53]. Extrema [BO59, BO61]. Extreme [Dav56f, Dro58, Lie54, Nai52, PT59, Pil59, Rub56a].

F [Dav57d, Dav59c, Est59, Joh55a, Joh58b, Jon56, Smi51, Stu54, Wil54, Wil58a]. Factor [Bar51, Box52]. Factorial [Dav58b, Joh54c, MM59, Wil52a, Wis54c]. Factors [Dav58b, Tho53]. Factory [Moo58b]. Failure [MH58]. Familiar [Chu55a]. Families [Bai56a, Bai56b]. Family [Mar52a]. Features [Pow55]. Federer [Pea56]. Fejezetek [BS57]. Felix [Bar58a]. Feller [Dav58a, Mor51a]. Fernordnung [Fuc54]. Fiducial [Ans57, Ste54, Ste50]. Field [Cra53a, HP52]. Finite [AL53, Ani55, Ani56, Dav59c, Ken52a, Mil59, Olk58, Wis52b, Wis59a].

Finney [Hea53, Pea56, Hea53, Lat53]. First [Box52, Pen59, Stu50, Pea51b]. Fish [Dav59g]. Fisher [Bar50a, Bar57b, Hyr50, Pea51a, Tuk57]. Fisher-Behrens [Bar50a]. Fisz [Ela59]. Fit [FH51, Pat54, Wal56, Wat57b, Whi52]. Fitting [Bra58, Gru52, Gue53, Gue54, Gue56, Har51, Har59, Hut56, Pat58, WM53, PL59]. Fixed [Hoo58]. fluctuations [Pat57]. Flying [Cra53b]. Forestry [Fin58a]. Form [Dav50a, Har50b, Lan57]. Forms [Ait50, Ait59, BO59, BO61, Huz55, Lei58b, Nel51]. Formula [Pla54]. Formulae [AA54a, DJ54, Dro58, Har50b, Bar55b]. Formulas [Moo53a]. Forsythe [Mal59d]. Fortet [Mal59d]. Foundations [Dav52d]. Fraction [She54]. Fractional [Dav58b]. Francesco [Dav58c]. Frank [Wis52a]. Franklin [Box55]. Franz [Mal59a]. Fraser [Stu58]. Free [Kam56, Suk58, Jon54]. Freedom [Tho53, Wis56b]. Frequencies [Goo53, Gru51, Tan58]. Frequency [Bar52a, Bar53b, Coh53, Gay51, Pat54]. Front [Ano50d, Ano50e, Ano51d, Ano51e, Ano52d, Ano52e, Ano53e, Ano53f, Ano54e, Ano54f, Ano55j, Ano55k, Ano55l, Ano56e, Ano56f, Ano57g, Ano57h, Ano58c, Ano58d, Ano59c, Ano59d]. Fulfilment [Joh57b]. Function [Ano55m, Aro50, Ber57, Cad52a, DJ51, Dav53b, Daw54, Gil56, HF51, Joh55c, Lei58a, Lev59, Pat48, Pat59, PH51, Wis50, Wis55, KL59, Sri58]. Functional [Bar57d, Edw59, Ken51, Ken52b, Lin53, Wil55b]. Functions [Bar54a, Bar55b, Bar56c, DK51a, DK53, DK55, DK58b, Joh52b, Kab58, Lei58b, Sim55, Wal56, Wil55b, Hyr50, Bar53a, Est59]. Fundamental [DJ54]. Fundamentals [Mal59b]. Funktsii [Joh55c]. Furnish [Ban50]. Further [Cox58b, Dav51b, Har57a, Hui58, MPW53, May54, May58, RG57, RG58, RG61, TW54, TWJ56].

G [Ano57e, Ans58, Ban58, Bar55b, Cor58, Dav50c, Dav54b, Dav54c, Dav55c, Dav57d, Gum58, Joh51, Joh58a, Mal59d, Moo55b, Pea59a, Pen59, Smi59a, Stu56b, Wel58, Whi51, Wis59a]. Galton [Hod55]. Gambler [Moh55]. Games [Mor56]. gaming [Dav55d]. Gas [Mal59d]. Gass [Vaj59a]. Gause [Les57]. Geary [Lah53]. General [Bai53a, Bro54, Has54, Law56a, Sun53b, Fos55c, Lor54a]. Generalization [AS57, PS59, DS54]. Generalizations [RM56]. Generalized [Ano56d, FR57, Fos57, Fos58, Gur57, Ram59, RK50, Wil55a, AA54b]. Generating [Lei58b]. Generation [Pow55]. Genetics [Kal57, Smi56, Owe59]. Geometric [Cle59]. Geometry [Dav59a, DK51b, Win57]. George [Dav59c, Dav59a, Mal59d]. Gerhard [Dur53]. Gertrude [Joh51]. Gesetze [Dav52d]. GI [Con59]. GI/M/1 [Con59]. Gini [Kam53c]. Girshick [Mor56]. Given [Dav50d, Saw59]. Gleser [Ban58]. Goldschmidt [Kal57]. Good [Dav51a]. Goodier [Mal59d]. Goodman [Bar59f]. Goodness [FH51, Pat54, Wal56, Wat57b]. Goodness-Of-Fit [Wat57b]. Gowen [Smi55b]. Gram [BD52, She51]. Gram-Charlier [BD52, She51]. Graphical [Pre53, Roy56]. Greenberg [Her58a]. Greenwood [McK51]. Grenander [Jen58]. Gross [Rei56]. Group [Hor53]. Grouped [Gru52, HM54]. Grouping [Gue54, Gue56, Kup52]. Groups [Has57, Jam51]. Growing [Dar53]. Growth [HP52, Whi51]. Gupta [Bar59e]. Gwyther [Bar59e].

H [Bar56a, Bar58d, Bar58c, Bar59e, Bar59d, Bea50, Bea51, Cox53a, Dav55c, Dav56c, Dav59a, Dav59g, Dur59a, Jen59, Joh51, Joh55a, Joh55b, Joh58b, Lew58, Mal59b, Moo56a, Moo58b, Mor57, Owe58, Pea52a, Wil58a]. H. [Joh58f]. Hack [Joh58f]. Hald [Moo53a, Moo53a]. Hall [Mal59d]. Handbook [Stu56a, Dav59d]. Hankel [Lor54a, Lor54b]. Hansen [Dav55c]. Harley [DK58a]. Harold [Ano57f]. Harry [Dav57c]. Hartley [Mul51, PL59]. Hartley-Khamis [Mul51]. Having [Gro51]. Haycocks [Moo56a]. Hazelwood [Wis59a]. Health [Dav57b]. Helen [Fos55b]. Herbert [Dav58d]. Heredity [Owe58]. Herman [Fos55a, Jen55]. Heterogeneity [GF57, Har50a, Lon51a]. Heterogeneous [Cur57, Ehr50]. Heteromorphic [Jow57]. Hewitt [Mal59d]. Higher [Coh53, Jen56]. History [BB58, Ken56, Ken57b, Lor58, Pla58, Roy56, Wil56a, Dav55d, Tha57]. Hodge [Mal59d]. Hoel [Moo55b]. Hogben [Dav52a, Dav56a]. Hohn [Mal59a]. Holt [Dav59g]. Homogeneity [Bar59a, Bar59b, Hal55b, Hor53, Stu55b]. Hooker [Joh55a, Joh58b]. Hotelling [PS59]. Household [Bai53b]. Housing [Bow59]. Hsu [DO51]. Hugh [Joh55b]. Hungarian [Ano57l]. Hurwitz [Dav55c]. Hutchinson [Owe59]. Hypergeometric [Ano55e, Wis54a]. Hypotheses [Bha59, Bro55, Gho55, Jam54, Jen54b, Jen56, Rao53, Rao59]. Hypothesis [Bat50, Bro56a, Jen54b, Dav52b]. Hypothesis-Testing [Dav52b]. Hyrenius [Bar58c].

II [Bra64, Dav55c, Dav57b, Bai56b, Bar59b, Bar53e, BD58a, Ber57, Bra54, Dar59a, DK51a, DW51b, Fos57, Jen56, Ken52b, Ken56, Les52, Lor54b, WH56]. III [Dav57b, Bar55a, Bra55, DK51a, Fos58, HP52, Joh58e, LCC53, Roy56]. Immigration [Dar59a]. Importance [Her58b]. Important [Rao53]. Improvement [Owe59]. Inaccurately [Lon51b]. Inadmissible [Cox54a]. Including [She54]. Incomplete [Aro50, Ban50, BT52, Bra54, Bra64, Cad52a, HF51, Joh55c, Kam53a, Sil57, VN59, Wil57, Wis50, Bra55, KL59]. Increase [GT56]. Increased [GT56]. Incubated [Mac53]. Independence [Ait50, Ait59, Mor50b, Nel51]. Independent [AL53, Ani55, Ani56, Wal50]. Index [Ano55a, Ano55p, Bat50, Dar57b, Her58a, Kat53]. Indicators [Stu56a]. Individual [Pow55, Wil58a]. Industrial [Cox53b, Dav57b, MPW52, McM55]. Industry [Box55, Joh59a, McM57, Vaj59b]. Inefficiency [Chu55a]. Inequalities [She54]. Infection [Bai53b, Bai56c, Has57]. Infectious [Bai56a, Bai56b, Irw54]. Inference [Ano57f, Bar57b, Dav59e, Fos55b]. Information [Ano50f, Ano51f, Ano52f, Ano53g, Ano54g, Ano55q, Ano56k, Ano57m, Ano58h, Ano59h, Bar52c, Lin57a]. Insects [Cra53b]. Inspection [Pag54]. Institute [Ano57l]. Integral [AA54a, Cad51, Chu55b, DJ50, Dav53a, Har54, HP50b, Joh52a, Mal59e, McF56, She54, Wis50, DS55, HP50a]. Integrals [Dav58d, Pla54]. Integration [Dav59f]. Interactions [Wil52a]. Interference [Tan53]. Intermediate [Bar58b]. Internal [Har59]. Interpolations [Tuk55]. Interpretation [Cox58a, Nel54, Wil52a]. Interrelations [Gur57, Sil51]. Interval [CR58]. Intervals [Bar53c, Bar53d, Bar53e, Bar55a, Bul58, Cro56, Cro58a, CG59, Joh57c, MPW52, Pal57, Ste57]. Intra [Bai53b]. Intra-Household [Bai53b]. Introduction [Ans58, Dav55a, Dav58a, Dav59b, RM56, Ano57c, Dav59e, Doi59, Edw59, Joh56, Moo55b, Wis52a]. Invariants [Huz55]. Inventories [Wis59b]. Inverse [Ano55e, Wis50, Wis54a]. Inversion [GP51, Goo50]. Inverting [RS56]. Investigation [Hac58, PL59]. Investigations [Moo56a]. Investment [Stu56b]. Involved [Med56]. Involving [Pat52, Yat55b, Rao59]. Irrationalzahlen [Wat57a]. Irving [Mal59d]. Irwin [Joh54a]. Isolated [Has57]. Iterative [TW54]. IV [Dav57b, DK53, Wil56a]. IX [BB58].

J [Ano57c, Bar59e, Cox53b, Dav50b, Dav51a, Dav52c, Dav54e, Dav56b, Dav59e, Dav59c, Dav59g, Hea53, Joh54a, Mal59d, Moo57a, Moo58c, Pea56, Pea51b, Smi55b, Wel54, Wel58, Whi51, Wis52a]. Jackknife [Efr79, Mil74]. Jacobians [DO51, Olk53]. Janko [Joh58c]. Jaroslav [Joh58c]. Jeffreys [Ano57f]. Jerzy [Dav57b]. John [Dav59b, Lew58, Ano56g, Pea57b]. Johnson [Bea50, Bea51, Bai58, Les59]. Joins [Wis54c]. Joint [Wat56b, Coo51a]. Jordan [BS57, BS57]. Joseph [Fos55b, Owe59]. Judges [Bos56]. Judging [Sch53, Sch69]. Jump [JW59]. Jureen [Fos55a]. Justification [Bar52a, Bar53b].

Kalmus [Owe58]. Kaplan [Bar56c]. Kaplansky [Mal59d]. Karl [Ken59, Hal57]. Karoly [BS57]. Kempthorne [Smi55b, Mos53]. Kendall [Ans58, Dav54b, Gum58, Pea59a, Ano57e, Cor58, Dav50c, Hab55, Sil50, Stu56c, Whi51]. Khamis [Mul51]. Khintchine [Dav52d]. Kingdom [Cor58]. Kitagawa [Joh54c]. Klasszikus [BS57]. Known [ZSP58, Mal56]. Kolmogorov [Dav52d, Ano57d, Joh55c]. Kopal [Moo56b]. Krisen [Stu52]. Kurtosis [Box53b].

L [Bai58, Ban58, Bar56a, Bar56d, Bar57f, Bar57g, Bar58b, Bea50, Bea51, Bow59, Box55, Dav52a, Dav54d, Dav54e, Dav56c, Dav59f, DO51, Doi59, Joh55a, Joh58b, Kes58, Lew57, McM55, McM57, Moo58b, Pow58, Smi55b, Wis59a]. Lag [Gib50]. Lancelot [Dav56a]. Large [AC55, Wis56b, Bra55, Tuk58]. Large-Sample [AC55, Bra55]. Largest [Mau58b, Pil56, PB59]. Lars [Fos55a]. Latent [Bai56a, Bai56b, Law56b]. Latin [Cox58a]. Learning [Jon56]. Least [Dow53, DW50, DW51a, DW51b, Han57, Llo52, Pla50]. Least-Squares [Dow53, Llo52]. Lecture [Efr79]. Lectures [DO51, Bar56c]. Leimanis [Mal59d]. Leipnik [Ken57a]. Length [Her58b]. Leonard [Doi59, Pea52a]. Lesnickeho [Fin58a]. Lev [Fos55b]. Levels [Aro50, Dav58b]. Li [Dav59e, Smi56]. Lieberman [Wel58]. Life [Men58, MH58, Joh55a, Joh58b]. Light [Hai59]. Likelihood [Ano56c, Bar56e, Ber57, Cha54, Gan55b, Gil56, Goo55, Goo57, Goo58, HS56, Law56a, LC51, She50, She58, She59a, She61, Yat55b]. Limit [RK50]. Limit-Theorem [RK50]. Limits [Bul57, Cle59, Hui55, Ste54, Ste50]. Lindley [Wel54, Bar57c]. Line [BD56, IK54, Wis54c]. Linear [Bar57d, BL59, BF58, Gho55, Hui55, Jam54, Jen54b, Jen56, Nai50, Sil51, Ste51b, Tay56, Wil55b, Rao59, Arc58, Vaj59a]. linearization [KL59]. Lines [Pag55a, Toc52b, Wil53]. Linguistic [Her58a]. Linguistics [Her58b]. Lipman [Mal59d]. Literary [Fuc54, Lor58, Wil56a]. Local [Jow55, Jow57]. Location [CS55, HM54, Llo52]. Loeve [Dav55b]. Log [Gru51]. Log-Normally [Gru51]. Logarithmic [Ans50]. Logic [Ken59]. Lognormal [Bro56b, Tho51b]. Long [HP52]. Long-Tailed [HP52]. Longley [Joh55a, Joh58b]. Longley-Cook [Joh55a, Joh58b]. Look [Efr79]. Lush [Smi55b]. Lyle [Moo58b].

M [Ano57e, Ans58, Bar56a, Bar58b, Con59, Cor58, Cox53a, Dav50c, Dav54b, Dav55c, Dav55a, Dav55b, Ela59, Fos55b, Gum58, Jen58, Jen59, Joh51, Joh54c, Joh58a, Kes58, Mal59b, Mor56, Stu56a, Whi51, Wil58a, Wis59b]. Made [Ban50, Mau58b]. Madow [Dav55c]. Maguire [Bar53c]. Mainland [Arm54]. Maintenance [Wis59b]. Major [LD53, McK51]. Manifold [Som59]. Mann [Joh51]. Many [Chu55a]. Marginal [Stu55b, Les55]. Mark [Mor52]. Marked [Cra53b, Gul55]. Markoff [Goo57, Wau59, Dar59c, Goo55, Goo58, Hoe54, Pat57]. Markov [Ano56c, Der56, Gan55b, Gan56, Wau58]. Marriott [Wil58a]. Marshall [Mal59d]. Massey [Ano57c, Wis52a]. Matched [Coc50]. Mathematical [Dav50b, Dav52c, Dav57b, Dav58d, Ela59, Fuc52, Her58a, Mal59d, Moo55b, Moo55c, Moo58c, Mor51b, Pea51a, Ano57l]. Mathematics [Mal59d, Smi55b]. Mathematische [Ano60, Bar57f, Ela59, Bar57g]. Matrices [Goo50, Law56b, RS56, Kes58]. Matrix [DO51, Mal59a, Pil56, PB59, Olk53]. Matter [Ano50b, Ano50c, Ano50d, Ano50e, Ano51a, Ano51b, Ano51d, Ano51e, Ano52a, Ano52b, Ano52d, Ano52e, Ano53a, Ano53b, Ano53e, Ano53f, Ano54a, Ano54b, Ano54e, Ano54f, Ano55b, Ano55c, Ano55d, Ano55j, Ano55k, Ano55l, Ano56a, Ano56b, Ano56e, Ano56f, Ano57a, Ano57b, Ano57g, Ano57h, Ano58a, Ano58b, Ano58c, Ano58d, Ano59a, Ano59b, Ano59c, Ano59d]. Maurice [Pea59a]. Maxima [Cha54, Pal57]. Maximum [Ani55, Ani56, Ano56c, Bar56e, Ber57, Gan55b, HS56, Har57c, Har50a, LC51, She50, She58, She59a, She61, Yat55b, Dav52e]. Maximum-Likelihood [Bar56e, Gan55b, HS56]. Maximum-Minimum [Har57c]. May [Ban50, Dav59d]. McCracken [Moo58a]. Mean [Ano55h, BP59, Cad54, Cox54b, Cro58b, Dav56f, Gup52, Has54, Joh57a, Joh58d, Joh58e, Kam53b, Kam53c, Kam54b, Kam58, Mau58b, Moo56c, Nai52, Pat50, PT59, Pil59, Pla58, Ram58, Ram59, Ste51a, Tay56, Wel51, Wil56b]. Means [BDS54, Dav57e, Gay50b, GR54, Gro51, HG58, Hor53, Hyr52, Jam56, LC51, Les52, LCC53, Mau58a, Mil59, TW54, TWJ56]. Measles [Bai56a, Bai56b]. Measure [PM51]. Measured [Lon51b]. Measurements [Ken54b, Cox53a]. Measures [Cad53, SK57]. Measuring [Stu56a]. Median [Chu55a]. Medical [Arm54]. Meijer [Kab58]. Members [Dar57a, Gul55]. Method [BT52, Bra54, Bra64, Cra53a, Dur59c, Ghu50, Gib50, Gue53, Hea53, Hui55, Law56a, LD53, LC51, Les52, LCC53, Pre53, Que53, Rao53, Sch53, Sch69, She50, Sun53a, vdH52, Bra55, Les55, PL59, She51]. Methodenlehre [Mos59]. Methodology [Stu52]. Methods [Bar57b, Bra58, Bur59, CM51, Dav54a, Dav55c, Dav55a, Efr79, Fin58a, Gue54, Gue56, Joh54b, Les58, McM57, Moo57a, Smi51, Stu54, Stu58, TW54, Vaj59a, Lor54b, Ano57e, Boo57, Bro56c, Dav50c, Fos56, Mor57, Lew57]. Metoder [Bar58c]. Metody [Fin58a]. Meyer [Mor57]. Miller [Bar59e, Dav59f, Wel54]. Million [Moo55a]. Mills [Bro56c]. Minima [Dav57e]. Minimum [Bar56e, Ber57, Har57c]. Minorsky [Mal59d]. Miss [DK58a]. Missing [Big59, Har51]. Mitome [Joh54c]. mittels [Bar56d]. Mixed [Big59, Hoo58, MH58]. Mixed-Up [Big59]. Mixtures [Ash58]. Mobile [Bai51]. Model [Bar58a, Cox58b, Jen54b, Jen56, LG58, Les58, LG59, RM55, Pat57]. Models [Bar57e, Dur59b, Hei55, Joh52b, Jon56, Mal57]. Modifications [Que53]. modified [Lor50]. Moment [Gay51, GW47, GW50, HP51, Kam53c, Ken52a, Lei58b, Mal56, She58, She59a, She59b, She61, Som59, Wis52b, Mul51]. Moment-Problem [Mal56, Mul51]. Moment-Statistics [Ken52a]. Moments [Ani56, Ano54d, BG59, Coh53, Coo57, Dou56, GW47, GW50, Kam53a, Kam54b, Ken57a, Kup52, Rub54, Rub56a, Saw58a, Saw58b, Saw79, She50, Sun53b, Tho51b, Wis54c, She51]. Monographs [Ano59e]. Monomial [Lev59]. Monomial-Monomial [Lev59]. Monte [Mor57]. Moore [Pen59]. Morgan [Lor58]. Morphological [Dav59f]. Morse [Wis59b]. Mortality [Moo56a]. Mortimer [Joh56]. most [Bla56]. Mosteller [Jon56]. Moulton [Dav59c]. Mouse [HP52]. Moving [Dav55e, Dur59b, Wis56a]. Moving-Average [Dur59b, Wis56a]. Moyal [Whi51]. Mr [Pea59c]. Multi [Box52]. Multi-Factor [Box52]. Multiple [AS57, BD57a, BD57b, BDS54, Dar59a, HG58, Jen59, Mor51c, Dar58, KR57]. Multiple-Recapture [Dar59a, Dar58]. Multivariate [Ans58, BF58, Cha50, Dav53a, DO51, Jam54, Kam53a, Olk58, Pil56, PB59, Pla54, RM56, RG57, RG58, RG61, Wat56a, Wil52b, Woo56, DS55, Olk53, Rao59, Dur59a]. Murdoch [Arc58]. Myslivec [Fin58a].

N [Ano57d, Ans58, Bai58, Bea50, Bea51, Box55, Dav52d, Dav55c, Joh55c, Mal59d, Moo58c, Wil54, Wis59a]. Nahordnung [Fuc54]. Nair [San58]. National [Ano55o, Ano56j, Ano57k, Har53, Pea54]. natural [Stu50]. Nature [Cor58]. Negative [Ans50, BO58, Blo54, Blo56, Bra58, Nel54, Sam55]. Neiswanger [Boo57]. Nepolnoi [Joh55c]. Neyman [Dav57b, Bat50, Dav50b, Dav52c, Pea51b, Ske58, Toc50]. Neyman-Pearson [Toc50]. Nielsen [Lew57]. Nievergelt [Bar56d]. Night [Wil58a].

Night-Vision [Wil58a]. Non [AA54a, Ano55g, Ano55h, BDM58, BD58a, Box53a, BL59, Cox54b, Cox58a, DJ51, Dav54f, Gay50a, Gay50b, Gay51, Gro53, Hac58, Hea56, Hil54, Hyr52, Jen56, Mal57, MP58, PH51, Pea59b, RM56, San59, Tuk57, Wel58, Fin50, Sri58, Sri59, Stu58]. Non-Additivity [Cox58a]. Non-Central [AA54a, MP58, PH51, Pea59b, San59, Tuk57, Wel58]. Non-Linear [BL59]. Non-Normal [Ano55g, Ano55h, Cox54b, Dav54f, Gay50a, Gay50b, Gay51, Hac58, Hil54, Hyr52]. Non-Normality [Box53a, DJ51, Gro53, Fin50, Sri58, Sri59]. Non-Null [Jen56, Mal57]. Non-Parametric [RM56, Stu58]. Non-Randomness [BDM58, BD58a]. Non-Zero [Hea56]. noncentral [Har57b]. Nonlinear [Mal59d]. Nonparametric [Ano58e, Dav57a]. Normal [Ait50, Ait59, AL53, Ani55, Ani56, Ano53c, Ano54d, Ano55g, Ano55h, Ano55m, BP59, BDS54, Ber57, BG59, Cad51, Chu55b, Coh55, Coh57, Cox54b, Dan52, Dav53a, Dav54f, DHP54, Dav57e, Dra52, Gay50a, Gay50b, Gay51, GW47, GW50, Gro51, Gru52, Gup52, Hac58, HP57, HP51, Hil54, Hyr52, Kam53a, Lan57, Lei58b, Mau58a, McF56, Moo55a, Moo56c, Nai50, Pea52b, Pla54, Pre53, Rub54, Rub56a, Rub56b, Saw58a, Saw58b, Saw59, Saw79, She54, Tay56, Tuk55, Wal56, Wat57b, Woo56, ZSP58, Hyr50, Rub65]. Normality [Box53a, DJ51, GW47, GW50, Fin50, Gro53, Sri58, Sri59]. Normally [Gru51]. Normit [Ber57]. North [DO51]. not-quite [Tuk58]. Notation [Cha50, Kap52]. Note [Ano52c, Ano58e, Ano59e, Bar53c, Bar54b, Bar57d, Box53b, Cha54, Cox56, Dar59c, Dar59b, Dav53a, Dav55e, Dav56d, Dow53, Dur59c, Fin52, Fra50, FH51, GP51, Gut58, HG58, Har54, Joh57a, Joh58d, Kat53, Ken54a, Ken57b, Ker58, Lei58a, Llo59, Mal56, Mar52a, Med56, Mer55, Mur51, Nel51, Olk53, Pea59b, Pea59c, Rei56, RG58, Roy56, Tho54, Toc52a, Wat56a, Whi55, Wil56a, Yat55a, Dav55d, Fos55c, Sil50, Tha57]. Notes [MPW53, Mor50a, Pea50b, Que56]. Null [Jen56, Mal57, Jen54b, Jen56]. Number [AL53, Ani55, Ani56, Der56, Gab59, GT56, Ste52]. Numbers [Hal55a, Les52, Mil59, Pal57, Wis56b, Stu50]. Numerical [Arm58, Les58, Lew57, Mal59d, Fos56, Lew58, Moo56b]. Nuzialita [Smi59a].

O [McM55, McM57, Mos53, Mos59, Smi55b, Wat57a]. Observations [Ban51, Big59, But56, Gue54, Gue56, Jam51, Que59, Wal50, Wil52d]. Observed [Dav57e]. Obtained [LC51, Les52, LCC53, Pat54]. Occupancy [BD59]. Occur [MS59]. Occurrence [Bar58e, Her58b]. Occurring [Bar56b, Bar58e, Pag55b]. Occurs [Pag57]. Odd [Bar52c]. O’Doherty [Wil58a]. Olson [Dav59f]. One [RM55, RB58, Tat55, Bar53e, Mal56]. One-Way [RB58]. Only [Bai56a, Mau58b]. Operation [Moo58b]. Operations [Doi59]. Optimal [Joh57b]. Optimum [Som54]. Order [Ano54d, BG59, Box52, Dar57a, Hod55, Jen56, Llo52, Rub54, Rub56a, Sun53a]. Ordered [Bar59a, Bar59b, Dow53, Hab55, Moo56c, Som59, Jon54]. Ordinate [Ano55m]. Organic [KH50]. Organisms [Pea57a]. Orthogonal [Ban50, WM53]. Oscar [Stu52]. Osgood [Est59]. Other [Ano55n, Ano56h, Ano56i, Ano57i, Ano57j, Ano58f, Ano58g, Ano59f, Ano59g, Dav58d, FH51, Joh55a, Joh58b, Moo56a]. Otherwise [Has57]. Outcome [Whi55]. Outline [Smi55a]. Overflow [Hai59]. Overlapping [Mac53].

P [Bar59e, Dur59a, Joh55a, Joh58b, Mal59d, Moo58b, Pen59, Wel54]. P. [DO51]. Paired [Bos56, BT52, Bra54, Bra64, Dav59j, Wil57, vdH52, Bra55]. Pairs [Mau58b, Tay50]. Pandurang [Dur55]. Panizzon [Smi59a]. Paper [Bar53c, Mac58, Whi55]. Papers [Lie54]. Paradox [Bar57c, Lin57b, Stu55a]. Parallel [Hai58]. Paramecium [Les57]. Parameter [Ano56c, Arm58, Bar53e, CR58, Gan55b, Moo52, Pag55b, Pag57, Ste50, Ste57]. Parameters [Coh53, Dur59b, Fin52, Ghu50, Goo53, Gul55, HM54, LC51, Les52, LCC53, Les59, Llo52, MH58, Phi59, Saw59]. Parametric [Joh52b, RM56, Stu58]. Part [Ber57, DK53, DJ54, DK55, DW51a, Gru51, Ken51, Ken52b]. Partial [AL53, Ani55, Ani56, Goo59, Mal59d, Mor51c, Rei56, Som59]. Partially [Hab55]. Partition [Lan49b, Lan50b, Lan50c]. Partitions [Bar59e]. Parts [DK51a]. Pattern [Ske52]. Patterned [RS56]. Paul [Mal59d, Moo55b]. Pearson [Bar53c, Coh53, Hal57, Joh58e, PM51, Toc50]. Peck [Wis59a]. Pedestrians [May54, May58, Tan51]. Percentage [AA54a, Ano53d, Dav56f, FR57, Fos57, Fos58, May52, Nai52, Pil59, DS55]. Percentages [Coc50]. Period [Con59]. Periodic [CS53, Dav59h]. Periodogram [Bar50b]. Periodograms [BM55]. Periods [Bai56a, Bai56b]. Perks [Moo56a]. Perron [Wat57a]. Personnel [Ban58]. Peter [Dav52b, Jen55]. Petrographic [Bar58a]. Phenomena [Mor50a]. Philip [Wis59b]. Physics [Dav57b, Whi51]. Pillai [Bar59c]. Pirenne [Wil58a]. Place [Lor52, Lor47]. Plane [Ano55f, Whi54]. Planning [Mal59c]. Plant [Tho51a]. Plantation [Fre53]. Plasticity [Mal59d]. play [Tha57]. Playing [Ken57b]. Plot [Cur57, Tay50, Whi56]. Point [Pag55b, Pag57, Tho55a]. Points [AA54a, Ano53d, BD56, Dav56f, FR57, Fos57, Fos58, IK54, Lew53, May52, Nai52, Pac59, PM51, Pil59, Tuk57, Dav52e, DS55]. Poisons [Ash58]. Poisson [Blo56, Bir53, Blo54, Cox53c, Cro58b, CG59, Dar57b, Dou56, HP50b, Joh59b, Kat53, Moo52, Pat57, Ram59, Ste57]. Poisson-Markoff [Pat57]. Polynomial [WM53]. Polynomials [Gue53, Gue54, Gue56, Har51]. Pooling [Lan50c]. Popper [Ken59]. Population [BG59, Cha55, Dar53, Dav50d, Ehr52, Goo53, GT56, Gru51, Gul55, Gup52, Hil54, Jam51, Jam54, Joh58e, Ken52a, LC51, Les52, LCC53, Mau58b, Saw58a, Saw58b, Saw59, Saw79, Sil51, Whi51, Dar58, Wis52b, Joh55b, Smi56]. Populations [Ano54d, Ano55g, Ano55h, AP54, Bai51, BDS54, Bir53, Coh57, Cox54b, Cra53a, Cra53b, Dav54f, Dav57e, Gro51, Hac58, Les57, Mau58a, Olk58, Pea52b, Rub54, Ske51, Tho51a, ZSP58, Dar59b, Fin50, Dav59g]. positive [BD52]. Power [Bat50, Cad53, Dar57b, DJ51, Dav53b, Dou56, HP52, Joh52b, Kha59, Lor50, Pat48, Pat59, PH51, Sun54, Bra55, Sri58, Sri59]. Power-Series [Kha59]. Powerful [Bla56]. Practical [Ban51, LCC53]. Predatory [Bar57e]. Prefatory [Mer55]. Present [Lon51a]. Presentation [Roy56]. Primer [Wil54]. Principle [Pla58]. Principles [Pen59]. Probabilities [Ano55e, Lan49a, Lan50a, RK50, Wis54a, Yat55a, Wis56b]. Probability [AA54a, Ano55m, BS57, BB58, Bar52b, DJ50, Dav56e, Dav57b, Dav58a, Dav59d, Ela59, Har54, IK54, Joh52a, Joh55c, Ken56, Ken57b, Ker58, Les59, Lor58, Mal59d, Mal59e, Pea51b, Pla58, Roy56, Wil56a, Dav55d, DS55, HP50a, Les55, Tha57, Dav50b, Dav52d, Dav52c, Dav57b, Ano57d, Bar58b, Dav51a, Dav55b, Mor51a]. Probit [AS57, Hea53]. Probits [Bla50, Hea56]. Problem [AB54, BB58, Gut58, Hal55a, Har57c, Lan50c, Llo59, Mal56, Moh55, Tan53, Tan58, Toc52a, TWJ56, Mul51, Tha57]. Probleme [Mos59]. Problems [Dav54b, Dav57b, FH51, Gan55a, Kab58, Ken54b, Lon51a, Pag57, Som54, Rao59]. Procedure [BDS54]. Procedures [Arm57, BM55, Bir53, GR54, Moo58d, Moo59]. Proceedings [Dav50b, Dav52c, Dav57b]. Process [Con59, Dar56, Wau55]. Processes [Ano55f, Bir53, CS54, Dar53, Dav55a, Hei55, Pat54, Smi59b, Smi59c, Tho55a, Wau58, Wau59, Whi51, Whi54, Wis56a, Whi51, Bar56a, Dav54e]. Product [Gay51, Rub56a, Som59]. Product-Moment [Gay51, Som59]. Production [McM57, Moo58b]. Products [Bro56b]. Programming [Vaj59b, Vaj59a, Moo58a]. Proneness [Gre50, Joh57d]. Proof [DK58a]. Properties [Ano53c, Bak50, Dra52, Gra52, Har56, Har57a, Jow55, Jow57, Kha59, Lan57, LG58, Les58, LG59, Moo58d, Moo59, Sun53a]. Property [Sil57]. Proportion [Cro56, Cro58a]. Proportions [Pat48, Pat59]. Protozoa [Les57]. Provisioning [Gan55a]. Psychological [Ban58]. Psychometry [Dav57b]. Publication [Ano55i]. Publications [Ano55o, Ano56j, Ano57k, Ano57l, Har53, Pea54].

Quadratic [Ait50, Ait59, BO59, BO61, Lei58b, Nag50, Nel51]. Quadrivariate [McF56]. Quality [Lon51b, Moo57a, Moo58d, Moo59, Cox53b]. Quantal [Ash58, Yat55b]. Quantiles [Cad52b]. Quantitative [AB54, Her58b]. Quantum [Bro55, Bro56a]. Quenouille [Cox53a, Cox53a, Dur59c, Jen59, Mal59b, Pea59c]. Questions [DKS51, Pea50a, Pea51c]. Queueing [Con59, Hai57, Wis59a]. Queues [Doi57, Hai58, HM59, Tan53, Wis59b]. Quick [BC58, CS55, Cox56]. Quickly [Ano55e, Wis50, Wis54a]. quite [Tuk58]. Quota [Joh57b]. Quotient [Bro54]. Quotients [Bro56b].

R [Bar56a, Bar59f, Dav54d, Dav54a, Dav56b, Dav59e, Dav59f, Dav59g, Dav59d, Gum58, Joh58f, Jon56, Kal57, Ken59, Kes58, Mal59c, Pea51a, Wis59a]. Raised [Pea50a, Pea51c]. Rand [Moo55a]. Random [AC55, Dar57a, Gan56, Gay50a, Gay51, GW47, GW50, Gra52, HM59, Joh52b, MT57, Mac58, MS59, Moo55a, Pal57, Que59, Ske51, Dav59i]. Randomization [Ano56d, Joh58f, Wil55a]. Randomized [Ano56d, Bla56, GF57, Wil55a]. Randomly [Bar58e, MS59]. Randomness [Bar54b, Bar56b, BD56, BDM58, BD58a, Moo53b, Tho51a]. Range [AL53, Ano53d, Ano55g, Ano55h, BCT56, Cox54b, Dav51b, Dav53b, Dav54f, DHP54, HP57, Har50c, HP51, Joh52a, LD53, Lor52, May52, Pac59, Pat50, Pea50b, Pea52b, Pil52, Rub56a, Tho53, Tho55b, Tuk55, Lor47, Lor50, Moo57b]. Ranges [Dav55e]. Rank [Ber58, BT52, Bra54, Bra55, Bra64, Coo57, DKS51, FHP57, Hod55, Mor51c, Som59, Stu56c, Sun53b, Ano57e, Dav50c]. Rank-Order [Hod55]. Rankers [Ehr52]. Ranking [BDS54, Mal57, Mau58a, Stu51]. Rao [Dav54a]. Rapid [CM51, Gib50, Hal55b, LD53]. Rate [Bar58e, LC51]. Rates [Mor52]. Ratio [DHP54, Gay50a, Goo55, Goo57, Goo58, Hac58, Har50a, Kam53b, Kam58, Law56a, Med56, Olk58, Tho55b, Dav52e]. Ratios [Ano55m, Dur59c, Jam51, Jam54, Jam56]. Real [Est59, Mil59]. Realizations [Dar59c]. Reasoning [Mal59b]. Recapture [Bai51, Dar59a, Gil56, Ham53, LC51, Les52, LCC53, Mor52, Dar58]. Received [Ano55n, Ano56h, Ano56i, Ano57i, Ano57j, Ano58f, Ano58g, Ano59f, Ano59g]. Reciprocals [Ano55m]. Rectangular [Bro54, Fre53]. Reduced [LD53]. Reduction [Dur59c, Pla54]. Reference [Dav55a, Joh58e, McM57, Nas59]. Regarding [Ano58e, Ano59e, Joh58f]. Region [BH54]. Regions [Box53b, Rao53]. Regression [BC58, Cox58b, Dav59h, DW50, DW51a, DW51b, Dur57, Fin58b, Han57, Har59, Hil54, Hut56, Ken51, Ken52b, Som59, Toc52b, Wat55, WH56, Wil53, Wil58b, KR57, Moo58b]. regressive [Mur51]. Regular [Gan56]. Rejection [BCT56]. Related [Ano58e, Har57c, Men58, Tuk55]. Relation [Bro57, Con59, Har57b, Her58a, Her58b, Lie54, Sun54]. Relations [Ans56, BF58]. Relationship [Ken51, Ken52b, Lin53]. Relationships [Bar57d, Wil55b]. Remark [Ber58]. Remarks [Ste54]. Removal [Cha55]. Renewal [CS54, Smi59b, Smi59c]. Repetitions [Wil57]. Requiring [MM59]. Research [Ano59e, Dav57b, Fin58a, McM57, Wis54b, Dav54d, Dav54a, Doi59]. Residuals [Mor50b, Wal56]. Resnikoff [Wel58]. Resort [Coh53]. Response [Ans56, Hea56]. Responses [AS57, Ash58]. Restricted [Arm57]. Result [Joh57d]. Resulting [Ano53c, Dra52]. Results [May54, May58, Bra55]. Review [Ano57e, Ano57d, Ano57f, Ano57c, Ans58, Arc58, Arm54, BS57, Bai58, Ban58, Bar53a, Bar56a, Bar57b, Bar55b, Bar56c, Bar56d, Bar57f, Bar57g, Bar58a, Bar58d, Bar58b, Bar58c, Bar59f, Bar59c, Bar59e, Bar59d, Bea50, Bea51, Boo57, Bow59, Box55, Bro56c, Bur59, Car59, Cor58, Cox53b, Cox53a, Dav50c, Dav50b, Dav51a, Dav52a, Dav52b, Dav52c, Dav54d, Dav54a, Dav54e, Dav54b, Dav54c, Dav55c, Dav55a, Dav55b, Dav56a, Dav56b, Dav56c, Dav57d, Dav57c, Dav57b, Dav57a, Dav58a, Dav58c, Dav58d, Dav58b, Dav59e, Dav59b, Dav59c, Dav59a, Dav59f, Dav59g, Dav59h, Dav59d, Doi59, Dur53, Dur55, Dur59a, Edw59, Ela59, Est59, Fin58a, Fos55a, Fos55b, Fos56, Gum58, Hea53, Jen55, Jen58, Jen59, Joh54b, Joh54a, Joh54c, Joh55a, Joh55b, Joh55c, Joh56, Joh58a, Joh58b, Joh58c, Joh59a, Jon56]. Review [Kal57, Ken55, Ken59, Kes57, Kes58, Lew57, Lew58, Mal59b, Mal59c, Mal59a, Mal59d, McM55, McM57, Mil74, Moo53a, Moo55b, Moo55c, Moo55a, Moo56a, Moo56b, Moo57a, Moo58b, Moo58a, Moo58c, Mor51a, Mor56, Mor57, Mos53, Mos59, Owe58, Owe59, Pea56, Pea51a, Pea51b, Pea52a, Pea59a, Pen59, Pow58, Smi51, Smi55a, Smi55b, Smi56, Smi59a, Stu52, Stu54, Stu56b, Stu56a, Stu58, Vaj59b, Vaj59a, Wat57a, Wel54, Wel58, Whi51, Wil54, Wil58a, Wis52a, Wis54b, Wis59b, Wis59a]. Reviews [Dav52d, Joh51, Stu52]. Revised [Dav56f]. Richard [Stu56a]. Richter [Bar58d]. Riemann [Win57]. Rietz [Efr79]. Ring [BD58b]. Riordan [Dav59b]. Risks [Med56]. Road [May54, May58, Tan51]. Robert [Mal59d]. Roberts [Dav57c]. Ronald [Bar57b]. Root [Kam53b, Kam58, Pil56]. Roots [Cha54, Law56b, PB59]. Rosenblatt [Jen58]. Rosenbloom [Mal59d]. Rowe [Dav59a]. Roy [Ans58]. Royal [Moo55c]. Ruin [Moh55]. Rule [Ans57]. Runs [BD57a, BD57b, BD58b, Goo58, Gra52, Moo58d, Moo59, BD58a]. Russell [Doi59].

S [Ans58, Bar59c, Bur59, Dav55a, Dav57a, Dav59g, Dav59d, Stu58, Whi51]. Sacle [Tho53]. Salmon [Dav59a]. Salzer [Bar59d]. Same [Ste52]. Sample [AC55, Bar56e, BDS54, Bha59, Chu55a, Dan52, Dav50d, DHP54, Dav56f, GT56, Gru51, Gup52, Kam56, Lon51a, Nai52, PT59, Pil59, Rao53, Saw58a, Saw58b, Saw59, Saw79, Suk58, Tho53, Bra55, Gro53, Jon54, Moo57b, Dav55c]. Samples [Ano54d, Ano55h, Bar56d, Cad52b, Coc50, Coh55, Coh57, Cox54b, Fuc54, Gay50a, Gay50b, Gay51, GW47, GW50, Gil56, Hac58, HP57, HP51, Hil54, Joh58e, Ken52a, Nai50, Pea52b, Rub54, Rub56a, Saw58a, Saw58b, Saw79, Sil51, Ste52, Wil56b, Hyr50, Tuk58, Wis52b]. Sampling [Ans50, Ans57, Bar52b, Cox52, Ehr52, FH54, HS56, Hoo58, Hyr52, Joh57b, Jow55, Lin57a, Mor52, Som54, Sun53a, Coo51a, Kap52, Dav54c, Dur55, Smi51, Stu54]. Saul [Vaj59a]. Scale [HM54, Llo52]. Scaling [Bur59]. Scheme [Jen54b]. Schemes [Jen56, Lin57a, MM59, Pag54]. Schmetterer [Bar57f]. Schools [Pen59]. Science [Dav56b]. Sciences [Ano57l, Wis54b, Dav57a]. Scientific [Bar57b, Ken59, Ano57f, Vaj59b]. Scores [Rub56b, Wil52c, Rub65]. Second [Dav52c, Kam54a, Mac58, RK50]. Secondary [Cha50]. Segments [Wis54c]. Selecting [Cox57b]. Selection [Mau58b]. Selective [Cha55]. Sequence [BDM58, BD58a, Gab59]. Sequences [CS53, Pat52]. Sequential [Ano54c, AP54, Arm57, Arm58, Bak50, Bar52a, Bar53b, Bar56b, BD59, Bha59, Med56, Moo53b, Ray56, Tay56, Rus50, Rus52]. Sequentially [Joh57c]. Serial [Dan56, DW50, DW51a, DW51b, Dur57, Gib50, Han55b, Han57, Jen54a, Lei58a, Mor50b, Wat55, Wat56b, WH56]. Serial-Correlation [Lei58a]. Serially [But56, Lei58b, Wil52d]. Series [Ans50, BM55, Bar58e, Dar57a, Dav52b, Ghu50, Gra52, Gut58, Han55a, Jow55, Jow57, Kha59, Pal57, Rud52, Wal50, Wal56, Whi52, Wis50, Mur51, Que49, Jen55, Jen58, Jen59, Kes57]. Set [BH54, Bro56a, Mil59, Toc52b, Yat55a]. Sets [Ken54b]. Seventh [GW47, GW50]. Several [Big59, CS53, Dar59c, Jam51, Wel51]. Sheppard [Har50b]. Sherman [Bar54b]. Short [DK58a, Har50a]. Short-Cut [Har50a]. Shorter [Ste57]. Sided [Ano54c, Rus52]. Siegel [Dav57a]. Sign [Bar56d, CS55]. Significance [Aro50, Bai56c, Bar52c, Bar57d, FH51, Gay50b, Gro51, Hal55a, Law56b, Law59, Mar52b, WW56, Wil53, Wil55b, Arm55, Lat53]. Significant [Stu56a]. Signs [Rei56]. Similar [Ash58, Rao53]. Simple [Ano53c, Ano56c, Ash58, Bai50, Bha59, Cox53c, Dar53, Dar56, Dra52, Gan55b, Has54, Lon51a, Rao53, RM55, Les55]. Simplified [Bar54a, Bra58, Daw54, Goo58, Har50b]. Simultaneous [BH54, Dur57, Gho55, Ste51b, Wil58b]. Single [Bro56a, DHP54]. Singly [Saw59]. Singularity [Stu57]. Situations [BL59]. Six [PB59]. Size [Bai51, Bai53a, Bha59, Dar57a, Gay50a, Gay51, Whi56]. Sizes [Bar56e, Tho53]. Skew [Sim55]. Slutskii [Joh55c]. Small [Ano55h, Cad52b, Cox54b, Hal55a, Lan50c, Nas59, Pea52b, Tho53, Les55]. Smallest [Pil56]. Smith [Ken55]. Smooth [Fra50]. Smoothed [Gra52]. Smoothing [BM55]. Smoothly [Jow57]. Snyder [Stu56a]. Social [Dav56b]. Society [Moo55c]. Solution [Bar59d, BH54, Coh57, Gut58, Mul51, Tho51b]. solutions [Tha57]. Solving [BB58, Ste51b]. Some [Ano56c, Ans58, Bak50, Ban51, Bir53, Cox51, Cox53c, CS55, DKS51, Dav53b, Der56, Gan55a, Gan55b, Gra52, Gur57, Har56, Huz55, Kab58, Kam53a, Lan57, Les57, MT57, Mal59d, May54, May58, MM59, Moo58d, Moo59, Pea50b, Pla50, Pow55, Ram58, Rao59, RM56, SK57, Som54, Ste54, Tho51a, Whi51, Wil52b, Bra55]. Sources [Cor58]. Sozialwissenschaften [Mos59]. Space [Bar56b, Rao53]. Spaced [Gue54, Gue56]. Spaces [Kes58]. Spatial [Tho55a, Ske52]. Spearman [Ber58]. Special [Dav55a, Joh58e, McM57, Nas59, DS54]. Species [Dar53, Goo53, GT56, Gru51, LG58, LG59]. Specimens [Cra53b]. Spectra [Bar50b, BM55]. Spectral [Wal56, Wis55]. Sphere [WW56]. Spiegelman [Joh56]. Split [Cur57, Tay50]. Split-Plot [Cur57, Tay50]. Spread [Fre53]. Sprowls [Dav56b]. Square [Cox58a, Kam53b, Kam58, San59]. Squares [Dow53, DW50, DW51a, DW51b, Han57, Llo52, Pla50, Rub56b, Rub65]. Stage [GR54]. Standard [DHP54, Gup52, Kam54a, Lor52, PM51, Que59, Tho55b, Lor47]. Standardization [Bar51]. Standards [Ano55o, Ano56j, Ano57k, Har53, Pea54]. States [Der56]. Stationarity [Wis56a]. Stationary [Ano55f, Hei55, Jen55, Jen58, Jow55, Wal56, Whi54]. Statistic [Bar54b, Stu56c]. Statistica [Dav58c]. Statistical [Ait50, Ait59, Ano57c, Ano59e, Ans58, Bar52c, Bar57c, Bar58a, Boo57, Cad54, Dav54a, Dav54b, DJ54, Dav56b, Dav59e, Dro58, Fin58a, GW47, GW50, Gum58, Har57c, Hea53, Joh58c, Jow57, Lan52, Lin57b, Lor58, Mal59b, Mor56, Nel51, Pat50, Pea56, Pen59, Pre53, Ske52, Stu55a, Wel54, Whi51, Wil54, Wil56a, Wis52a, Wis54b, Moo53a, Bai58, Bar53a, Bar57b, Box55, Bro56c, Dav54d, Fos55b, Jen58, Joh54b, McM57, Moo53a, Moo57a, Stu52, Wis54b]. Statisticians [Bar59c]. Statisticke [Fin58a, Joh58c]. Statistics [Ano54d, Ano58e, BB58, BDM58, BG59, BO59, BO61, Cor58, Dav50b, Dav52c, Dav56e, Dav56b, Dav57b, Dav57a, Dav59d, Ela59, Hea53, Huz55, Joh55b, Jow55, Kab58, Ken52a, Ken56, Ken57b, Llo52, Lor54b, Lor58, MT57, Mac58, Nai50, Pea59a, Pla58, Roy56, Rub54, Rub56a, Sil51, Wil56a, AA54b, Coo51a, Coo51b, Dav55d, Kap52, Lor54a, Tha57, Wis52b, Dav57b, Arm54, Bar59f, Cox53b, Dav54b, Dav57d, Moo55b, Pea51a, Pea51b, Pea52a, Smi55a, Stu58, Bea50, Bea51, Dav56c, Dav57c, Smi55b]. Statistik [Ela59, Ano60, Bar57f, Bar57g]. Statistika [Bar58c]. Statistischen [Mos59]. Steps [MS59]. Stichproben [Bar56d]. Stochastic [Bai50, Bai53a, Dav55a, Has54, Has57, Hei55, Jow55, Jow57, LG58, Les58, LG59, Mor50a, Pat54, Phi59, Whi51, Whi55, Wis56a, Fos55c, Dav54e, Jon56]. Stratum [Ste52]. Strength [Lie54]. Strengths [Stu53]. Strictly [CS53]. Structural [Bro57, BF58]. Structure [Ken51, Ken52b]. Stuart [Pea59a]. Student [DS54, DS55, Hyr50, Win57]. Student’-Fisher [Hyr50]. Studentized [Ano53d, Dav56f, May52, Nai52, Pac59, Pil52, PT59, Pil59]. Students [Dav56b]. Studies [Arm58, BB58, Dav55d, Ken56, Ken57b, Lor58, Pla58, Roy56, Ske52, Tha57, Wil56a]. Studio [Smi59a]. Study [Fos55a, Her58b, Irw54, Jen55, Lor58, Wil56a]. Studying [Les58]. Style [Fuc52, Lor58, Wil56a]. Subject [Ano55p]. Subscripts [Cha50]. Subsonic [Mal59d]. Substitutes [Hal55c]. Successes [Gab59]. Successive [Kam53b, Kam58, Pal57, SK57]. Sufficiency [Ano56c, Gan55b, Gan56]. Sufficient [Huz55]. Sukhatme [Dur55]. Sum [Hui58, Rub56b, Rub65]. Sums [AL53, Ani55, Ani56]. Superposition [CS53, CS54]. Survey [Dav55c, MM59]. Surveys [Dur55, Mal59d, Smi51, Stu54, Mal59d]. Survival [HP52]. Susceptibles [Bai56a, Bai56b]. Sylvaticus [HP52]. Symmetric [Chu55a, DK51a, DK53, DK55, DK58b, Lev59, McF56]. Symmetrical [Cad52a]. symmetrically [Dav59i]. Symposium [Dav50b, Dav52c, Dav57b, Whi51, Mor57, Whi51]. System [Joh58f, Les59]. Systematic [Cox51, Nai50, Sil51, Sun53a, Wil56b]. Systems [Bar57e, Dur57, Hab55, Les58, Lon51b, Phi59].

T [Ans58, Ban58, Bar58b, Dav54d, Joh54c, Moo58c, Pea56, Smi55b]. Table [Daw54, Hal55b, HP50a, Lan57, Pac59, Pat48, Pat59, Ste51a, Lat53, Bar59d, Moo55c]. Tables [AA54b, Ano53d, Ano55m, Arm55, BP59, Bar56d, Bar59c, Ber57, Bra54, Bra64, DK51a, DK53, Dav54b, DK55, DK58b, Dou56, HP50b, Joh55c, Joh58c, Lev59, May52, Moo53a, Nai52, Nas59, PM51, Que59, Som59, Ste53, Stu53, TW54, Wil52c, Yat55a, DS54, Les55, Dav59d, Moo55c, Wel54, Wis54b, Wis59a, Bar59e, Dav58d, Joh54c, Wel58]. Tablitsi [Joh55c]. Tabulky [Joh58c]. Tafeln [Bar56d]. Tailed [HP52]. Tau [Hab55]. Taus [Goo59]. Tavole [Smi59a]. Taylor [Edw59]. Teach [Bar59f]. Techniques [AS54, Pen59, Dav54c]. Tensile [Lie54]. Tensor [Kap52]. Term [vdH52]. Terminals [Mal56]. Terms [Gum58]. Test [Ano54c, Bar54b, Bar56b, Bar59a, Bar59b, Bar52b, Bar56d, BDM58, BD58a, Bat50, Bla56, DJ51, Fra50, Goo55, Goo57, Gro51, Hal55b, Han55a, Har50a, Hod55, Hoe54, Hor53, Kam56, Lor52, Med56, MH58, Moo53b, Mor50b, Nas59, Pag55b, Pat48, Pat59, Stu55b, Suk58, VN59, Wal56, Wat57b, Yat55a, Bar50a, Dar59b, Dav59i, Fin50, Jon54, Lor47, Lor50, Moo57b, Rus50, Rus52, Sri58, Sri59]. Testing [Bar57a, Bos56, DW50, DW51a, DW51b, Dur57, Han57, Men58, Dav52b]. Tests [AP54, Bai56c, Bak50, Ban58, Bar52a, Bar53b, BP59, Bar57d, BD56, Bar58e, Bha59, Box53a, Box53b, Cox53c, CS55, Cox56, Dav50d, Dav53b, Dav56d, FHP57, GW47, GW50, Gho55, Goo58, Goo59, Han55b, Jam54, Jen54b, Jen56, Lat53, Law56b, Law59, Mar52b, Pat50, Pea50a, PH51, Pea51c, Rao53, Sun54, Tay56, Tho51a, WW56, Whi52, Wil52b, Wil53, Wil55b, Arm55, Gro53, Que49, Toc50, Bha59]. Tetley [Bea50, Bea51]. Textbook [Bar58b]. Texts [Fuc54]. Their [Ano55m, Bla50, Que57, Coo51a, Joh55b]. Theorem [Cha50, DK58a, Dav56e, GP51, Lah53, RK50]. Theorems [Ano56c, Gan55b, Pla50]. Theoretical [Bar57e, Dav54b, Joh58f, Ske51, Kal57]. Theory [Ans50, Bai56c, Bar52b, Cox56, DKS51, Dav52d, Dav54d, Dav55c, Dav57b, Dav58a, Der56, Doi57, Dur55, Gan55a, HM59, Joh57d, Kam54a, May54, May58, Moo53a, Moo58c, Mor51a, Mor51b, Pea56, Pea59a, Tat55, Lor54a, Toc50, Dav57b, Dav55c, Dav55b, Pea59a, Bur59, Mor56]. There [Dar59a]. Third [Dav57b, Kam53c]. Thomas [BB58]. Thrall [Bar56a, Kes58]. Three [Bai56b, Dav59a, ZSP58, Dar59b]. three-decision [Dar59b]. Ties [Coo57]. Time [Bar53c, Bar56b, BM55, Bar58e, Dav52b, Ghu50, Han55a, HM59, Jen58, MPW52, MH58, MS59, Rud52, Wal56, Whi52, Que49, Jen55, Jen59]. Time-Series [Ghu50, Rud52, Que49, Jen55, Jen59]. Times [Pow55]. Tintner [Dur53]. Tippett [Dav56c, Moo58b, Pea52a]. Tolerance [Fin52]. Topics [Ano58e, Men58]. Torgerson [Bur59]. Tornheim [Kes58]. Total [Bai53a, Les52]. totals [Les55]. Tournaments [Dav59j]. Traffic [Hai59, May54, May58]. Transform [Lor54b, Lor54a]. Transformation [DJ50, Har56, Har57a, Jen54a, Mal59e, San58, Ste53, Pea59c]. Transformations [Ano53c, Blo54, Blo56, DO51, Dra52, Que57, Yat55b, Olk53]. Transformed [Jam55, Har57b]. Transonic [Mal59d]. Trapping [Mor51b]. Treatise [Dav59a]. Treatment [But56, Cad54, Cox54a, DJ54, FH51, MM59, Fos55c]. Treatments [Pat52, Tay50, Yat55b]. Trend [CS55]. Trials [Gab59]. Triangular [Bro54]. Tribolium [CS57]. Trigonometrical [Kes57]. Tropical [Owe59]. Truncated [Bra58, Coh53, Coh55, Coh57, Gru52, Moo52, Sam55, Tho51b]. Two [Ano54c, Bai56a, Bar56d, Bar58e, BDM58, BD58a, BDS54, Coo51b, Cox58b, Dav50d, Dav58b, Gay50b, GR54, Gro51, Hac58, Hai58, Har57c, Has57, Hei55, Hui58, Kam54a, Kam56, Ken54b, LG58, LG59, Lie54, Mau58a, Pat48, Pat59, Pea52b, RB58, Stu55b, Suk58, Tan53, Tat55, TW54, TWJ56, Yat55b, Dar59b, Fin50, Gro53, Moo57b, Pre53, Rus52]. Two-Dimensional [Hei55]. Two-Means [TWJ56]. Two-Sample [BDS54, Kam56, Gro53, Moo57b]. Two-Sided [Ano54c, Rus52]. Two-Stage [GR54]. Two-Way [RB58, Stu55b]. Type [Lon51a, Ske58, Wis56a, Joh58e, Pea59c, She51].

U [Jen58]. U.S. [Dav58b, Ano55o, Ano56j, Ano57k, Har53, Pea54]. Unbiased [Bar56e, Bla56, Ehr50, Nag50]. Uncertain [AC55]. Uncontrolled [Ash58]. Undergraduates [Arc58]. Unequal [Gro51, Hor53]. Unequally [Gue56]. Uniformly [Bla56]. Unimodal [Mal56, BD52]. Uniqueness [Bla56, Joh57d]. United [Cor58]. Univariate [Jam54]. Universal [GW47, GW50]. Universe [GW47, GW50]. Universes [Gay50a, Gay51, Hyr52]. Universities [Pen59]. University [DO51]. Unknown [Ano56c, Bar53e, BDS54, Gan55b, Jam51, Jam54, Mau58a, Pag55b, Pag57]. Upon [BCT56, Que57]. Upper [Dav52e, Dav56f, FR57, Fos57, Fos58, May52, Pac59, Pil59, Tuk57]. Use [Bar54b, Ber57, Bla50, Cox57a, Cox57b, Har50c, Lor52, Lor54b, Pat50, Pat58, Pea50b, Wil52c, Yat55b, Bai53b, Lor47, Lor54a]. used [Fin50]. Useful [DO51, Olk53]. Uses [Bow59]. Using [Cha50, Jow57, Llo52]. Utilization [Cra53b].

V [Bar57c, Dav57c, Dav57b, Dur55, Joh59a, Wel54, DK55, Ken57b]. Vacancies [Fre53]. Vaclav [Fin58a]. Valoszinusegszamitasbol [BS57]. Values [Dro58, Har51, Hui58, Lie54, TWJ56, Wel51]. Variable [Bai53b, Bai56c, Cox57a, Cox57b, CG59, DJ50, Bar56c]. Variables [Dar57a, FH54, Lei58b, Lon51b, Moo56c, Que57, Tat55, Woo56, Dav59i, Est59]. Variance [Ani55, Bar56e, BDS54, Bul57, Bul58, DJ51, Dav51b, Daw54, Gay50a, Har50a, Har50c, Hoo58, Joh52b, Nag50, Nel51, Nel54, Pat50, PH51, Ray56, RM56, Sch53, Sch69, Ste51a, Stu56c, Stu57, Whi56, Wil56b, Sri59]. Variances [Box53a, But56, Cur57, Ehr50, GF57, Gro51, Hor53, Hui55, Hui58, Jam51, Jam54, Mau58a, RB58, Suk58, ZSP58, Fin50]. Variate [Bro54, Jam55, Kam54a, Que53, Coo51a, Mur51]. Variate-Difference [Que53]. Variates [Ait50, Ait59, AL53, Ani55, Ani56, Cox53c, Hoo58, Toc50]. Variation [Ano55h, Cox54b, HP52, Whi56, Owe58]. Vazsonyi [Vaj59b]. Vector [Kes58, Pow58]. Vergleich [Bar56d]. Veroyatnosti [Joh55c]. VI [Tha57]. Vichisleniya [Joh55c]. VII [Pla58]. VIII [Lor58]. Vision [Wil58a]. Vol [Dav55c, Mal59d]. Vol. [Dav57b]. Volume [Ano50f, Ano51f, Ano52f, Ano53g, Ano54g, Ano55q, Ano56k, Ano57m, Ano58h, Ano59h, Mor51a, Pea59a]. Vyzkumnictivi [Fin58a].

W [Ans58, Bar55b, Boo57, Dav54c, Dav55c, Dav57c, Dav58a, Est59, Gum58, Joh58a, Moo56a, Mor51a, Pea56, Smi55a, Smi55b]. Waerden [Bar56d, Bar57g]. Wahrscheinlichkeitsrechnung [Ano60, Dav52d, Ela59]. Wahrscheinlichkeitstheorie [Bar58d]. Wald [Bar53a, BP59]. Walk [HM59, MS59]. Walker [Fos55b]. Walks [Gan56]. Wallis [Dav57c]. Wandering [HP52, HP52]. Warning [Pag55a]. Warren [Bur59]. Waugh [Dav54b]. Way [RB58, Stu55b]. Weighing [Ban50, Ban51]. Weighted [Bla50, Gue53, Hea56, Jam56, Mil59]. Weighting [Dav51a]. West [Doi59]. Whether [Dav50d]. Which [Gru51, MS59, BD52, Bar56e, Cox54a, Mul51, Pag57]. Whittle [Dav52b, Fos55c, Jen55]. Wilcoxon [BDM58, Dav56d]. Wilfred [Ano57c, Bar56c, Wis52a]. William [Joh51]. Winnick [Bow59]. Wishart [Ano56g, Pea57b, Ste51b]. without [Coh53]. Wold [Fos55a, Jen55]. Wolfenden [Joh55b]. Word [Her58b]. Wynn [Bar53c].

Yates [Smi51, Stu54]. Yield [Whi56]. Yourself [Bar59f]. Yule [Cha50, Her58a, Jen54b].

Z [Moo56b]. Zeichentest [Bar56d]. Zemedelskeho [Fin58a]. Zero [Hea56]. zum [Bar56d]. zweier [Bar56d]. Zygmund [Kes57].

References

Abdel-Aty:1954:AFP

[AA54a] S. H. Abdel-Aty. Approximate formulae for the percentage points and the probability integral of the non-central χ^{2} distribution. Biometrika, 41(3/4):538–540, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332731.

Abdel-Aty:1954:TGS

[AA54b] S. H. Abdel-Aty. Tables of generalized k-statistics. Biometrika, 41(1/2):253–260, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333021.

Aitchison:1954:EPQ

[AB54] J. Aitchison and J. A. C. Brown. An estimation problem in quantitative assay. Biometrika, 41(3/4):338–343, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332714.

Andrews:1955:LSB

[AC55] Fred C. Andrews and Herman Chernoff. A large-sample bioassay design with random doses and uncertain concentration. Biometrika, 42(3/4):307–315, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333380.

Aitken:1950:SIQ

[Ait50] A. C. Aitken. On the statistical independence of quadratic forms in normal variates. Biometrika, 37(1/2):93–96, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332151. See note [Nel51] and corrigenda

[Ait59].

Aitken:1959:CSI

[Ait59] A. C. Aitken. Corrigenda: “On the Statistical Independence of Quadratic Forms in Normal Variates”. Biometrika, 46(1/2):279, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332856. See [Ait59].

Anis:1953:RPS

[AL53] A. A. Anis and E. H. Lloyd. On the range of partial sums of a finite number of independent normal variates. Biometrika, 40(1/2):35–42, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333093.

Anis:1955:VMP

[Ani55] A. A. Anis. The variance of the maximum of partial sums of a finite number of independent normal variates. Biometrika, 42(1/2):96–101, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333425.

Anis:1956:MMP

[Ani56] A. A. Anis. On the moments of the maximum of partial sums of a finite number of independent normal variates. Biometrika, 43(1/2): 79–84, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333579.

Anonymous:1950:A

[Ano50a] Anonymous. Addendum. Biometrika, 37(1/2):129, June 1950. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332155.

Anonymous:1950:BMa

[Ano50b] Anonymous. Back matter. Biometrika, 37(1/2):??, June 1950. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332170.

Anonymous:1950:BMb

[Ano50c] Anonymous. Back matter. Biometrika, 37(3/4):??, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332402.

Anonymous:1950:FMa

[Ano50d] Anonymous. Front matter. Biometrika, 37(1/2):??, June 1950. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332140.

Anonymous:1950:FMb

[Ano50e] Anonymous. Front matter. Biometrika, 37(3/4):??, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332370.

Anonymous:1950:VI

[Ano50f] Anonymous. Volume information. Biometrika, 37(3/4):i–vii, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332369.

Anonymous:1951:BMa

[Ano51a] Anonymous. Back matter. Biometrika, 38(1/2):??, June 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332339.

Anonymous:1951:BMb

[Ano51b] Anonymous. Back matter. Biometrika, 38(3/4):??, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332602.

Anonymous:1951:B

[Ano51c] Anonymous. Biometrika: 1901–1951. Biometrika, 38(3/4):267–268, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332573.

Anonymous:1951:FMa

[Ano51d] Anonymous. Front matter. Biometrika, 38(1/2):??, June 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332310.

Anonymous:1951:FMb

[Ano51e] Anonymous. Front matter. Biometrika, 38(3/4):??, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332572.

Anonymous:1951:VI

[Ano51f] Anonymous. Volume information. Biometrika, 38(3/4):i–vii, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332571.

Anonymous:1952:BMa

[Ano52a] Anonymous. Back matter. Biometrika, 39(1/2):??, April 1952. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332491.

Anonymous:1952:BMb

[Ano52b] Anonymous. Back matter. Biometrika, 39(3/4):??, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334046.

Anonymous:1952:EN

[Ano52c] Anonymous. Editorial note. Biometrika, 39(1/2):215, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332490.

Anonymous:1952:FMa

[Ano52d] Anonymous. Front matter. Biometrika, 39(1/2):i–viii, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332457.

Anonymous:1952:FMb

[Ano52e] Anonymous. Front matter. Biometrika, 39(3/4):??, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334017.

Anonymous:1952:VI

[Ano52f] Anonymous. Volume information. Biometrika, 39(3/4):i–viii, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334016.

Anonymous:1953:BMa

[Ano53a] Anonymous. Back matter. Biometrika, 40(1/2):??, June 1953. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333123.

Anonymous:1953:BMb

[Ano53b] Anonymous. Back matter. Biometrika, 40(3/4):??, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333375.

Anonymous:1953:CPD

[Ano53c] Anonymous. Corrigenda: “Properties of Distributions Resulting from Certain Simple Transformations of the Normal Distribution”. Biometrika, 40(1/2):236, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333122. See [Dra52].

Anonymous:1953:CTP

[Ano53d] Anonymous. Corrigenda: “Tables of Percentage Points of the ‘Studentized’ Range”. Biometrika, 40(1/2):236, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333121. See [Nai52].

Anonymous:1953:FMa

[Ano53e] Anonymous. Front matter. Biometrika, 40(1/2):??, June 1953. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333089.

Anonymous:1953:FMb

[Ano53f] Anonymous. Front matter. Biometrika, 40(3/4):??, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333343.

Anonymous:1953:VI

[Ano53g] Anonymous. Volume information. Biometrika, 40(3/4):i–viii, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333342.

Anonymous:1954:BMa

[Ano54a] Anonymous. Back matter. Biometrika, 41(1/2):??, June 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333039.

Anonymous:1954:BMb

[Ano54b] Anonymous. Back matter. Biometrika, 41(3/4):??, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332745.

Anonymous:1954:CTS

[Ano54c] Anonymous. Corrigenda: “On a Two-Sided Sequential t-Test”. Biometrika, 41(3/4):??, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332707. See [Rus52].

Anonymous:1954:CMO

[Ano54d] Anonymous. Corrigenda: “On the Moments of Order Statistics in Samples from Normal Populations”. Biometrika, 41(3/4):??, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332708. See [Rub54].

Anonymous:1954:FMa

[Ano54e] Anonymous. Front matter. Biometrika, 41(1/2):??, June 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333000.

Anonymous:1954:FMb

[Ano54f] Anonymous. Front matter. Biometrika, 41(3/4):??, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332706.

Anonymous:1954:VI

[Ano54g] Anonymous. Volume information. Biometrika, 41(3/4):i–viii, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332705.

Anonymous:1955:AI

[Ano55a] Anonymous. Author index. Biometrika, 42(??):1–22, 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331624.

Anonymous:1955:BMa

[Ano55b] Anonymous. Back matter. Biometrika, 42(??):??, 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331626.

Anonymous:1955:BMb

[Ano55c] Anonymous. Back matter. Biometrika, 42(1/2):??, June 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333462.

Anonymous:1955:BMc

[Ano55d] Anonymous. Back matter. Biometrika, 42(3/4):??, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333417.

Anonymous:1955:CQC

[Ano55e] Anonymous. Corrigenda: “A Quickly Convergent Expansion for Cumulative Hypergeometric Probabilities, Direct and Inverse”. Biometrika, 42(1/2):277, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333458. See [Wis54a].

Anonymous:1955:CSP

[Ano55f] Anonymous. Corrigenda: “On Stationary Processes in the Plane”. Biometrika, 42(1/2):277, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333459. See [Whi54].

Anonymous:1955:CDR

[Ano55g] Anonymous. Corrigenda: “The Distribution of Range in Certain Non-Normal Populations”. Biometrika, 42(1/2):277, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333460. See [Dav54f].

Anonymous:1955:CMC

[Ano55h] Anonymous. Corrigenda: “The Mean and Coefficient of Variation of Range in Small Samples from Non-Normal Populations”. Biometrika, 42(1/2):277, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333461. See [Cox54b].

Anonymous:1955:DPB

[Ano55i] Anonymous. Dates of publication of Biometrika. Biometrika, 42(??): ??, 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331623.

Anonymous:1955:FMa

[Ano55j] Anonymous. Front matter. Biometrika, 42(??):??, 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331621.

Anonymous:1955:FMb

[Ano55k] Anonymous. Front matter. Biometrika, 42(1/2):??, June 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333418.

Anonymous:1955:FMc

[Ano55l] Anonymous. Front matter. Biometrika, 42(3/4):??, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333377.

Anonymous:1955:NPF

[Ano55m] Anonymous. The normal probability function: Tables of certain areaordinate ratios and of their reciprocals. Biometrika, 42(1/2):217–222, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333438.

Anonymous:1955:OBR

[Ano55n] Anonymous. Other books received. Biometrika, 42(3/4):544, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333416.

Anonymous:1955:PID

[Ano55o] Anonymous. Publications of U.S. Department of Commerce, National Bureau of Standards. Biometrika, 42(1/2):276–277, June 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333457.

Anonymous:1955:SI

[Ano55p] Anonymous. Subject index. Biometrika, 42(??):23–59, 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331625.

Anonymous:1955:VI

[Ano55q] Anonymous. Volume information. Biometrika, 42(3/4):i–viii, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333376.

Anonymous:1956:BMa

[Ano56a] Anonymous. Back matter. Biometrika, 43(1/2):??, June 1956. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333614.

Anonymous:1956:BMb

[Ano56b] Anonymous. Back matter. Biometrika, 43(3/4):??, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332950.

Anonymous:1956:CST

[Ano56c] Anonymous. Corrigenda: “Some Theorems and Sufficiency Conditions for the Maximum Likelihood Estimator of an Unknown Parameter in a Simple Markov Chain”. Biometrika, 43(3/4):496–498, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332948. See [Gan55b].

Anonymous:1956:CRA

[Ano56d] Anonymous. Corrigenda: “The Randomization Analysis of a Generalized Randomized Block Design”. Biometrika, 43(1/2):235, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333604. See [Wil55a].

Anonymous:1956:FMa

[Ano56e] Anonymous. Front matter. Biometrika, 43(1/2):??, June 1956. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333572.

Anonymous:1956:FMb

[Ano56f] Anonymous. Front matter. Biometrika, 43(3/4):??, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332902.

Anonymous:1956:JW

[Ano56g] Anonymous. John Wishart. Biometrika, 43(3/4):498, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332949.

Anonymous:1956:OBRa

[Ano56h] Anonymous. Other books received. Biometrika, 43(1/2):240, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333613.

Anonymous:1956:OBRb

[Ano56i] Anonymous. Other books received. Biometrika, 43(3/4):495, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332947.

Anonymous:1956:PID

[Ano56j] Anonymous. Publications of the U.S. Department of Commerce, National Bureau of Standards. Biometrika, 43(3/4):494, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332946.

Anonymous:1956:VI

[Ano56k] Anonymous. Volume information. Biometrika, 43(3/4):i–viii, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332901.

Anonymous:1957:BMa

[Ano57a] Anonymous. Back matter. Biometrika, 44(1/2):??, June 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333291.

Anonymous:1957:BMb

[Ano57b] Anonymous. Back matter. Biometrika, 44(3/4):??, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332900.

Anonymous:1957:BRBc

[Ano57c] Anonymous. Book review: Introduction to Statistical Analysis. by Wilfred J. Dixon; J. Massey. Biometrika, 44(3/4):539, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332898.

Anonymous:1957:BRBa

[Ano57d] Anonymous. Book review: Probability by A. N. Kolmogorov. Biometrika, 44(1/2):298, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333283.

Anonymous:1957:BRB

[Ano57e] Anonymous. Book review: Rank Correlation Methods by M. G. Kendall. Biometrika, 44(1/2):298, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333282.

Anonymous:1957:BRBb

[Ano57f] Anonymous. Book review: Scientific Inference by Harold Jeffreys. Biometrika, 44(3/4):538–539, December 1957. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332897.

Anonymous:1957:FMa

[Ano57g] Anonymous. Front matter. Biometrika, 44(1/2):??, June 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333235.

Anonymous:1957:FMb

[Ano57h] Anonymous. Front matter. Biometrika, 44(3/4):??, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332862.

Anonymous:1957:OBRa

[Ano57i] Anonymous. Other books received. Biometrika, 44(1/2):302, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333290.

Anonymous:1957:OBRb

[Ano57j] Anonymous. Other books received. Biometrika, 44(3/4):540, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332899.

Anonymous:1957:PID

[Ano57k] Anonymous. Publications of the U.S. Department of Commerce, National Bureau of Standards. Biometrika, 44(1/2):300–301, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333288.

Anonymous:1957:PMI

[Ano57l] Anonymous. Publications of the Mathematical Institute of the Hungarian Academy of Sciences. Biometrika, 44(1/2):299, June 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333286.

Anonymous:1957:VI

[Ano57m] Anonymous. Volume information. Biometrika, 44(3/4):i–viii, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332861.

Anonymous:1958:BMa

[Ano58a] Anonymous. Back matter. Biometrika, 45(1/2):??, June 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333088.

Anonymous:1958:BMb

[Ano58b] Anonymous. Back matter. Biometrika, 45(3/4):??, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333234.

Anonymous:1958:FMa

[Ano58c] Anonymous. Front matter. Biometrika, 45(1/2):??, June 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333040.

Anonymous:1958:FMb

[Ano58d] Anonymous. Front matter. Biometrika, 45(3/4):??, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333179.

Anonymous:1958:NRB

[Ano58e] Anonymous. Note regarding bibliography of nonparametric statistics and related topics. Biometrika, 45(3/4):599, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333232.

Anonymous:1958:OBRa

[Ano58f] Anonymous. Other books received. Biometrika, 45(1/2):290, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333084.

Anonymous:1958:OBRb

[Ano58g] Anonymous. Other books received. Biometrika, 45(3/4):600, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333233.

Anonymous:1958:VI

[Ano58h] Anonymous. Volume information. Biometrika, 45(3/4):i–viii, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333178.

Anonymous:1959:BMa

[Ano59a] Anonymous. Back matter. Biometrika, 46(1/2):??, June 1959. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332860.

Anonymous:1959:BMb

[Ano59b] Anonymous. Back matter. Biometrika, 46(3/4):??, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333571.

Anonymous:1959:FMa

[Ano59c] Anonymous. Front matter. Biometrika, 46(1/2):??, June 1959. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332803.

Anonymous:1959:FMb

[Ano59d] Anonymous. Front matter. Biometrika, 46(3/4):??, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333524.

Anonymous:1959:NRS

[Ano59e] Anonymous. Note regarding statistical research monographs. Biometrika, 46(1/2):264, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332834.

Anonymous:1959:OBRa

[Ano59f] Anonymous. Other books received. Biometrika, 46(1/2):277–278, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332855.

Anonymous:1959:OBRb

[Ano59g] Anonymous. Other books received. Biometrika, 46(3/4):501, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333567.

Anonymous:1959:VI

[Ano59h] Anonymous. Volume information. Biometrika, 46(3/4):i–viii, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333523.

Anonymous:1960:CWM

[Ano60] Anonymous. Corrigenda: Wahrscheinlichkeitsrechnung und Mathematische Statistik. Biometrika, 47(1/2):216, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332997. See [Ela59].

Anscombe:1950:STN

[Ans50] F. J. Anscombe. Sampling theory of the negative binomial and logarithmic series distributions. Biometrika, 37(3/4):358–382, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332388.

Anscombe:1956:EBR

[Ans56] F. J. Anscombe. On estimating binomial response relations. Biometrika, 43(3/4):461–464, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332926.

Anscombe:1957:DFA

[Ans57] F. J. Anscombe. Dependence of the fiducial argument on the sampling rule. Biometrika, 44(3/4):464–469, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332877.

Anscombe:1958:BRB

[Ans58] F. J. Anscombe. Book review: A Course in Multivariate Analysis, by M. G. Kendall; An Introduction to Multivariate Statistical Analysis. by T. W. Anderson; Some Aspects of Multivariate Analysis. by S. N. Roy. Biometrika, 45(3/4):594–595, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333223.

Anscombe:1954:STB

[AP54] F. J. Anscombe and E. S. Page. Sequential tests for binomial and exponential populations. Biometrika, 41(1/2):252–253, June 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333020.

Archbold:1958:BRB

[Arc58] J. W. Archbold. Book review: Linear Algebra for Undergraduates, by D. C. Murdoch. Biometrika, 45(3/4):598–599, December 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333230.

Armitage:1954:BRB

[Arm54] P. Armitage. Book review: Elementary Medical Statistics by D. Mainland. Biometrika, 41(1/2):281, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333029.

Armsen:1955:TST

[Arm55] P. Armsen. Tables for significance tests of 2 × 2 contingency tables. Biometrika, 42(3/4):494–511, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333395.

Armitage:1957:RSP

[Arm57] P. Armitage. Restricted sequential procedures. Biometrika, 44(1/2):9– 26, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333237.

Armitage:1958:NSS

[Arm58] P. Armitage. Numerical studies in the sequential estimation of a binomial parameter. Biometrika, 45(1/2):1–15, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333041.

Aroian:1950:LSI

[Aro50] Leo A. Aroian. On the levels of significance of the incomplete beta function and the F -distributions. Biometrika, 37(3/4):219–223, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332374.

Arbous:1954:NTA

[AS54] A. G. Arbous and H. S. Sichel. New techniques for the analysis of absenteeism data. Biometrika, 41(1/2):77–90, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333007.

Aitchison:1957:GPA

[AS57] J. Aitchison and S. D. Silvey. The generalization of probit analysis to the case of multiple responses. Biometrika, 44(1/2):131–140, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333245.

Ashford:1958:QRM

[Ash58] J. R. Ashford. Quantal responses to mixtures of poisons under conditions of simple similar action — the analysis of uncontrolled data. Biometrika, 45(1/2):74–88, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333047.

Bailey:1950:SSE

[Bai50] Norman T. J. Bailey. A simple stochastic epidemic. Biometrika, 37(3/4):193–202, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332371.

Bailey:1951:ESM

[Bai51] Norman T. J. Bailey. On estimating the size of mobile populations from recapture data. Biometrika, 38(3/4):293–306, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332575.

Bailey:1953:TSG

[Bai53a] Norman T. J. Bailey. The total size of a general stochastic epidemic. Biometrika, 40(1/2):177–185, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333107.

Bailey:1953:UCB

[Bai53b] Norman T. J. Bailey. The use of chain-binomials with a variable chance of infection for the analysis of intra-household epidemics. Biometrika, 40(3/4):279–286, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333346.

Bailey:1956:ELIa

[Bai56a] Norman T. J. Bailey. On estimating the latent and infectious periods of measles: I. Families with two susceptibles only. Biometrika, 43(1/2): 15–22, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333574.

Bailey:1956:ELIb

[Bai56b] Norman T. J. Bailey. On estimating the latent and infectious periods of measles: II. Families with three or more susceptibles. Biometrika, 43(3/4):322–331, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332910.

Bailey:1956:STV

[Bai56c] Norman T. J. Bailey. Significance tests for a variable chance of infection in chain-binomial theory. Biometrika, 43(3/4):332–336, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332911.

Bailey:1958:BRB

[Bai58] Norman T. J. Bailey. Book review: Statistical Exercises, by N. L. Johnson. Biometrika, 45(3/4):595–596, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333224.

Baker:1950:PST

[Bak50] A. G. Baker. Properties of some tests in sequential analysis. Biometrika, 37(3/4):334–346, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332386.

Banerjee:1950:HBI

[Ban50] K. S. Banerjee. How balanced incomplete block designs May be made to furnish orthogonal estimates in weighing designs. Biometrika, 37(1/2):50–58, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332146.

Banerjee:1951:SOP

[Ban51] K. S. Banerjee. Some observations on the practical aspects of weighing designs. Biometrika, 38(1/2):248–251, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332330.

Banks:1958:BRB

[Ban58] Charlotte Banks. Book review: Psychological Tests and Personnel Decisions, by L. T. Cronbach; G. C. Gleser. Biometrika, 45(1/2):284– 286, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333075.

Barnard:1950:FBT

[Bar50a] G. A. Barnard. On the Fisher-Behrens test. Biometrika, 37(3/4):203– 207, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332372.

Bartlett:1950:PAC

[Bar50b] M. S. Bartlett. Periodogram analysis and continuous spectra. Biometrika, 37(1/2):1–16, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332141.

Bartlett:1951:ESA

[Bar51] M. S. Bartlett. The effect of standardization on a χ^{2} approximation in factor analysis. Biometrika, 38(3/4):337–344, December 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332580. With an appendix by W.

Ledermann.

Barnard:1952:FJC

[Bar52a] G. A. Barnard. The frequency justification of certain sequential tests. Biometrika, 39(1/2):144–150, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332473.

Bartlett:1952:STT

[Bar52b] M. S. Bartlett. A sampling test of the χ^{2} theory for probability chains. Biometrika, 39(1/2):118–121, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332469.

Bartlett:1952:SSO

[Bar52c] M. S. Bartlett. The statistical significance of odd bits of information. Biometrika, 39(3/4):228–237, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334019.

Barnard:1953:BRB

[Bar53a] G. A. Barnard. Book review: Statistical Decision Functions, by A. Wald. Biometrika, 40(3/4):475–477, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333373.

Barnard:1953:FJS

[Bar53b] G. A. Barnard. The frequency justification of sequential tests — addendum. Biometrika, 40(3/4):468–469, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333368.

Barnard:1953:TIB

[Bar53c] G. A. Barnard. Time intervals between accidents – a note on Maguire, Pearson and Wynn’s paper. Biometrika, 40(1/2):212–213, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333109.

Bartlett:1953:ACIa

[Bar53d] M. S. Bartlett. Approximate confidence intervals. Biometrika, 40(1/2): 12–19, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333091.

Bartlett:1953:ACIb

[Bar53e] M. S. Bartlett. Approximate confidence intervals. II. More than one unknown parameter. Biometrika, 40(3/4):306–317, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333349.

Barnard:1954:SDF

[Bar54a] G. A. Barnard. Simplified decision functions. Biometrika, 41(1/2):241– 251, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333019.

Bartholomew:1954:NUS

[Bar54b] D. J. Bartholomew. Note on the use of Sherman’s statistic as a test for randomness. Biometrika, 41(3/4):556–558, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332738.

Bartlett:1955:ACI

[Bar55a] M. S. Bartlett. Approximate confidence intervals: III. A bias correction. Biometrika, 42(1/2):201–204, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333436.

Barton:1955:BRB

[Bar55b] D. E. Barton. Book review: Bessel Functions and Formulae, by W. G. Bickley. Biometrika, 42(1/2):275, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333456.

Barnard:1956:BRB

[Bar56a] G. A. Barnard. Book review: Decision Processes by R. M. Thrall; C. H. Coombs; R. L. Davis. Biometrika, 43(1/2):236–237, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333606.

Bartholomew:1956:STR

[Bar56b] D. J. Bartholomew. A sequential test of randomness for events occurring in time or space. Biometrika, 43(1/2):64–78, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333578.

Barton:1956:BRBa

[Bar56c] D. E. Barton. Book review: Lectures on Functions of a Complex Variable by Wilfred Kaplan. Biometrika, 43(1/2):239, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333612.

Barton:1956:BRBb

[Bar56d] D. E. Barton. Book review: Tafeln zum Vergleich zweier Stichproben mittels X-Test und Zeichentest (Tables for Comparing Two Samples by X-Test and Sign Test), by B. L. Van Der Waerden; E. Nievergelt. Biometrika, 43(3/4):493–494, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332945.

Barton:1956:CDW

[Bar56e] D. E. Barton. A class of distributions for which the maximumlikelihood estimator is unbiased and of minimum variance for all sample sizes. Biometrika, 43(1/2):200–202, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333592.

Bartholomew:1957:TDE

[Bar57a] D. J. Bartholomew. Testing for departure from the exponential distribution. Biometrika, 44(1/2):253–257, June 1957. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333259.

Bartlett:1957:BRB

[Bar57b] M. S. Bartlett. Book review: Statistical Methods and Scientific Inference by Ronald A. Fisher. Biometrika, 44(1/2):293–295, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333276.

Bartlett:1957:CDV

[Bar57c] M. S. Bartlett. A comment on D. V. Lindley’s statistical paradox. Biometrika, 44(3/4):533–534, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332888.

Bartlett:1957:NTS

[Bar57d] M. S. Bartlett. A note on tests of significance for linear functional relationships. Biometrika, 44(1/2):268–269, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333265.

Bartlett:1957:TMC

[Bar57e] M. S. Bartlett. On theoretical models for competitive and predatory biological systems. Biometrika, 44(1/2):27–42, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333238.

Barton:1957:BRBa

[Bar57f] D. E. Barton. Book review: Einführung in die Mathematische Statistik by L. Schmetterer. Biometrika, 44(1/2):293, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333275.

Barton:1957:BRBb

[Bar57g] D. E. Barton. Book review: Mathematische Statistik, by B. L. Van der Waerden. Biometrika, 44(3/4):536–537, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332893.

Barton:1958:BRBa

[Bar58a] D. E. Barton. Book review: Petrographic Model Analysis — an Elementary Statistical Appraisal, by Felix Chayes. Biometrika, 45(3/4):

591, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333217.

Barton:1958:BRBc

[Bar58b] D. E. Barton. Book review: Probability, An Intermediate Textbook, by M. T. L. Bizley. Biometrika, 45(3/4):593, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333221.

Barton:1958:BRBd

[Bar58c] D. E. Barton. Book review: Statistika Metoder, by H. Hyrenius. Biometrika, 45(3/4):593–594, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333222.

Barton:1958:BRBb

[Bar58d] D. E. Barton. Book review: Wahrscheinlichkeitstheorie, by H. Richter. Biometrika, 45(3/4):592, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333219.

Barton:1958:ETT

[Bar58e] D. E. Barton. On the equivalence of two tests of equality of rate of occurrence in two series of events occurring randomly in time. Biometrika, 45(1/2):267–268, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333066.

Bartholomew:1959:THOa

[Bar59a] D. J. Bartholomew. A test of homogeneity for ordered alternatives. Biometrika, 46(1/2):36–48, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332806.

Bartholomew:1959:THOb

[Bar59b] D. J. Bartholomew. A test of homogeneity for ordered alternatives. II. Biometrika, 46(3/4):328–335, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333530.

Barton:1959:BRBb

[Bar59c] D. E. Barton. Book review: Concise Tables for Statisticians, by K. C. S. Pillai. Biometrika, 46(1/2):270, June 1959. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332842.

Barton:1959:BRBd

[Bar59d] D. E. Barton. Book review: Table for the Solution of Cubic Equations, by H. E. Salzer. Biometrika, 46(3/4):499, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333565.

Barton:1959:BRBc

[Bar59e] D. E. Barton. Book review: Tables of Partitions, by H. Gupta; C. E. Gwyther; J. C. P. Miller. Biometrika, 46(3/4):498–499, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333564.

Barton:1959:BRBa

[Bar59f] D. E. Barton. Book review: Teach Yourself Statistics, by R. Goodman. Biometrika, 46(1/2):269–270, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332841.

Bateman:1950:PID

[Bat50] G. I. Bateman. The power of the χ^{2} index of dispersion test when Neyman’s contagious distribution is the alternate hypothesis. Biometrika, 37(1/2):59–63, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332147.

Barnard:1958:SHP

[BB58] G. A. Barnard and Thomas Bayes. Studies in the history of probability and statistics: IX. Thomas Bayes’s essay towards solving a problem in the doctrine of chances. Biometrika, 45(3/4):293–315, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333180.

Barton:1958:QER

[BC58] D. E. Barton and D. J. Casley. A quick estimate of the regression coefficient. Biometrika, 45(3/4):431–435, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333190.

Bliss:1956:RCB

[BCT56] C. I. Bliss, W. G. Cochran, and J. W. Tukey. A rejection criterion based upon the range. Biometrika, 43(3/4):418–422, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332919.

Barton:1952:CUW

[BD52] D. E. Barton and K. E. Dennis. The conditions under which Gram-Charlier and Edgeworth curves are positive definite and unimodal. Biometrika, 39(3/4):425–427, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334037.

Barton:1956:TRP

[BD56] D. E. Barton and F. N. David. Tests for randomness of points on a line. Biometrika, 43(1/2):104–112, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333583.

Barton:1957:CMR

[BD57a] D. E. Barton and F. N. David. Corrigenda: “Multiple Runs”. Biometrika, 44(3/4):534, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332890. See [BD57b].

Barton:1957:MR

[BD57b] D. E. Barton and F. N. David. Multiple runs. Biometrika, 44(1/2): 168–178, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333249. See

corrigenda [BD57a].

Barton:1958:NRSb

[BD58a] D. E. Barton and F. N. David. Non-randomness in a sequence of two alternatives: II. Runs test. Biometrika, 45(1/2):253–256, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333062.

Barton:1958:RR

[BD58b] D. E. Barton and F. N. David. Runs in a ring. Biometrika, 45 (3/4):572–578, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333207.

Barton:1959:SO

[BD59] D. E. Barton and F. N. David. Sequential occupancy. Biometrika, 46(1/2):218–223, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332824.

Barton:1958:NRSa

[BDM58] D. E. Barton, F. N. David, and C. L. Mallows. Non-randomness in a sequence of two alternatives: I. Wilcoxon’s and allied test statistics. Biometrika, 45(1/2):166–180, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333054.

Bechhofer:1954:TSM

[BDS54] Robert E. Bechhofer, Charles W. Dunnett, and Milton Sobel. A twosample multiple decision procedure for ranking means of normal populations with a common unknown variance. Biometrika, 41(1/2):170– 176, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333014.

Beard:1950:BRB

[Bea50] R. E. Beard. Book review: Statistics, by N. L. Johnson; H. Tetley. Biometrika, 37(3/4):453–454, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332401.

Beard:1951:BRB

[Bea51] R. E. Beard. Book review: Statistics, by N. L. Johnson; H. Tetley. Biometrika, 38(3/4):484–485, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332600.

Berkson:1957:TUE

[Ber57] Joseph Berkson. Tables for use in estimating the normal distribution function by normit analysis: Part I. Description and use of tables. Part II. Comparison between minimum normit χ^{2} estimate and the maximum likelihood estimate. Biometrika, 44(3/4):411–435, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332873.

Bergstrom:1958:RSR

[Ber58] Harald Bergstrom. A remark on Spearman’s rank correlation coefficient. Biometrika, 45(1/2):273–274, June 1958. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333069.

Brown:1958:MLS

[BF58] R. L. Brown and F. Fereday. Multivariate linear structural relations. Biometrika, 45(1/2):136–153, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333052.

Bose:1959:MOS

[BG59] R. C. Bose and Shanti S. Gupta. Moments of order statistics from a normal population. Biometrika, 46(3/4):433–440, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333540.

Box:1954:CRS

[BH54] G. E. P. Box and J. S. Hunter. A confidence region for the solution of a set of simultaneous equations with an application to experimental design. Biometrika, 41(1/2):190–199, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333016.

Bhate:1959:ADS

[Bha59] D. H. Bhate. Approximation to the distribution of sample size for sequential tests: I. Tests of simple hypotheses. Biometrika, 46(1/2): 130–138, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332815.

Biggers:1959:EMM

[Big59] J. D. Biggers. The estimation of missing and mixed-up observations in several experimental designs. Biometrika, 46(1/2):91–105, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332811.

Birnbaum:1953:SPC

[Bir53] Allan Birnbaum. Some procedures for comparing Poisson processes or populations. Biometrika, 40(3/4):447–449, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333360.

Box:1959:DEN

[BL59] G. E. P. Box and H. L. Lucas. Design of experiments in non-linear situations. Biometrika, 46(1/2):77–90, June 1959. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332810.

Black:1950:WPT

[Bla50] A. N. Black. Weighted probits and their use. Biometrika, 37(1/2):158– 167, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332159.

Blank:1956:EUU

[Bla56] A. A. Blank. Existence and uniqueness of a uniformly most powerful randomized unbiased test for the binomial. Biometrika, 43(3/4):465– 466, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332927.

Blom:1954:TBN

[Blo54] Gunnar Blom. Transformations of the binomial, negative binomial, Poisson and χ^{2} distributions. Biometrika, 41(3/4):302–316, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332711. See corri-

genda [Blo56].

Blom:1956:CTB

[Blo56] Gunnar Blom. Corrigenda: “Transformations of the Binomial, Negative Binomial, Poisson and χ^{2} Distributions”. Biometrika, 43(1/2): 235, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333603. See

[Blo54].

Bartlett:1955:EPS

[BM55] M. S. Bartlett and J. Medhi. On the efficiency of procedures for smoothing periodograms from time series with continuous spectra. Biometrika, 42(1/2):143–150, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333431.

Bliss:1958:NBD

[BO58] C. I. Bliss and A. R. G. Owen. Negative binomial distributions with a common k. Biometrika, 45(1/2):37–58, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333044.

Bush:1959:EQF

[BO59] K. A. Bush and I. Olkin. Extrema of quadratic forms with applications to statistics. Biometrika, 46(3/4):483–486, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333549. See corrigenda [BO61].

Bush:1961:CEQ

[BO61] K. A. Bush and I. Olkin. Corrigenda: “Extrema of Quadratic Forms with Applications to Statistics”. Biometrika, 48(3/4):474–475, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332777. See [BO59].

Booker:1957:BRB

[Boo57] H. S. Booker. Book review: Elementary Statistical Methods by W. A. Neiswanger. Biometrika, 44(1/2):296–297, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333279.

Bose:1956:PCD

[Bos56] R. C. Bose. Paired comparison designs for testing concordance between judges. Biometrika, 43(1/2):113–121, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333584.

Bowley:1959:BRB

[Bow59] Marian Bowley. Book review: American Housing and Its Uses, by L. Winnick. Biometrika, 46(1/2):273, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332849.

Box:1952:MFD

[Box52] G. E. P. Box. Multi-factor designs of first order. Biometrika, 39(1/2): 49–57, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332462.

Box:1953:NNT

[Box53a] G. E. P. Box. Non-normality and tests on variances. Biometrika, 40(3/4):318–335, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333350.

Box:1953:NRT

[Box53b] G. E. P. Box. A note on regions for tests of kurtosis. Biometrika, 40(3/4):465–468, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333367.

Box:1955:BRB

[Box55] G. E. P. Box. Book review: Statistical Analysis in Chemistry and the Chemical Industry, by C. A. Bennett; N. L. Franklin. Biometrika, 42(3/4):542, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333413.

Barraclough:1959:TWT

[BP59] Elizabeth D. Barraclough and E. S. Page. Tables for Wald tests for the mean of a normal distribution. Biometrika, 46(1/2):169–177, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332819.

Bradley:1954:RAI

[Bra54] Ralph Allan Bradley. Rank analysis of incomplete block designs: II. Additional tables for the method of paired comparisons. Biometrika, 41(3/4):502–537, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332730. See corrigenda [Bra64].

Bradley:1955:RAI

[Bra55] Ralph Allan Bradley. Rank analysis of incomplete block designs. III. Some large-sample results on estimation and power for a method of paired comparisons. Biometrika, 42(3/4):450–470, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333391.

Brass:1958:SMF

[Bra58] W. Brass. Simplified methods of fitting the truncated negative binomial distribution. Biometrika, 45(1/2):59–68, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333045.

Bradley:1964:CRA

[Bra64] Ralph A. Bradley. Corrigendum: “The Rank Analysis of Incomplete Block Designs. II. Additional Tables for the Method of Paired Comparisons”. Biometrika, 51(1/2):288, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334226. See [Bra54].

Broadbent:1954:QRT

[Bro54] S. R. Broadbent. The quotient of a rectangular or triangular and a general variate. Biometrika, 41(3/4):330–337, December 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332713.

Broadbent:1955:QH

[Bro55] S. R. Broadbent. Quantum hypotheses. Biometrika, 42(1/2):45–57, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333421.

Broadbent:1956:EQH

[Bro56a] S. R. Broadbent. Examination of a quantum hypothesis based on a single set of data. Biometrika, 43(1/2):32–44, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333576.

Broadbent:1956:LAP

[Bro56b] S. R. Broadbent. Lognormal approximation to products and quotients. Biometrika, 43(3/4):404–417, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332918.

Brown:1956:BRB

[Bro56c] Frederick Brown. Book review: Statistical Methods, by Frederick C. Mills. Biometrika, 43(3/4):492–493, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332941.

Brown:1957:BSR

[Bro57] R. L. Brown. Bivariate structural relation. Biometrika, 44(1/2):84– 96, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333242.

B:1957:BRB

[BS57] Cedric A. B. and Piri Smith. Book review: Fejezetek a Klasszikus Valoszinusegszamitasbol. (Chapters on the Classical Probability Calculus) by Jordan Karoly; Charles Jordan. Biometrika, 44(3/4):536, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332892.

Bradley:1952:RAI

[BT52] Ralph Allan Bradley and Milton E. Terry. Rank analysis of incomplete block designs: I. The method of paired comparisons. Biometrika, 39(3/4):324–345, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334029.

Bulmer:1957:ACL

[Bul57] M. G. Bulmer. Approximate confidence limits for components of variance. Biometrika, 44(1/2):159–167, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333248.

Bulmer:1958:CID

[Bul58] M. G. Bulmer. Confidence intervals for distance in the analysis of variance. Biometrika, 45(3/4):360–369, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333184.

Burt:1959:BRB

[Bur59] Cyril Burt. Book review: Theory and Methods of Scaling, by Warren S. Torgerson. Biometrika, 46(3/4):493, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333553.

Butcher:1956:TVE

[But56] J. C. Butcher. Treatment variances for experimental designs with serially correlated observations. Biometrika, 43(1/2):208–212, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333596.

Cadwell:1951:BNI

[Cad51] J. H. Cadwell. The bivariate normal integral. Biometrika, 38(3/4):475– 479, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332596.

Cadwell:1952:ASI

[Cad52a] J. H. Cadwell. An approximation to the symmetrical incomplete beta function. Biometrika, 39(1/2):204–207, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332483.

Cadwell:1952:DQS

[Cad52b] J. H. Cadwell. The distribution of quantiles of small samples. Biometrika, 39(1/2):207–211, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332484.

Cadwell:1953:ADM

[Cad53] J. H. Cadwell. Approximating to the distributions of measures of dispersion by a power of χ^{2}. Biometrika, 40(3/4):336–346, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333351.

Cadwell:1954:STM

[Cad54] J. H. Cadwell. The statistical treatment of mean deviation. Biometrika, 41(1/2):12–18, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333002.

Carrier:1959:BRB

[Car59] N. H. Carrier. Book review: Demografia. Biometrika, 46(1/2):272–273, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332848.

Crow:1959:CIE

[CG59] Edwin L. Crow and Robert S. Gardner. Confidence intervals for the expectation of a Poisson variable. Biometrika, 46(3/4):441–453, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333541.

Chandler:1950:TCS

[Cha50] K. N. Chandler. On a theorem concerning the secondary subscripts of deviations in multivariate correlation using Yule’s notation. Biometrika, 37(3/4):451–452, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332398.

Chanda:1954:NCM

[Cha54] K. C. Chanda. A note on the consistency and maxima of the roots of likelihood equations. Biometrika, 41(1/2):56–61, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333005.

Chapman:1955:PEB

[Cha55] Douglas G. Chapman. Population estimation based on change of composition caused by a selective removal. Biometrika, 42(3/4):279–290, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333378.

Chu:1955:ISM

[Chu55a] John T. Chu. The ‘inefficiency’ of the sample median for many familiar symmetric distributions. Biometrika, 42(3/4):520–521, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333399.

Chu:1955:BNI

[Chu55b] John T. Chu. On bounds for the normal integral. Biometrika, 42(1/2): 263–265, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333443.

Clemans:1959:CLC

[Cle59] Kermit G. Clemans. Confidence limits in the case of the geometric distribution. Biometrika, 46(1/2):260–264, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332833.

Chown:1951:RME

[CM51] L. N. Chown and P. A. P. Moran. Rapid methods for estimating correlation coefficients. Biometrika, 38(3/4):464–467, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332592.

Cochran:1950:CPM

[Coc50] W. G. Cochran. The comparison of percentages in matched samples. Biometrika, 37(3/4):256–266, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332378.

Cohen:1953:EPT

[Coh53] A. C. Cohen, Jr. Estimating parameters in truncated Pearson frequency distributions without resort to higher moments. Biometrika, 40(1/2):50–57, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333096.

Cohen:1955:CST

[Coh55] A. Clifford Cohen, Jr. Censored samples from truncated normal distributions. Biometrika, 42(3/4):516–519, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333397.

Cohen:1957:SEE

[Coh57] A. Clifford Cohen, Jr. On the solution of estimating equations for truncated and censored samples from normal populations. Biometrika, 44(1/2):225–236, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333256.

Conolly:1959:BPR

[Con59] B. W. Conolly. The busy period in relation to the queueing process GI/M/1. Biometrika, 46(1/2):246–251, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332831.

Cook:1951:BVS

[Coo51a] M. B. Cook. Bi-variate k-statistics and cumulants of their joint sampling distribution. Biometrika, 38(1/2):179–195, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332327.

Cook:1951:TAB

[Coo51b] M. B. Cook. Two applications of bivariate k-statistics. Biometrika, 38(3/4):368–376, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332582.

Cooper:1957:ETM

[Coo57] B. E. Cooper. The effect of ties on the moments of rank criteria. Biometrika, 44(3/4):526–527, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332884.

Corlett:1958:BRB

[Cor58] W. J. Corlett. Book review: Sources and Nature of the Statistics of the United Kingdom, by M. G. Kendall. Biometrika, 45(3/4):590–591, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333216.

Cox:1951:SSE

[Cox51] D. R. Cox. Some systematic experimental designs. Biometrika, 38(3/4):312–323, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332577.

Cox:1952:EDS

[Cox52] D. R. Cox. Estimation by double sampling. Biometrika, 39(3/4):217– 227, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334018.

Cox:1953:BRBb

[Cox53a] D. R. Cox. Book review: Associated Measurements, by M. H. Quenouille; The Design and Analysis of Experiment. by M. H. Quenouille. Biometrika, 40(3/4):471–472, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333370.

Cox:1953:BRBa

[Cox53b] D. R. Cox. Book review: Quality Control and Industrial Statistics, by Acheson J. Duncan. Biometrika, 40(1/2):234–235, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333120.

Cox:1953:SSA

[Cox53c] D. R. Cox. Some simple approximate tests for Poisson variates. Biometrika, 40(3/4):354–360, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333353.

Cox:1954:DEW

[Cox54a] D. R. Cox. The design of an experiment in which certain treatment arrangements are inadmissible. Biometrika, 41(3/4):287–295, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332709.

Cox:1954:MCV

[Cox54b] D. R. Cox. The mean and coefficient of variation of range in small samples from non-normal populations. Biometrika, 41(3/4):469–481, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332727. See

corrigenda [Ano55h].

Cox:1956:NTQ

[Cox56] D. R. Cox. A note on the theory of quick tests. Biometrika, 43 (3/4):478–480, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332931.

Cox:1957:CUC

[Cox57a] D. R. Cox. Corrigenda: “The Use of a Concomitant Variable in Collecting an Experimental Design”. Biometrika, 44(3/4):534, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332889. See

[Cox57b].

Cox:1957:UCV

[Cox57b] D. R. Cox. The use of a concomitant variable in selecting an experimental design. Biometrika, 44(1/2):150–158, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333247. See corrigenda [Cox57a].

Cox:1958:IEN

[Cox58a] D. R. Cox. The interpretation of the effects of non-additivity in the Latin square. Biometrika, 45(1/2):69–73, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333046.

Cox:1958:TFA

[Cox58b] D. R. Cox. Two further applications of a model for binary regression. Biometrika, 45(3/4):562–565, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333203.

Clunies-Ross:1958:IEP

[CR58] C. W. Clunies-Ross. Interval estimation for the parameter of a binomial distribution. Biometrika, 45(1/2):275–279, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333070.

Craig:1953:MEB

[Cra53a] C. C. Craig. On a method of estimating biological populations in the field. Biometrika, 40(1/2):216–218, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333111.

Craig:1953:UMS

[Cra53b] C. C. Craig. On the utilization of marked specimens in estimating populations of flying insects. Biometrika, 40(1/2):170–176, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333106.

Crow:1956:CIP

[Cro56] Edwin L. Crow. Confidence intervals for a proportion. Biometrika, 43(3/4):423–435, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332920. See corrigenda [Cro58a].

Crow:1958:CCI

[Cro58a] Edwin L. Crow. Corrigenda: “Confidence Intervals for a Proportion”. Biometrika, 45(1/2):291, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333086. See [Cro56].

Crow:1958:MDP

[Cro58b] Edwin L. Crow. The mean deviation of the Poisson distribution. Biometrika, 45(3/4):556–559, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333201.

Cox:1953:SSS

[CS53] D. R. Cox and W. L. Smith. The superposition of several strictly periodic sequences of events. Biometrika, 40(1/2):1–11, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333090.

Cox:1954:SRP

[CS54] D. R. Cox and Walter L. Smith. On the superposition of renewal processes. Biometrika, 41(1/2):91–99, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333008.

Cox:1955:SQS

[CS55] D. R. Cox and A. Stuart. Some quick sign tests for trend in location and dispersion. Biometrika, 42(1/2):80–95, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333424.

Cox:1957:DBC

[CS57] D. R. Cox and W. L. Smith. On the distribution of Tribolium Confusum in a container. Biometrika, 44(3/4):328–335, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332865.

Curnow:1957:HEV

[Cur57] R. N. Curnow. Heterogeneous error variances in split-plot experiments. Biometrika, 44(3/4):378–383, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332870.

Daniels:1952:CCS

[Dan52] H. E. Daniels. The covering circle of a sample from a circular normal distribution. Biometrika, 39(1/2):137–143, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332472.

Daniels:1956:ADS

[Dan56] H. E. Daniels. The approximate distribution of serial correlation coefficients. Biometrika, 43(1/2):169–185, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333590.

Darwin:1953:PDB

[Dar53] J. H. Darwin. Population differences between species growing according to simple birth and death processes. Biometrika, 40(3/4):370–382, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333355.

Darwin:1956:BES

[Dar56] J. H. Darwin. The behaviour of an estimator for a simple birth and death process. Biometrika, 43(1/2):23–31, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333575.

Darwin:1957:DBC

[Dar57a] J. H. Darwin. The difference between consecutive members of a series of random variables arranged in order of size. Biometrika, 44(1/2):211– 218, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333254.

Darwin:1957:PPI

[Dar57b] J. H. Darwin. The power of the Poisson index of dispersion. Biometrika, 44(1/2):286–289, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333273.

Darroch:1958:MRC

[Dar58] J. N. Darroch. The multiple-recapture census. I. Estimation of a closed population. Biometrika, 45(3/4):343–359, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333183.

Darroch:1959:MRC

[Dar59a] J. N. Darroch. The multiple-recapture census: II. Estimation when there is immigration or death. Biometrika, 46(3/4):336–351, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333531.

Darwin:1959:NTD

[Dar59b] J. H. Darwin. Note on a three-decision test for comparing two binomial populations. Biometrika, 46(1/2):106–113, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332812.

Darwin:1959:NCS

[Dar59c] J. H. Darwin. Note on the comparison of several realizations of a Markoff chain. Biometrika, 46(3/4):412–419, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333537.

David:1950:AF

[Dav50a] F. N. David. An alternative form of χ^{2}. Biometrika, 37(3/4):448–451, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332397.

David:1950:BRBb

[Dav50b] F. N. David. Book review: Proceedings of the Berkeley Symposium on Mathematical Statistics and Probability, by J. Neyman. Biometrika, 37(1/2):190–191, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332169.

David:1950:BRBa

[Dav50c] F. N. David. Book review: Rank Correlation Methods, by M. G. Kendall. Biometrika, 37(1/2):190, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332168.

David:1950:TCT

[Dav50d] F. N. David. Two combinatorial tests of whether a sample has come from a given population. Biometrika, 37(1/2):97–110, June 1950. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332152.

David:1951:BRB

[Dav51a] F. N. David. Book review: Probability and the Weighting of Evidence, by I. J. Good. Biometrika, 38(3/4):485, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332601.

David:1951:FAR

[Dav51b] H. A. David. Further applications of range to the analysis of variance. Biometrika, 38(3/4):393–409, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332585.

David:1952:BRBa

[Dav52a] F. N. David. Book review: Chance and Choice by Cardpack and Chessboard, by L. Hogben. Biometrika, 39(1/2):213, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332486.

David:1952:BRBb

[Dav52b] F. N. David. Book review: Hypothesis-Testing in Time Series Analysis, by Peter Whittle. Biometrika, 39(1/2):213–214, April 1952. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332487.

David:1952:BRBd

[Dav52c] F. N. David. Book review: Proceedings of the Second Berkeley Symposium on Mathematical Statistics and Probability, by J. Neyman. Biometrika, 39(3/4):440, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334042.

David:1952:BRBc

[Dav52d] F. N. David. Book reviews: Foundations of the Theory of Probability, by A. N. Kolmogorov; Asymptotische Gesetze der Wahrscheinlichkeitsrechnung, by A. Khintchine. Biometrika, 39(1/2):214, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332488.

David:1952:UPM

[Dav52e] H. A. David. Upper 5 and 1% points of the maximum F -ratio. Biometrika, 39(3/4):422–424, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334036.

David:1953:NEM

[Dav53a] F. N. David. A note on the evaluation of the multivariate normal integral. Biometrika, 40(3/4):458–459, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333365.

David:1953:PFS

[Dav53b] H. A. David. The power function of some tests based on range. Biometrika, 40(3/4):347–353, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333352.

David:1954:BRBb

[Dav54a] F. N. David. Book review: Advanced Statistical Methods in Biometric Research by C. R. Rao. Biometrika, 41(1/2):283–284, June 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333034.

David:1954:BRBd

[Dav54b] F. N. David. Book review: Exercises in Theoretical Statistics, by M. G. Kendall; Statistical Tables and Problems. by A. E. Waugh. Biometrika, 41(3/4):565, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332741.

David:1954:BRBe

[Dav54c] F. N. David. Book review: Sampling Techniques, by W. G. Cochran. Biometrika, 41(3/4):565–566, December 1954. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332742.

David:1954:BRBa

[Dav54d] F. N. David. Book review: Statistical Theory in Research by R. L. Anderson; T. A. Bancroft. Biometrika, 41(1/2):283, June 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333033.

David:1954:BRBc

[Dav54e] F. N. David. Book review: Stochastic Processes by J. L. Doob. Biometrika, 41(1/2):284–285, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333036.

David:1954:DRC

[Dav54f] H. A. David. The distribution of range in certain non-normal populations. Biometrika, 41(3/4):463–468, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332726. See corrigenda [Ano55g].

David:1955:BRBb

[Dav55a] F. N. David. Book review: An Introduction to Stochastic Processes with Special Reference to Methods and Applications, by M. S. Bartlett. Biometrika, 42(3/4):539, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333408.

David:1955:BRBc

[Dav55b] F. N. David. Book review: Probability Theory, by M. Loeve. Biometrika, 42(3/4):540, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333409.

David:1955:BRBa

[Dav55c] F. N. David. Book review: Sample Survey Methods and Theory: Vol. I, Methods and Applications: Vol. II, Theory by M. H. Hansen; W. N. Hurwitz; W. G. Madow. Biometrika, 42(1/2):272–273, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333451.

David:1955:SHP

[Dav55d] F. N. David. Studies in the history of probability and statistics. I. Dicing and gaming (a note on the history of probability). Biometrika, 42(1/2):1–15, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333419.

David:1955:NMR

[Dav55e] H. A. David. A note on moving ranges. Biometrika, 42(3/4):512–515, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333396.

David:1956:BRBa

[Dav56a] F. N. David. Book review: Choice and Chance by Cardpack and Chessboard by Lancelot Hogben. Biometrika, 43(1/2):236, June 1956. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333605.

David:1956:BRBb

[Dav56b] F. N. David. Book review: Elementary Statistics for Students of Social Science and Business, by R. C. Sprowls; Basic Statistical Concepts. by J. K. Adams. Biometrika, 43(3/4):490–491, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332938.

David:1956:BRBc

[Dav56c] F. N. David. Book review: Statistics, by L. H. C. Tippett. Biometrika, 43(3/4):493, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332944.

David:1956:NWA

[Dav56d] F. N. David. A note on Wilcoxon’s and allied tests. Biometrika, 43(3/4):485–488, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332934.

David:1956:ASE

[Dav56e] H. A. David. On the application to statistics of an elementary theorem in probability. Biometrika, 43(1/2):85–91, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333580.

David:1956:RUP

[Dav56f] H. A. David. Revised upper percentage points of the extreme Studentized deviate from the sample mean. Biometrika, 43(3/4):449–451, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332922.

David:1957:BRBd

[Dav57a] F. N. David. Book review: Nonparametric Statistics for the Behavioral Sciences, by S. Siegel. Biometrika, 44(3/4):538, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332896.

David:1957:BRBc

[Dav57b] F. N. David. Book review: Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability by Jerzy Neyman; Vol. I. Theory of Statistics. Vol. II. Probability Theory. Vol. III. Astronomy and Physics. Vol. IV: Biology and Problems of Health. Vol. V. Econometrics, Industrial Research and Psychometry. Biometrika, 44(3/4):535–536, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332891.

David:1957:BRBb

[Dav57c] F. N. David. Book review: Statistics: A New Approach by W. Allen Wallis; Harry V. Roberts. Biometrika, 44(1/2):297–298, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333280.

David:1957:BRBa

[Dav57d] F. N. David. Book review: The Essentials of Educational Statistics by F. G. Cornell. Biometrika, 44(1/2):296, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333278.

David:1957:EMN

[Dav57e] H. A. David. Estimation of means of normal populations from observed minima. Biometrika, 44(1/2):282–286, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333272.

David:1958:BRBa

[Dav58a] F. N. David. Book review: An Introduction to Probability Theory and its Applications, by W. Feller. Biometrika, 45(1/2):287, June 1958.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333079.

David:1958:BRBd

[Dav58b] F. N. David. Book review: Fractional Factorial Experiment Designs for Factors at Two Levels by U.S. Dept. of Commerce. Biometrika, 45(3/4):598, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333228.

David:1958:BRBb

[Dav58c] F. N. David. Book review: Statistica, by Francesco Brambilla. Biometrika, 45(3/4):592, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333220.

David:1958:BRBc

[Dav58d] F. N. David. Book review: Tables of Integrals and Other Mathematical Data, by Herbert B. Dwight. Biometrika, 45(3/4):597, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333227.

David:1959:BRBd

[Dav59a] F. N. David. Book review: A Treatise on the Analytic Geometry of Three Dimensions, by George Salmon; C. H. Rowe. Biometrika, 46(1/2):271–272, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332846.

David:1959:BRBb

[Dav59b] F. N. David. Book review: An Introduction to Combinatorial Analysis, by John Riordan. Biometrika, 46(1/2):271, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332844.

David:1959:BRBc

[Dav59c] F. N. David. Book review: Calculus of Finite Differences, by George Boole; J. F. Moulton. Biometrika, 46(1/2):271, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332845.

David:1959:BRBh

[Dav59d] F. N. David. Book review: Handbook of Probability and Statistics with Tables, by R. S. Burington; D. C. May. Biometrika, 46(3/4):497, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333561.

David:1959:BRBa

[Dav59e] F. N. David. Book review: Introduction to Statistical Inference, by J. C. R. Li. Biometrika, 46(1/2):269, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332840.

David:1959:BRBe

[Dav59f] F. N. David. Book review: Morphological Integration, by E. C. Olson; R. L. Miller. Biometrika, 46(1/2):273–274, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332851.

David:1959:BRBf

[Dav59g] F. N. David. Book review: On the Dynamics of Exploited Fish Populations, by R. J. H. Beverton; S. J. Holt. Biometrika, 46(3/4):495, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333556.

David:1959:BRBg

[Dav59h] F. N. David. Book review: Periodic Regression in Biology and Climatology, by C. I. Bliss. Biometrika, 46(3/4):495, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333558.

David:1959:TSD

[Dav59i] F. N. David. The z-test and symmetrically distributed random variables. Biometrika, 46(1/2):123–129, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332814.

David:1959:TPC

[Dav59j] H. A. David. Tournaments and paired comparisons. Biometrika, 46(1/2):139–149, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332816.

Dawson:1954:SEV

[Daw54] Reed B. Dawson, Jr. A simplified expression for the variance of the χ^{2}- function on a contingency table. Biometrika, 41(1/2):280, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333028.

Derman:1956:SAD

[Der56] Cyrus Derman. Some asymptotic distribution theory for Markov chains with a denumerable number of states. Biometrika, 43(3/4):285–294, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332907.

David:1954:DRS

[DHP54] H. A. David, H. O. Hartley, and E. S. Pearson. The distribution of the ratio, in a single normal sample, of range to standard deviation. Biometrika, 41(3/4):482–493, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332728.

David:1950:PIT

[DJ50] F. N. David and N. L. Johnson. The probability integral transformation when the variable is discontinuous. Biometrika, 37(1/2):42–49, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332145.

David:1951:ENN

[DJ51] F. N. David and N. L. Johnson. The effect of non-normality on the power function of the F -test in the analysis of variance. Biometrika, 38(1/2):43–57, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332316.

David:1954:STC

[DJ54] F. N. David and N. L. Johnson. Statistical treatment of censored data. Part I. Fundamental formulae. Biometrika, 41(1/2):228–240, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333018.

David:1951:TSF

[DK51a] F. N. David and M. G. Kendall. Tables of symmetric functions. Parts II and III. Biometrika, 38(3/4):435–462, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332590.

Durbin:1951:GE

[DK51b] J. Durbin and M. G. Kendall. The geometry of estimation. Biometrika, 38(1/2):150–158, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332324.

David:1953:TSF

[DK53] F. N. David and M. G. Kendall. Tables of symmetric functions. Part IV. Biometrika, 40(3/4):427–446, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333359.

David:1955:TSF

[DK55] F. N. David and M. G. Kendall. Tables of symmetric functions. Part V. Biometrika, 42(1/2):223–242, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333439. See errata [DK58b].

Daniels:1958:SPM

[DK58a] H. E. Daniels and M. G. Kendall. Short proof of miss Harley’s theorem on the correlation coefficient. Biometrika, 45(3/4):571–572, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333206.

David:1958:TSF

[DK58b] F. N. David and M. G. Kendall. Tables of symmetric functions — errata. Biometrika, 45(1/2):292, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333087. See [DK55].

David:1951:SQD

[DKS51] S. T. David, M. G. Kendall, and A. Stuart. Some questions of distribution in the theory of rank correlation. Biometrika, 38(1/2):131–140, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332322.

Deemer:1951:JCM

[DO51] Walter L. Deemer and Ingram Olkin. The Jacobians of certain matrix transformations useful in multivariate analysis: Based on lectures of P. L. Hsu at the University of North Carolina, 1947. Biometrika, 38(3/4):345–367, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332581. Based on lectures of P. L. Hsu at the University of North Carolina, 1947.

Doig:1957:BTQ

[Doi57] Alison Doig. A bibliography on the theory of queues. Biometrika, 44(3/4):490–514, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332880.

Doig:1959:BRB

[Doi59] Alison Doig. Book review: Introduction to Operations Research, by C. West Churchman; Russell L. Ackoff; E. Leonard Arnoff. Biometrika, 46(1/2):274–275, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332853.

Douglas:1956:TPP

[Dou56] J. B. Douglas. Tables of Poisson power moments. Biometrika, 43(3/4): 489, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332936.

Downton:1953:NOL

[Dow53] F. Downton. A note on ordered least-squares estimation. Biometrika, 40(3/4):457–458, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333364.

Draper:1952:PDR

[Dra52] J. Draper. Properties of distributions resulting from certain simple transformations of the normal distribution. Biometrika, 39(3/4):290– 301, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334025. See

corrigenda [Ano53c].

Dronkers:1958:AFS

[Dro58] J. J. Dronkers. Approximate formulae for the statistical distributions of extreme values. Biometrika, 45(3/4):447–470, December 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333192.

Dunnett:1954:BGS

[DS54] Charles W. Dunnett and Milton Sobel. A bivariate generalization of Student’s t-distribution, with tables for certain special cases. Biometrika, 41(1/2):153–169, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333013.

Dunnett:1955:API

[DS55] C. W. Dunnett and M. Sobel. Approximations to the probability integral and certain percentage points of a multivariate analogue of Student’s t-distribution. Biometrika, 42(1/2):258–260, June 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333441.

Durbin:1953:BRB

[Dur53] J. Durbin. Book review: Econometrics, by Gerhard Tintner. Biometrika, 40(1/2):234, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333119.

Durbin:1955:BRB

[Dur55] J. Durbin. Book review: Sampling Theory of Surveys with Applications, by Pandurang V. Sukhatme. Biometrika, 42(1/2):273–274, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333452.

Durbin:1957:TSC

[Dur57] J. Durbin. Testing for serial correlation in systems of simultaneous regression equations. Biometrika, 44(3/4):370–377, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332869.

Durbin:1959:BRB

[Dur59a] J. Durbin. Book review: Multivariate Correlational Analysis, by P. H. DuBois. Biometrika, 46(1/2):268–269, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332839.

Durbin:1959:EEP

[Dur59b] J. Durbin. Efficient estimation of parameters in moving-average models. Biometrika, 46(3/4):306–316, December 1959. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333528.

Durbin:1959:NAQ

[Dur59c] J. Durbin. A note on the application of Quenouille’s method of bias reduction to the estimation of ratios. Biometrika, 46(3/4):477–480, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333547.

Durbin:1950:TSC

[DW50] J. Durbin and G. S. Watson. Testing for serial correlation in least squares regression: I. Biometrika, 37(3/4):409–428, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332391. See correc-

tions [DW51a].

Durbin:1951:CPT

[DW51a] J. Durbin and G. S. Watson. Corrections to Part I: “Testing for Serial Correlation in Least Squares Regression: I”. Biometrika, 38(1/2):177– 178, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332326. See

[DW50].

Durbin:1951:TSC

[DW51b] J. Durbin and G. S. Watson. Testing for serial correlation in least squares regression. II. Biometrika, 38(1/2):159–177, June 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332325.

Edwards:1959:BRB

[Edw59] R. E. Edwards. Book review: Introduction to Functional Analysis, by A. E. Taylor. Biometrika, 46(3/4):497–498, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333563.

Efron:1979:RLB

[Efr79] B. Efron. The 1977 Rietz Lecture: Bootstrap methods: Another look at the jackknife. Annals of Statistics, 7(1):1–26, January 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://www.jstor.org/stable/2958830. See [Que49, Que56, Tuk58,

Mil74].

Ehrenberg:1950:UEH

[Ehr50] A. S. C. Ehrenberg. The unbiased estimation of heterogeneous error variances. Biometrika, 37(3/4):347–357, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332387.

Ehrenberg:1952:SPR

[Ehr52] A. S. C. Ehrenberg. On sampling from a population of rankers. Biometrika, 39(1/2):82–87, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332465.

Elandt:1959:BRB

[Ela59] Regina C. Elandt. Book review: Wahrscheinlichkeitsrechnung und Mathematische Statistik (Probability and Mathematical Statistics), by M. Fisz. Biometrika, 46(3/4):494, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333555. See corrigenda [Ano60].

Estermann:1959:BRB

[Est59] T. Estermann. Book review: Functions of Real and Complex Variables, by W. F. Osgood. Biometrika, 46(1/2):270, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332843.

Evans:1953:EEC

[Eva53] D. A. Evans. Experimental evidence concerning contagious distributions in ecology. Biometrika, 40(1/2):186–211, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333108.

Freeman:1951:NET

[FH51] G. H. Freeman and J. H. Halton. Note on an exact treatment of contingency, goodness of fit and other problems of significance. Biometrika, 38(1/2):141–149, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332323.

Fieller:1954:SCV

[FH54] E. C. Fieller and H. O. Hartley. Sampling with control variables. Biometrika, 41(3/4):494–501, December 1954. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332729.

Fieller:1957:TRC

[FHP57] E. C. Fieller, H. O. Hartley, and E. S. Pearson. Tests for rank correlation coefficients. I. Biometrika, 44(3/4):470–481, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332878.

Finch:1950:ENN

[Fin50] D. J. Finch. The effect of non-normality on the z-test, when used to compare the variances in two populations. Biometrika, 37(1/2):186– 189, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332166.

Finney:1952:NEP

[Fin52] D. J. Finney. A note on ‘the estimation of the parameters of tolerance distributions’. Biometrika, 39(3/4):439, December 1952. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334041.

Finney:1958:BRB

[Fin58a] D. J. Finney. Book review: Statisticke Metody Zemedelskeho a Lesnickeho Vyzkumnictivi. [Statistical Methods of Agricultural and Forestry Research], by Vaclav Myslivec. Biometrika, 45(1/2):287, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333078.

Finney:1958:EAE

[Fin58b] D. J. Finney. The efficiencies of alternative estimators for an asymptotic regression equation. Biometrika, 45(3/4):370–388, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333185.

Foster:1955:BRBa

[Fos55a] F. G. Foster. Book review: Demand Analysis. A Study in Econometrics, by Herman Wold; Lars Jureen. Biometrika, 42(3/4):538, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333406.

Foster:1955:BRBb

[Fos55b] F. G. Foster. Book review: Statistical Inference, by Helen M. Walker; Joseph Lev. Biometrika, 42(3/4):540–541, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333410.

Foster:1955:NBW

[Fos55c] F. G. Foster. A note on Bailey’s and Whittle’s treatment of a general stochastic epidemic. Biometrika, 42(1/2):123–125, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333428.

Foster:1956:BRB

[Fos56] F. G. Foster. Book review: Numerical Methods by Andrew D. Booth. Biometrika, 43(1/2):238–239, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333610.

Foster:1957:UPPb

[Fos57] F. G. Foster. Upper percentage points of the generalized beta distribution. II. Biometrika, 44(3/4):441–453, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332875.

Foster:1958:UPP

[Fos58] F. G. Foster. Upper percentage points of the generalized beta distribution. III. Biometrika, 45(3/4):492–503, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333196.

Foster:1957:UPPa

[FR57] F. G. Foster and D. H. Rees. Upper percentage points of the generalized beta distribution. I. Biometrika, 44(1/2):237–247, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333257.

Fraser:1950:NST

[Fra50] D. A. S. Fraser. Note on the χ^{2} smooth test. Biometrika, 37(3/4):447– 448, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332396.

Freeman:1953:SDR

[Fre53] G. H. Freeman. Spread of diseases in a rectangular plantation with vacancies. Biometrika, 40(3/4):287–296, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333347.

Fucks:1952:MAS

[Fuc52] Wilhelm Fucks. On mathematical analysis of style. Biometrika, 39(1/2):122–129, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332470.

Fucks:1954:NFS

[Fuc54] Wilhelm Fucks. On Nahordnung and Fernordnung in samples of literary texts. Biometrika, 41(1/2):116–132, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333010.

Gabriel:1959:DNS

[Gab59] K. R. Gabriel. The distribution of the number of successes in a sequence of dependent trials. Biometrika, 46(3/4):454–460, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333542.

Gani:1955:SPT

[Gan55a] J. Gani. Some problems in the theory of provisioning and of dams. Biometrika, 42(1/2):179–200, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333435.

Gani:1955:STS

[Gan55b] J. Gani. Some theorems and sufficiency conditions for the maximumlikelihood estimator of an unknown parameter in a simple Markov chain. Biometrika, 42(3/4):342–359, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333383. See corrigenda [Ano56c].

Gani:1956:SCR

[Gan56] J. Gani. Sufficiency conditions in regular Markov chains and certain random walks. Biometrika, 43(3/4):276–284, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332906.

Gayen:1950:DVR

[Gay50a] A. K. Gayen. The distribution of the variance ratio in random samples of any size drawn from non-normal universes. Biometrika, 37(3/4):236– 255, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332377.

Gayen:1950:SDB

[Gay50b] A. K. Gayen. Significance of difference between the means of two non-normal samples. Biometrika, 37(3/4):399–408, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332390.

Gayen:1951:FDP

[Gay51] A. K. Gayen. The frequency distribution of the product-moment correlation coefficient in random samples of any size drawn from nonnormal universes. Biometrika, 38(1/2):219–247, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332329.

Graybill:1957:HEV

[GF57] Franklin A. Graybill and John Leroy Folks. Heterogeneity of error variances in a randomized block design. Biometrika, 44(1/2):275–277, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333269.

Ghosh:1955:STL

[Gho55] M. N. Ghosh. Simultaneous tests of linear hypotheses. Biometrika, 42(3/4):441–449, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333390.

Ghurye:1950:MEP

[Ghu50] S. G. Ghurye. A method of estimating the parameters of an autoregressive time-series. Biometrika, 37(1/2):173–178, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332161.

Gibson:1950:RMA

[Gib50] Gordon D. Gibson. A rapid method for ascertaining serial lag correlation. Biometrika, 37(3/4):288–307, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332382.

Gilbert:1956:LFC

[Gil56] N. E. G. Gilbert. Likelihood function for capture-recapture samples. Biometrika, 43(3/4):488–489, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332935.

Good:1950:ICM

[Goo50] I. J. Good. On the inversion of circulant matrices. Biometrika, 37(1/2): 185–186, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332165.

Good:1953:PFS

[Goo53] I. J. Good. The population frequencies of species and the estimation of population parameters. Biometrika, 40(3/4):237–264, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333344.

Good:1955:LRT

[Goo55] I. J. Good. The likelihood ratio test for Markoff chains. Biometrika, 42(3/4):531–533, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333404. See corrigenda [Goo57].

Good:1957:CLR

[Goo57] I. J. Good. Corrigenda: “The Likelihood Ratio Test for Markoff Chains”. Biometrika, 44(1/2):301, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333289. See [Goo55].

Goodman:1958:SRT

[Goo58] Leo A. Goodman. Simplified runs tests and likelihood ratio tests for Markoff chains. Biometrika, 45(1/2):181–197, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333055.

Goodman:1959:PTP

[Goo59] Leo A. Goodman. Partial tests for partial taus. Biometrika, 46 (3/4):425–432, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333539.

Gil-Pelaez:1951:NIT

[GP51] J. Gil-Pelaez. Note on the inversion theorem. Biometrika, 38(3/4):481– 482, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332598.

Ghurye:1954:TSP

[GR54] S. G. Ghurye and Herbert Robbins. Two-stage procedures for estimating the difference between means. Biometrika, 41(1/2):146–152, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333012.

Grant:1952:SPR

[Gra52] Alison M. Grant. Some properties of runs in smoothed random series. Biometrika, 39(1/2):198–204, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332482.

Greenwood:1950:AP

[Gre50] Major Greenwood. Accident proneness. Biometrika, 37(1/2):24–29, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332143.

Gronow:1951:TSD

[Gro51] D. G. C. Gronow. Test for the significance of the difference between means in two normal populations having unequal variances. Biometrika, 38(1/2):252–256, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332331.

Gronow:1953:NNT

[Gro53] D. G. C. Gronow. Non-normality in two-sample t-tests. Biometrika, 40(1/2):222–225, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333114.

Grundy:1951:EFS

[Gru51] P. M. Grundy. The expected frequencies in a sample of an animal population in which the abundances of species are log-normally distributed. Part I. Biometrika, 38(3/4):427–434, December 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332589.

Grundy:1952:FGT

[Gru52] P. M. Grundy. The fitting of grouped truncated and grouped censored normal distributions. Biometrika, 39(3/4):252–259, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334022.

Good:1956:NNS

[GT56] I. J. Good and G. H. Toulmin. The number of new species, and the increase in population coverage, when a sample is increased.

Biometrika, 43(1/2):45–63, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333577.

Guest:1953:DMF

[Gue53] P. G. Guest. The Doolittle method and the fitting of polynomials to weighted data. Biometrika, 40(1/2):229–231, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333117.

Guest:1954:GMF

[Gue54] P. G. Guest. Grouping methods in the fitting of polynomials to equally spaced observations. Biometrika, 41(1/2):62–76, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333006.

Guest:1956:GMF

[Gue56] P. G. Guest. Grouping methods in the fitting of polynomials to unequally spaced observations. Biometrika, 43(1/2):149–160, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333588.

Gulland:1955:EPP

[Gul55] J. A. Gulland. On the estimation of population parameters from marked members. Biometrika, 42(1/2):269–270, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333447.

Gumbel:1958:BRB

[Gum58] E. J. Gumbel. Book review: Dictionary of Statistical Terms, by M. G. Kendall; W. R. Buckland. Biometrika, 45(1/2):283, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333073.

Gupta:1952:EMS

[Gup52] A. K. Gupta. Estimation of the mean and standard deviation of a normal population from a censored sample. Biometrika, 39(3/4):260– 273, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334023.

Gurland:1957:SIA

[Gur57] John Gurland. Some interrelations among compound and generalized distributions. Biometrika, 44(1/2):265–268, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333264.

Guttman:1958:NSS

[Gut58] Irwin Guttman. A note on a series solution of a problem in estimation. Biometrika, 45(3/4):565–567, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333204.

Geary:1947:CUM

[GW47] R. C. Geary and J. P. G. Worlledge. On the computation of universal moments of tests of statistical normality derived from samples drawn at random from a normal universe. Application to the calculation of the seventh moment of b_{2}. Biometrika, 34(1/2):98–110, January 1947. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332514. See corrigenda [GW50].

Geary:1950:CCU

[GW50] R. C. Geary and J. P. G. Worlledge. Corrigenda: “On the Computation of Universal Moments of Tests of Statistical Normality Derived from Samples Drawn at Random from a Normal Universe. Application to the Calculation of the Seventh Moment of b_{2}”. Biometrika, 37(1/2): 189, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332167. See

[GW47].

Haberman:1955:DKT

[Hab55] Sol Haberman. Distributions of Kendall’s Tau based on partially ordered systems. Biometrika, 42(3/4):417–424, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333388.

Hack:1958:EID

[Hac58] H. R. B. Hack. An empirical investigation into the distribution of the F -ratio in samples from two non-normal populations. Biometrika, 45(1/2):260–265, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333064.

Haight:1957:QB

[Hai57] Frank A. Haight. Queueing with balking. Biometrika, 44(3/4):360–369, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332868.

Haight:1958:TQP

[Hai58] Frank A. Haight. Two queues in parallel. Biometrika, 45(3/4):401–410, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333187.

Haight:1959:OTL

[Hai59] Frank A. Haight. Overflow at a traffic light. Biometrika, 46(3/4):420– 424, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333538.

Haldane:1955:PSS

[Hal55a] J. B. S. Haldane. A problem in the significance of small numbers. Biometrika, 42(1/2):266–267, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333445.

Haldane:1955:RCT

[Hal55b] J. B. S. Haldane. The rapid calculation of χ^{2} as a test of homogeneity from a 2 × n table. Biometrika, 42(3/4):519–520, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333398.

Haldane:1955:S

[Hal55c] J. B. S. Haldane. Substitutes for χ^{2}. Biometrika, 42(1/2):265–266, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333444.

Haldane:1957:KP

[Hal57] J. B. S. Haldane. Karl Pearson, 1857–1957. Biometrika, 44(3/4):303– 313, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332863.

Hammersley:1953:CRA

[Ham53] J. M. Hammersley. Capture-recapture analysis. Biometrika, 40 (3/4):265–278, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333345.

Hannan:1955:ETC

[Han55a] E. J. Hannan. An exact test for correlation between time series. Biometrika, 42(3/4):316–326, December 1955. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333381.

Hannan:1955:ETS

[Han55b] E. J. Hannan. Exact tests for serial correlation. Biometrika, 42(1/2): 133–142, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333430.

Hannan:1957:TSC

[Han57] E. J. Hannan. Testing for serial correlation in least squares regression. Biometrika, 44(1/2):57–66, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333240.

Hartley:1950:MRS

[Har50a] H. O. Hartley. The maximum F -ratio as a short-cut test for heterogeneity of variance. Biometrika, 37(3/4):308–312, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332383.

Hartley:1950:SFS

[Har50b] H. O. Hartley. A simplified form of Sheppard’s correction formulae. Biometrika, 37(1/2):145–148, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332157.

Hartley:1950:URA

[Har50c] H. O. Hartley. The use of range in analysis of variance. Biometrika, 37(3/4):271–280, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332380.

Hartley:1951:FPE

[Har51] H. O. Hartley. The fitting of polynomials to equidistant data with missing values. Biometrika, 38(3/4):410–413, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332586.

Hartley:1953:PID

[Har53] H. O. Hartley. Publications of U.S. Department of Commerce, National Bureau of Standards. Biometrika, 40(3/4):477–478, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333374.

Harley:1954:NPI

[Har54] B. I. Harley. A note on the probability integral of the correlation coefficient. Biometrika, 41(1/2):278–280, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333027.

Harley:1956:SPA

[Har56] B. I. Harley. Some properties of an angular transformation for the correlation coefficient. Biometrika, 43(1/2):219–224, June 1956. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333599.

Harley:1957:FPA

[Har57a] B. I. Harley. Further properties of an angular transformation of the correlation coefficient. Biometrika, 44(1/2):273–275, June 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333268.

Harley:1957:RBD

[Har57b] B. I. Harley. Relation between the distributions of noncentral t and of a transformed correlation coefficient. Biometrika, 44(1/2):219–224, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333255.

Harris:1957:MMP

[Har57c] A. J. Harris. A maximum-minimum problem related to statistical distributions in two dimensions. Biometrika, 44(3/4):384–398, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332871.

Hartley:1959:EIR

[Har59] H. O. Hartley. The efficiency of internal regression for the fitting of the exponential regression. Biometrika, 46(3/4):293–295, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333526.

Haskey:1954:GEM

[Has54] H. W. Haskey. A general expression for the mean in a simple stochastic epidemic. Biometrika, 41(1/2):272–275, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333025.

Haskey:1957:SCI

[Has57] H. W. Haskey. Stochastic cross-infection between two otherwise isolated groups. Biometrika, 44(1/2):193–204, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333252.

Healy:1953:BRB

[Hea53] M. J. R. Healy. Book review: The Statistics of Bio-Assay, by C. I. Bliss; Statistical Method in Biological Assay. by D. J. Finney; Probit Analysis. by D. J. Finney. Biometrika, 40(3/4):473, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333371.

Healy:1956:WPA

[Hea56] M. J. R. Healy. Weighted probits allowing for a non-zero response in the controls. Biometrika, 43(1/2):207–208, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333595.

Heine:1955:MTD

[Hei55] V. Heine. Models for two-dimensional stationary stochastic processes. Biometrika, 42(1/2):170–178, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333434.

Herdan:1958:MRB

[Her58a] G. Herdan. The mathematical relation between Greenberg’s index of linguistic diversity and Yule’s characteristic. Biometrika, 45(1/2):268– 270, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333067.

Herdan:1958:RBD

[Her58b] G. Herdan. The relation between the dictionary distribution and the occurrence distribution of word length and its importance for the study of quantitative linguistics. Biometrika, 45(1/2):222–228, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333058.

Hartley:1951:CIB

[HF51] H. O. Hartley and E. R. Fitch. A chart for the incomplete beta-function and the cumulative binomial distribution. Biometrika, 38(3/4):423– 426, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332588.

Halperin:1958:NMC

[HG58] Max Halperin and S. W. Greenhouse. Note on multiple comparisons for adjusted means in the analysis of covariance. Biometrika, 45(1/2): 256–259, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333063.

Hill:1954:DRC

[Hil54] I. D. Hill. The distribution of the regression coefficient in samples from a non-normal population. Biometrika, 41(3/4):548–552, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332735.

Hammersley:1954:ELS

[HM54] J. M. Hammersley and K. W. Morton. The estimation of location and scale parameters from grouped data. Biometrika, 41(3/4):296–301, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332710.

Heathcote:1959:RWC

[HM59] C. R. Heathcote and J. E. Moyal. The random walk [in continuous time] and its application to the theory of queues. Biometrika, 46(3/4):400–411, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333536.

Hodges:1955:GRO

[Hod55] J. L. Hodges, Jr. Galton’s rank-order test. Biometrika, 42(1/2):261– 262, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333442.

Hoel:1954:TMC

[Hoe54] Paul G. Hoel. A test for Markoff chains. Biometrika, 41(3/4):430–433, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332723.

Hooper:1958:SVC

[Hoo58] John W. Hooper. The sampling variance of correlation coefficients under assumptions of fixed and mixed variates. Biometrika, 45(3/4):471– 477, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333193.

Horsnell:1953:EUG

[Hor53] G. Horsnell. The effect of unequal group variances on the F -test for the homogeneity of group means. Biometrika, 40(1/2):128–136, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333104.

Hartley:1950:TPI

[HP50a] H. O. Hartley and E. S. Pearson. Table of the probability integral of the t-distribution. Biometrika, 37(1/2):168–172, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332160.

Hartley:1950:TIC

[HP50b] H. O. Hartley and E. S. Pearson. Tables of the χ^{2}-integral and of the cumulative Poisson distribution. Biometrika, 37(3/4):313–325, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332384.

Hartley:1951:MCD

[HP51] H. O. Hartley and E. S. Pearson. Moment constants for the distribution of range in normal samples. Biometrika, 38(3/4):463–464, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332591.

Hacker:1952:GSW

[HP52] H. P. Hacker and H. S. Pearson. The growth, survival, wandering and variation of the long-tailed field mouse, Apodemus Sylvaticus. III. Wandering power and distribution. Biometrika, 39(3/4):389–413, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334032.

Harley:1957:DRN

[HP57] B. I. Harley and E. S. Pearson. The distribution of range in normal samples with n = 200. Biometrika, 44(1/2):257–260, June 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333260.

Haldane:1956:SDM

[HS56] J. B. S. Haldane and Sheila Maynard Smith. The sampling distribution of a maximum-likelihood estimate. Biometrika, 43(1/2):96–103, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333582.

Huitson:1955:MAC

[Hui55] A. Huitson. A method of assigning confidence limits to linear combinations of variances. Biometrika, 42(3/4):471–479, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333392.

Huitson:1958:FCV

[Hui58] A. Huitson. Further critical values for the sum of two variances. Biometrika, 45(1/2):279–282, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333071.

Huttly:1956:FRC

[Hut56] N. A. Huttly. The fitting of regression curves with autocorrelated data. Biometrika, 43(3/4):468–474, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332929.

Huzurbazar:1955:EFS

[Huz55] V. S. Huzurbazar. Exact forms of some invariants for distributions admitting sufficient statistics. Biometrika, 42(3/4):533–537, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333405.

Hyrenius:1950:DSF

[Hyr50] Hannes Hyrenius. Distribution of ‘Student’-Fisher’s t in samples from compound normal functions. Biometrika, 37(3/4):429–442, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332392.

Hyrenius:1952:SBN

[Hyr52] Hannes Hyrenius. Sampling from bivariate non-normal universes by means of compound normal distributions. Biometrika, 39(3/4):238– 246, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334020.

Iyer:1954:PDA

[IK54] P. V. Krishna Iyer and M. N. Kapur. Probability distributions arising from points on a line. Biometrika, 41(3/4):553–554, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332736.

Irwin:1954:DAS

[Irw54] J. O. Irwin. A distribution arising in the study of infectious diseases. Biometrika, 41(1/2):266–268, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333023.

James:1951:CSG

[Jam51] G. S. James. The comparison of several groups of observations when the ratios of the population variances are unknown. Biometrika, 38(3/4):324–329, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332578.

James:1954:TLH

[Jam54] G. S. James. Tests of linear hypotheses in univariate and multivariate analysis when the ratios of the population variances are unknown. Biometrika, 41(1/2):19–43, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333003.

James:1955:CTV

[Jam55] G. S. James. Cumulants of a transformed variate. Biometrika, 42 (3/4):529–531, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333403.

James:1956:AWM

[Jam56] G. S. James. On the accuracy of weighted means and ratios. Biometrika, 43(3/4):304–321, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332909.

Jenkins:1954:ATS

[Jen54a] G. M. Jenkins. An angular transformation for the serial correlation coefficient. Biometrika, 41(1/2):261–265, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333022.

Jenkins:1954:THL

[Jen54b] G. M. Jenkins. Tests of hypotheses in the linear autoregressive model: 1. Null hypothesis distributions in the Yule scheme. Biometrika, 41(3/4):405–419, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332721.

Jenkins:1955:BRB

[Jen55] G. M. Jenkins. Book review: A Study in the Analysis of Stationary Time-Series, by Herman Wold; Peter Whittle. Biometrika, 42 (3/4):538–539, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333407.

Jenkins:1956:THL

[Jen56] G. M. Jenkins. Tests of hypotheses in the linear autoregressive model: II. Null distributions for higher order schemes: Non-null distributions. Biometrika, 43(1/2):186–199, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333591.

Jenkins:1958:BRB

[Jen58] G. M. Jenkins. Book review: Statistical Analysis of Stationary Time Series, by U. Grenander; M. Rosenblatt. Biometrika, 45(1/2):283–284, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333074.

Jenkins:1959:BRB

[Jen59] G. M. Jenkins. Book review: Analysis of Multiple Time-Series, by M. H. Quenouille. Biometrika, 46(1/2):267, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332837.

Johnson:1950:CE

[Joh50] N. L. Johnson. On the comparison of estimators. Biometrika, 37 (3/4):281–287, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332381.

Johnson:1951:BRB

[Joh51] N. L. Johnson. Book reviews: Experimental Designs by William G. Cochran; Gertrude M. Cox; Analysis and Design of Experiments by H. B. Mann. Biometrika, 38(1/2):260–261, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332334.

Johnson:1952:API

[Joh52a] N. L. Johnson. Approximations to the probability integral of the distribution of range. Biometrika, 39(3/4):417–419, December 1952. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334034.

Johnson:1952:CAV

[Joh52b] N. L. Johnson. Comparison of analysis of variance power functions in the parametric and random models. Biometrika, 39(3/4):427–429, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334038.

Johnson:1954:BRBb

[Joh54a] N. L. Johnson. Book review: Design for Decision by Irwin D. J. Bross. Biometrika, 41(1/2):283, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333032.

Johnson:1954:BRBa

[Joh54b] N. L. Johnson. Book review: Statistical Methods in Electrical Engineering by D. A. Bell. Biometrika, 41(1/2):282, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333031.

Johnson:1954:BRBc

[Joh54c] N. L. Johnson. Book review: Tables for the Design of Factorial Experiments, by T. Kitagawa; M. Mitome. Biometrika, 41(3/4):567, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332744.

Johnson:1955:BRBa

[Joh55a] N. L. Johnson. Book review: Life and Other Contingencies, by P. F. Hooker; L. H. Longley-Cook. Biometrika, 42(1/2):274, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333453.

Johnson:1955:BRBb

[Joh55b] N. L. Johnson. Book review: Population Statistics and their Compilation, by Hugh H. Wolfenden. Biometrika, 42(1/2):274–275, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333454.

Johnson:1955:BRBc

[Joh55c] N. L. Johnson. Book review: Tablitsi Dlya Vichisleniya Nepolnoi Γ-Funktsii Veroyatnosti χ^{2} (Tables for the Calculation of the Incomplete Γ-Function and the χ^{2}-Probability Function). by E. E. Slutskii; A. N. Kolmogorov. Biometrika, 42(3/4):543, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333415.

Johnson:1956:BRB

[Joh56] N. L. Johnson. Book review: Introduction to Demography by Mortimer Spiegelman. Biometrika, 43(1/2):237–238, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333608.

Johnson:1957:NMD

[Joh57a] N. L. Johnson. A note on the mean deviation of the binomial distribution. Biometrika, 44(3/4):532–533, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332887. See corrigenda [Joh58d].

Johnson:1957:OSQ

[Joh57b] N. L. Johnson. Optimal sampling for quota fulfilment. Biometrika, 44(3/4):518–523, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332882.

Johnson:1957:SDC

[Joh57c] N. L. Johnson. Sequentially determined confidence intervals. Biometrika, 44(1/2):279–281, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333271.

Johnson:1957:URT

[Joh57d] N. L. Johnson. Uniqueness of a result in the theory of accident proneness. Biometrika, 44(3/4):530–531, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332886.

Johnson:1958:BRBa

[Joh58a] N. L. Johnson. Book review: Experimental Designs, by W. G. Cochran; G. M. Cox. Biometrika, 45(1/2):287, June 1958. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333077.

Johnson:1958:BRBb

[Joh58b] N. L. Johnson. Book review: Life and Other Contingencies, by P. F. Hooker; L. H. Longley-Cook. Biometrika, 45(1/2):289, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333083.

Johnson:1958:BRBc

[Joh58c] N. L. Johnson. Book review: Statisticke Tabulky. (Statistical Tables), by Jaroslav Janko. Biometrika, 45(3/4):597, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333226.

Johnson:1958:CNM

[Joh58d] N. L. Johnson. Corrigenda: “A Note on the Mean Deviation of the Binomial Distribution”. Biometrika, 45(3/4):587, December 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333211. See [Joh57a].

Johnson:1958:MDS

[Joh58e] N. L. Johnson. The mean deviation, with special reference to samples from a Pearson Type III population. Biometrika, 45(3/4):478–483, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333194.

Johnson:1958:TCR

[Joh58f] N. L. Johnson. Theoretical considerations regarding H. R. B. Hack’s system of randomization for cross-classifications. Biometrika, 45(1/2): 265–266, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333065.

Johnson:1959:BRB

[Joh59a] N. L. Johnson. Book review: Experimental Designs in Industry, by V. Clew. Biometrika, 46(1/2):266, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332836.

Johnson:1959:ECB

[Joh59b] N. L. Johnson. On an extension of the connexion between Poisson and χ^{2} distributions. Biometrika, 46(3/4):352–363, December 1959.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333532.

Jonckheere:1954:DFS

[Jon54] A. R. Jonckheere. A distribution-free k-sample test against ordered alternatives. Biometrika, 41(1/2):133–145, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333011.

Jonckheere:1956:BRB

[Jon56] A. R. Jonckheere. Book review: Stochastic Models for Learning by R. R. Bush; F. Mosteller. Biometrika, 43(1/2):237, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333607.

Jowett:1955:SPL

[Jow55] G. H. Jowett. Sampling properties of local statistics in stationary stochastic series. Biometrika, 42(1/2):160–169, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333433.

Jowett:1957:SAU

[Jow57] G. H. Jowett. Statistical analysis using local properties of smoothly heteromorphic stochastic series. Biometrika, 44(3/4):454–463, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332876.

Jowett:1959:JA

[JW59] G. H. Jowett and Wendy M. Wright. Jump analysis. Biometrika, 46(3/4):386–399, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333535.

Kabe:1958:SAM

[Kab58] D. G. Kabe. Some applications of Meijer–G functions to distribution problems in statistics. Biometrika, 45(3/4):578–580, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333208.

Kalmus:1957:BRB

[Kal57] H. Kalmus. Book review: Theoretical Genetics by R. B. Goldschmidt. Biometrika, 44(1/2):298, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333281.

Kamat:1953:IAM

[Kam53a] A. R. Kamat. Incomplete and absolute moments of the multivariate normal distribution with some applications. Biometrika, 40(1/2):20– 34, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333092.

Kamat:1953:MSD

[Kam53b] A. R. Kamat. On the mean successive difference and its ratio to the root mean square. Biometrika, 40(1/2):116–127, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333103. See corrigenda [Kam58].

Kamat:1953:TMG

[Kam53c] A. R. Kamat. The third moment of Gini’s mean difference. Biometrika, 40(3/4):451–452, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333362.

Kamat:1954:DTT

[Kam54a] A. R. Kamat. Distribution theory of two estimates for standard deviation based on second variate differences. Biometrika, 41(1/2):1– 11, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333001.

Kamat:1954:MMD

[Kam54b] A. R. Kamat. Moments of the mean deviation. Biometrika, 41 (3/4):541–542, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332732.

Kamat:1956:TSD

[Kam56] A. R. Kamat. A two-sample distribution-free test. Biometrika, 43(3/4):377–387. Addendum 386–387, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332916. Addendum by D. E. Barton.

Kamat:1958:CMS

[Kam58] A. R. Kamat. Corrigenda: “On the Mean Successive Difference and its Ratio to the Root Mean Square”. Biometrika, 45(3/4):587, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333212. See

[Kam53b].

Kaplan:1952:TNS

[Kap52] E. L. Kaplan. Tensor notation and the sampling cumulants of kstatistics. Biometrika, 39(3/4):319–323, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2334028.

Kathirgamatamby:1953:NPI

[Kat53] N. Kathirgamatamby. Note on the Poisson index of dispersion. Biometrika, 40(1/2):225–228, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333115.

Kendall:1951:RSF

[Ken51] M. G. Kendall. Regression, structure and functional relationship. Part I. Biometrika, 38(1/2):11–25, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332313.

Kendall:1952:MSS

[Ken52a] M. G. Kendall. Moment-statistics in samples from a finite population. Biometrika, 39(1/2):14–16, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332459.

Kendall:1952:RSF

[Ken52b] M. G. Kendall. Regression, structure and functional relationship. Part II. Biometrika, 39(1/2):96–108, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332467.

Kendall:1954:NBE

[Ken54a] M. G. Kendall. Note on bias in the estimation of autocorrelation. Biometrika, 41(3/4):403–404, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332720.

Kendall:1954:TPS

[Ken54b] M. G. Kendall. Two problems in sets of measurements. Biometrika, 41(3/4):560–564, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332740.

Kendall:1955:BRB

[Ken55] M. G. Kendall. Book review: Biomathematics, by Cedric A. B. Smith. Biometrika, 42(1/2):271, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333448.

Kendall:1956:SHP

[Ken56] M. G. Kendall. Studies in the history of probability and statistics: II. The beginnings of a probability calculus. Biometrika, 43(1/2):1–14 (2 plates), June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333573.

Kendall:1957:MLD

[Ken57a] M. G. Kendall. The moments of the Leipnik distribution. Biometrika, 44(1/2):270–272, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333266.

Kendall:1957:SHP

[Ken57b] M. G. Kendall. Studies in the history of probability and statistics. V. A note on playing cards. Biometrika, 44(1/2):260–262, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333261.

Kendall:1959:BRB

[Ken59] M. G. Kendall. Book review: The Logic of Scientific Discovery, by Karl R. Popper. Biometrika, 46(1/2):265–266, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332835.

Kerrich:1958:NDP

[Ker58] J. E. Kerrich. Note on a discontinuous probability density. Biometrika, 45(1/2):270–273, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333068.

Kestelman:1957:BRB

[Kes57] H. Kestelman. Book review: Trigonometrical Series by A. Zygmund. Biometrika, 44(1/2):299, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333285.

Kestelman:1958:BRB

[Kes58] H. Kestelman. Book review: Vector Spaces and Matrices, by R. M. Thrall; L. Tornheim. Biometrika, 45(1/2):287–288, June 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333080.

Kermack:1950:OCA

[KH50] K. A. Kermack and J. B. S. Haldane. Organic correlation and allometry. Biometrika, 37(1/2):30–41, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332144.

Khatri:1959:CPP

[Kha59] C. G. Khatri. On certain properties of power-series distributions. Biometrika, 46(3/4):486–490, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333550.

Kimball:1959:ALI

[KL59] A. W. Kimball and E. Leach. Approximate linearization of the incomplete β-function. Biometrika, 46(1/2):214–218, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332823.

Kullback:1957:AMR

[KR57] S. Kullback and H. M. Rosenblatt. On the analysis of multiple regression in k categories. Biometrika, 44(1/2):67–83, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333241.

Kupperman:1952:EGC

[Kup52] Morton Kupperman. On exact grouping corrections to moments and cumulants. Biometrika, 39(3/4):429–434, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2334039.

Laha:1953:EGT

[Lah53] R. G. Laha. On an extension of Geary’s theorem. Biometrika, 40(1/2): 228–229, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333116.

Lancaster:1949:CPA

[Lan49a] H. O. Lancaster. The combination of probabilities arising from data in discrete distributions. Biometrika, 36(3/4):370–382, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332674. See corri-

genda [Lan50a].

Lancaster:1949:DPC

[Lan49b] H. O. Lancaster. The derivation and partition of χ^{2} in certain discrete distributions. Biometrika, 36(1/2):117–129, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332535. See corrigenda [Lan50b].

Lancaster:1950:CCP

[Lan50a] H. O. Lancaster. Corrigenda: “The Combination of Probabilities Arising from Data in Discrete Distributions”. Biometrika, 37(3/4):452, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332400. See

[Lan49a].

Lancaster:1950:CDP

[Lan50b] H. O. Lancaster. Corrigenda: “The Derivation and Partition of χ^{2} in Certain Discrete Distributions”. Biometrika, 37(3/4):452, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332399. See [Lan49b].

Lancaster:1950:EPA

[Lan50c] H. O. Lancaster. The exact partition of χ^{2} and its application to the problem of the pooling of small expectations. Biometrika, 37(3/4):267– 270, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332379.

Lancaster:1952:SCC

[Lan52] H. O. Lancaster. Statistical control of counting experiments. Biometrika, 39(3/4):419–422, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334035.

Lancaster:1957:SPB

[Lan57] H. O. Lancaster. Some properties of the bivariate normal distribution considered in the form of a contingency table. Biometrika, 44(1/2):289– 292, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333274.

Latscha:1953:TSC

[Lat53] R. Latscha. Tests of significance in a 2×2 contingency table: Extension of Finney’s table. Biometrika, 40(1/2):74–86, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333099.

Lawley:1956:GMA

[Law56a] D. N. Lawley. A general method for approximating to the distribution of likelihood ratio criteria. Biometrika, 43(3/4):295–303, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332908.

Lawley:1956:TSL

[Law56b] D. N. Lawley. Tests of significance for the latent roots of covariance and correlation matrices. Biometrika, 43(1/2):128–136, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333586.

Lawley:1959:TSC

[Law59] D. N. Lawley. Tests of significance in canonical analysis. Biometrika, 46(1/2):59–66, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332808.

Leslie:1951:EPP

[LC51] P. H. Leslie and Dennis Chitty. The estimation of population parameters from data obtained by means of the capture-recapture method: I. The maximum likelihood equations for estimating the death-rate. Biometrika, 38(3/4):269–292, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332574.

Leslie:1953:EPP

[LCC53] P. H. Leslie, Dennis Chitty, and Helen Chitty. The estimation of population parameters from data obtained by means of the capture– recapture method: III. An example of the practical applications of the method. Biometrika, 40(1/2):137–169, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333105.

Leigh-Dugmore:1953:RME

[LD53] C. H. Leigh-Dugmore. A rapid method for estimating the correlation coefficient from the range of the deviations about the reduced major axis. Biometrika, 40(1/2):218–219, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333112.

Leipnik:1958:NCF

[Lei58a] Roy Leipnik. Note on the characteristic function of a serial-correlation distribution. Biometrika, 45(3/4):559–562, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333202.

Leipnik:1958:MGF

[Lei58b] Roy B. Leipnik. Moment generating functions of quadratic forms in serially correlated normal variables. Biometrika, 45(1/2):198–210, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333056.

Leslie:1952:EPP

[Les52] P. H. Leslie. The estimation of population parameters from data obtained by means of the capture-recapture method. II. The estimation of total numbers. Biometrika, 39(3/4):363–388, December 1952. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334031.

Leslie:1955:SMC

[Les55] P. H. Leslie. A simple method of calculating the exact probability in 2 × 2 contingency tables with small marginal totals. Biometrika, 42(3/4):522–523, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333400.

Leslie:1957:ADS

[Les57] P. H. Leslie. An analysis of the data for some experiments carried out by Gause with populations of the protozoa, Paramecium Aurelia and Paramecium Caudatum. Biometrika, 44(3/4):314–327, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332864.

Leslie:1958:SMS

[Les58] P. H. Leslie. A stochastic model for studying the properties of certain biological systems by numerical methods. Biometrika, 45(1/2):16–31, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333042.

Leslie:1959:DPJ

[Les59] D. C. M. Leslie. Determination of parameters in the Johnson system of probability distributions. Biometrika, 46(1/2):229–231, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332826.

Levine:1959:MMS

[Lev59] Jack Levine. Monomial-monomial symmetric function tables. Biometrika, 46(1/2):205–213, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332822.

Lewis:1953:PD

[Lew53] T. Lewis. 99.9 and 0.1% points of the χ^{2} distribution. Biometrika, 40(3/4):421–426, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333358.

Lewis:1957:BRB

[Lew57] T. Lewis. Book review: Methods in Numerical Analysis by K. L. Nielsen. Biometrika, 44(1/2):299–300, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333287.

Lewis:1958:BRB

[Lew58] T. Lewis. Book review: Numerical Analysis, by John H. Curtiss. Biometrika, 45(3/4):599, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333231.

Leslie:1958:PSM

[LG58] P. H. Leslie and J. C. Gower. The properties of a stochastic model for two competing species. Biometrika, 45(3/4):316–330, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333181. See corri-

genda [LG59].

Leslie:1959:CPS

[LG59] P. H. Leslie and J. C. Gower. Corrigenda: “The Properties of a Stochastic Model for Two Competing Species”. Biometrika, 46(1/2):

279, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332859. See

[LG58].

Lieblein:1954:TEP

[Lie54] Julius Lieblein. Two early papers on the relation between extreme values and tensile strength. Biometrika, 41(3/4):559–560, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332739.

Lindley:1953:EFR

[Lin53] D. V. Lindley. Estimation of a functional relationship. Biometrika, 40(1/2):47–49, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333095.

Lindley:1957:BSS

[Lin57a] D. V. Lindley. Binomial sampling schemes and the concept of information. Biometrika, 44(1/2):179–186, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333250.

Lindley:1957:SP

[Lin57b] D. V. Lindley. A statistical paradox. Biometrika, 44(1/2):187–192, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333251.

Lloyd:1952:LSE

[Llo52] E. H. Lloyd. Least-squares estimation of location and scale parameters using order statistics. Biometrika, 39(1/2):88–95, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332466.

Lloyd:1959:NPE

[Llo59] D. E. Lloyd. Note on a problem of estimation. Biometrika, 46(1/2): 231–235, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332827.

Long:1951:EPW

[Lon51a] W. M. Long. Estimation problems when a simple type of heterogeneity is present in the sample. Biometrika, 38(1/2):90–101, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332319.

Long:1951:QCS

[Lon51b] W. M. Long. Quality control systems based on inaccurately measured variables. Biometrika, 38(3/4):472–475, December 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332595.

Lord:1947:URP

[Lor47] E. Lord. The use of range in place of standard deviation in the t-test. Biometrika, 34(1/2):41–67, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2332512. See corrigenda [Lor52].

Lord:1950:PMT

[Lor50] E. Lord. Power of the modified t-test ( u-test) based on range. Biometrika, 37(1/2):64–77, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332148.

Lord:1952:CUR

[Lor52] E. Lord. Corrigenda: “The Use of Range in Place of Standard Deviation in the t-Test”. Biometrika, 39(3/4):442, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2334045. See [Lor47].

Lord:1954:UHTa

[Lor54a] R. D. Lord. The use of the Hankel transform in statistics. I. General theory and examples. Biometrika, 41(1/2):44–55, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333004.

Lord:1954:UHTb

[Lor54b] R. D. Lord. The use of the Hankel transform in statistics. II. Methods of computation. Biometrika, 41(3/4):344–350, December 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332715.

Lord:1958:SHP

[Lor58] R. D. Lord. Studies in the history of probability and statistics.: VIII. De Morgan and the statistical study of literary style. Biometrika, 45(1/2):282, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333072.

Mack:1953:EOB

[Mac53] C. Mack. The effect of overlapping in bacterial counts of incubated colonies. Biometrika, 40(1/2):220–222, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333113.

Mackenzie:1958:SPS

[Mac58] J. K. Mackenzie. Second paper on statistics associated with the random disorientation of cubes. Biometrika, 45(1/2):229–240, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333059.

Mallows:1956:NMP

[Mal56] C. L. Mallows. Note on the moment-problem for unimodal distributions when one or both terminals are known. Biometrika, 43(1/2):224– 227, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333600.

Mallows:1957:NNR

[Mal57] C. L. Mallows. Non-null ranking models. I. Biometrika, 44(1/2):114– 130, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333244.

Mallows:1959:BRBc

[Mal59a] C. L. Mallows. Book review: Elementary Matrix Algebra, by Franz E. Hohn. Biometrika, 46(3/4):497, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333562.

Mallows:1959:BRBa

[Mal59b] C. L. Mallows. Book review: Fundamentals of Statistical Reasoning, by M. H. Quenouille. Biometrika, 46(1/2):268, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332838.

Mallows:1959:BRBb

[Mal59c] C. L. Mallows. Book review: Planning of Experiments, by D. R. Cox. Biometrika, 46(3/4):492–493, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333552.

Mallows:1959:BRBd

[Mal59d] C. L. Mallows. Book review: Surveys in Applied Mathematics. Vol. 1: Elasticity and Plasticity, by J. N. Goodier; P. G. Hodge,; Surveys in Applied Mathematics. Vol. 2: Dynamics and Nonlinear by E. Leimanis; N. Minorsky; Surveys in Applied Mathematics. Vol. 3: Mathematical Aspects of Subsonic and Transonic Gas Dynamics. by Lipman Bers; Surveys in Applied Mathematics. Vol. 4: Some Aspects of Analysis and Probability. by Irving Kaplansky; Marshall Hall; Edwin Hewitt; Robert Fortet; Surveys in Applied Mathematics. Vol. 5: Numerical Analysis and Partial Differential Equations. by George E. Forsythe; Paul C. Rosenbloom. Biometrika, 46(3/4):499–500, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333566.

Mallows:1959:PIT

[Mal59e] C. L. Mallows. On the probability integral transformation. Biometrika, 46(3/4):481–483, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333548.

Maritz:1952:NCF

[Mar52a] J. S. Maritz. Note on a certain family of discrete distributions. Biometrika, 39(1/2):196–198, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332481.

Marriott:1952:TSC

[Mar52b] F. H. C. Marriott. Tests of significance in canonical analysis. Biometrika, 39(1/2):58–64, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332463.

Maurice:1958:RMT

[Mau58a] Rita Maurice. Ranking means of two normal populations with unknown variances. Biometrika, 45(1/2):250–252, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333061.

Maurice:1958:SPL

[Mau58b] Rita J. Maurice. Selection of the population with the largest mean when comparisons can be made only in pairs. Biometrika, 45(3/4):581– 586, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333210.

May:1952:ECT

[May52] Joyce M. May. Extended and corrected tables of the upper percentage points of the ‘Studentized’ range. Biometrika, 39(1/2):192–193, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332479.

Mayne:1954:SFR

[May54] Alan J. Mayne. Some further results in the theory of pedestrians and road traffic. Biometrika, 41(3/4):375–389, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332718. See corrigenda [May58].

Mayne:1958:CSF

[May58] Alan J. Mayne. Corrigenda: “Some Further Results in the Theory of Pedestrians and Road Traffic”. Biometrika, 45(1/2):291, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333085. See [May54].

McFadden:1956:ASQ

[McF56] J. A. McFadden. An approximation for the symmetric, quadrivariate normal integral. Biometrika, 43(1/2):206–207, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333594.

McKinlay:1951:MG

[McK51] P. L. McKinlay. Major Greenwood: 1880–1949. Biometrika, 38(1/2): 1–3, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332311.

McMullen:1955:BRB

[McM55] L. McMullen. Book review: Design and Analysis of Industrial Experiments, by O. L. Davies. Biometrika, 42(1/2):272, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333450.

McMullen:1957:BRB

[McM57] L. McMullen. Book review: Statistical Methods in Research and Production with Special Reference to the Chemical Industry, by O. L. Davies. Biometrika, 44(3/4):537, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332894.

Medhi:1956:NRE

[Med56] J. Medhi. A note on the risks of error involved in the sequential ratio test. Biometrika, 43(1/2):231–234, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333602.

Mendenhall:1958:BLT

[Men58] William Mendenhall. A bibliography on life testing and related topics. Biometrika, 45(3/4):521–543, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333198.

Merrington:1955:PN

[Mer55] M. Merrington. Prefatory note. Biometrika, 42(??):??, 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331622.

Mendenhall:1958:EPM

[MH58] William Mendenhall and R. J. Hader. Estimation of parameters of mixed exponentially distributed failure time distributions from censored life test data. Biometrika, 45(3/4):504–520, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333197.

Miles:1959:CAB

[Mil59] R. E. Miles. The complete amalgamation into blocks, by weighted means, of a finite set of real numbers. Biometrika, 46(3/4):317–327, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333529.

Miller:1974:JR

[Mil74] Rupert G. Miller. The jackknife — a review. Biometrika, 61(1):1– 15, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334280. See

[Que49, Que56, Efr79].

Mitton:1959:DFE

[MM59] R. G. Mitton and F. R. Morgan. The design of factorial experiments: A survey of some schemes requiring not more than 256 treatment combinations. Biometrika, 46(1/2):251–259, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332832.

Mohan:1955:GRP

[Moh55] C. Mohan. The Gambler’s ruin problem with correlation. Biometrika, 42(3/4):486–493, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333394.

Moore:1952:EPP

[Moo52] P. G. Moore. The estimation of the Poisson parameter from a truncated distribution. Biometrika, 39(3/4):247–251, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2334021.

Moore:1953:BRB

[Moo53a] P. G. Moore. Book review: Statistical Theory with Engineering Applications, by A. Hald; Statistical Tables and Formulas, by A. Hald. Biometrika, 40(3/4):473–474, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333372.

Moore:1953:STR

[Moo53b] P. G. Moore. A sequential test for randomness. Biometrika, 40(1/2): 111–115, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333102.

Moore:1955:BRBc

[Moo55a] P. G. Moore. Book review: A Million Random Digits with 100,000 Normal Deviates, by The Rand Corporation. Biometrika, 42(3/4):543, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333414.

Moore:1955:BRBa

[Moo55b] P. G. Moore. Book review: Introduction to Mathematical Statistics, by Paul G. Hoel. Biometrika, 42(1/2):271–272, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333449.

Moore:1955:BRBb

[Moo55c] P. G. Moore. Book review: Table of Binomial Coefficients by Royal Society Mathematical Tables. Biometrika, 42(1/2):275, June 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333455.

Moore:1956:BRBa

[Moo56a] P. G. Moore. Book review: Mortality and Other Investigations by H. W. Haycocks; W. Perks. Biometrika, 43(1/2):238, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333609.

Moore:1956:BRBb

[Moo56b] P. G. Moore. Book review: Numerical Analysis by Z. Kopal. Biometrika, 43(1/2):239, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333611.

Moore:1956:EMC

[Moo56c] P. G. Moore. The estimation of the mean of a censored normal distribution by ordered variables. Biometrika, 43(3/4):482–485, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332933.

Moore:1957:BRB

[Moo57a] P. G. Moore. Book review: Statistical Methods in Quality Control, by D. J. Cowden. Biometrika, 44(3/4):537–538, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332895.

Moore:1957:TST

[Moo57b] P. G. Moore. The two-sample t-test based on range. Biometrika, 44(3/4):482–489, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332879.

Moore:1958:BRBb

[Moo58a] P. G. Moore. Book review: Digital Computer Programming, by D. D. McCracken. Biometrika, 45(1/2):288, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333081.

Moore:1958:BRBa

[Moo58b] P. G. Moore. Book review: Regression Analysis of Production Costs and Factory Operation, by P. Lyle; L. H. C. Tippett. Biometrika, 45(1/2):286, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333076.

Moore:1958:BRBc

[Moo58c] P. G. Moore. Book review: The Mathematical Theory of Epidemics, by N. T. J. Bailey. Biometrika, 45(3/4):589, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333214.

Moore:1958:SPR

[Moo58d] P. G. Moore. Some properties of runs in quality control procedures. Biometrika, 45(1/2):89–95, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333048. See corrigenda [Moo59].

Moore:1959:CSP

[Moo59] P. G. Moore. Corrigenda: “Some Properties of Runs in Quality Control Procedures”. Biometrika, 46(1/2):279, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332857. See [Moo58d].

Moran:1950:NCS

[Mor50a] P. A. P. Moran. Notes on continuous stochastic phenomena. Biometrika, 37(1/2):17–23, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332142.

Moran:1950:TSI

[Mor50b] P. A. P. Moran. A test for the serial independence of residuals. Biometrika, 37(1/2):178–181, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332162.

Moran:1951:BRB

[Mor51a] P. A. P. Moran. Book review: Probability Theory and its Applications. Volume I by W. Feller. Biometrika, 38(1/2):259–260, June 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332333.

Moran:1951:MTA

[Mor51b] P. A. P. Moran. A mathematical theory of animal trapping. Biometrika, 38(3/4):307–311, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332576.

Moran:1951:PMR

[Mor51c] P. A. P. Moran. Partial and multiple rank correlation. Biometrika, 38(1/2):26–32, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332314.

Moran:1952:EDR

[Mor52] P. A. P. Moran. The estimation of death-rates from capture-mark-recapture sampling. Biometrika, 39(1/2):181–188, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332476.

Morton:1956:BRB

[Mor56] G. Morton. Book review: Theory of Games and Statistical Decisions, by D. Blackwell; M. A. Girshick. Biometrika, 43(3/4):490, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332937.

Morton:1957:BRB

[Mor57] K. W. Morton. Book review: Symposium on Monte Carlo Methods by H. A. Meyer. Biometrika, 44(1/2):295–296, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333277.

Moser:1953:BRB

[Mos53] C. A. Moser. Book review: The Design and Analysis of Experiments, by O. Kempthorne. Biometrika, 40(3/4):470–471, December 1953. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333369.

Moser:1959:BRB

[Mos59] C. A. Moser. Book review: Probleme der Statistischen Methodenlehre in den Sozialwissenschaften, by O. Anderson. Biometrika, 46(1/2): 272, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332847.

Marriott:1954:BEA

[MP54] F. H. C. Marriott and J. A. Pope. Bias in the estimation of autocorrelations. Biometrika, 41(3/4):390–402, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332719.

Merrington:1958:ADN

[MP58] Maxine Merrington and E. S. Pearson. An approximation to the distribution of non-central t. Biometrika, 45(3/4):484–491, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333195.

Maguire:1952:TIB

[MPW52] B. A. Maguire, E. S. Pearson, and A. H. A. Wynn. The time intervals between industrial accidents. Biometrika, 39(1/2):168–180, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332475.

Maguire:1953:FNA

[MPW53] B. A. Maguire, E. S. Pearson, and A. H. A. Wynn. Further notes on the analysis of accident data. Biometrika, 40(1/2):213–216, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333110.

Mercer:1959:RWW

[MS59] A. Mercer and C. S. Smith. A random walk in which the steps occur randomly in time. Biometrika, 46(1/2):30–35, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332805.

Mackenzie:1957:SSA

[MT57] J. K. Mackenzie and M. J. Thomson. Some statistics associated with the random disorientation of cubes. Biometrika, 44(1/2):205–210, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333253.

Mulholland:1951:DWH

[Mul51] H. P. Mulholland. On distributions for which the Hartley-Khamis solution of the moment-problem is exact. Biometrika, 38(1/2):74–89, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332318.

Murteira:1951:NVD

[Mur51] Bento Murteira. Note on the variate differences of auto-regressive series. Biometrika, 38(3/4):479–480, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332597.

Nagler:1950:BUQ

[Nag50] H. Nagler. On the best unbiased quadratic estimate of the variance. Biometrika, 37(3/4):444–445, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332394.

Nair:1950:ECL

[Nai50] K. R. Nair. Efficiencies of certain linear systematic statistics for estimating dispersion from normal samples. Biometrika, 37(1/2):182–183, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332163.

Nair:1952:TPP

[Nai52] K. R. Nair. Tables of percentage points of the ‘Studentized’ extreme deviate from the sample mean. Biometrika, 39(1/2):189–191, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332478. See corri-

genda [Ano53d].

Nass:1959:TSE

[Nas59] C. A. G. Nass. The χ^{2} test for small expectations in contingency tables, with special reference to accidents and absenteeism. Biometrika, 46(3/4):365–385, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333534.

Nelder:1951:NSI

[Nel51] J. A. Nelder. A note on the statistical independence of quadratic forms in the analysis of variance. Biometrika, 38(3/4):482–483, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332599. See [Ait50].

Nelder:1954:INC

[Nel54] J. A. Nelder. The interpretation of negative components of variance. Biometrika, 41(3/4):544–548, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332734.

Olkin:1953:NJC

[Olk53] Ingram Olkin. Note on ‘The Jacobians of certain matrix transformations useful in multivariate analysis’. Biometrika, 40(1/2):43–46, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333094.

Olkin:1958:MRE

[Olk58] Ingram Olkin. Multivariate ratio estimation for finite populations. Biometrika, 45(1/2):154–165, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333053.

Owen:1958:BRB

[Owe58] A. R. G. Owen. Book review: Variation and Heredity, by H. Kalmus. Biometrika, 45(3/4):589–590, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333215.

Owen:1959:BRB

[Owe59] A. R. G. Owen. Book review: Genetics and the Improvement of Tropical Crops, by Joseph Hutchinson. Biometrika, 46(1/2):273, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332850.

Pachares:1959:TUP

[Pac59] J. Pachares. Table of the upper 10% points of the Studentized range. Biometrika, 46(3/4):461–466, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2333543. Accompanied by revised tables of 5% and 1% points.

Page:1954:CIS

[Pag54] E. S. Page. Continuous inspection schemes. Biometrika, 41(1/2):100– 115, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333009.

Page:1955:CCW

[Pag55a] E. S. Page. Control charts with warning lines. Biometrika, 42(1/2): 243–257, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333440.

Page:1955:TCP

[Pag55b] E. S. Page. A test for a change in a parameter occurring at an unknown point. Biometrika, 42(3/4):523–527, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333401.

Page:1957:PWC

[Pag57] E. S. Page. On problems in which a change in a parameter occurs at an unknown point. Biometrika, 44(1/2):248–252, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333258.

Palmer:1957:DIB

[Pal57] D. S. Palmer. The distribution of intervals between successive maxima in a series of random numbers. Biometrika, 44(3/4):524–526, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332883.

Patnaik:1948:PFT

[Pat48] P. B. Patnaik. The power function of the test for the difference between two proportions in a 2 × 2 table. Biometrika, 35(1/2):157–175, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332636. See cor-

rigenda [Pat59].

Patnaik:1950:UMR

[Pat50] P. B. Patnaik. The use of mean range as an estimator of variance in statistical tests. Biometrika, 37(1/2):78–87, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332149.

Patterson:1952:CBD

[Pat52] H. D. Patterson. The construction of balanced designs for experiments involving sequences of treatments. Biometrika, 39(1/2):32–48, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332461.

Patankar:1954:GFF

[Pat54] V. N. Patankar. The goodness of fit of frequency distributions obtained from stochastic processes. Biometrika, 41(3/4):450–462, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332725.

Patil:1957:CAP

[Pat57] V. T. Patil. The consistency and adequacy of the Poisson-Markoff model for density fluctuations. Biometrika, 44(1/2):43–56, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333239.

Patterson:1958:UAF

[Pat58] H. D. Patterson. The use of autoregression in fitting an exponential curve. Biometrika, 45(3/4):389–400, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333186.

Patnaik:1959:CPF

[Pat59] P. B. Patnaik. Corrigenda: “The Power Function of the Test for the Difference Between Two Proportions in a 2 × 2 Table”. Biometrika, 46(3/4):502, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333568. See [Pat48].

Pillai:1959:DLS

[PB59] K. C. Sreedharan Pillai and Celia G. Bantegui. On the distribution of the largest of six roots of a matrix in multivariate analysis. Biometrika, 46(1/2):237–240, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332829.

Pearson:1950:QRC

[Pea50a] E. S. Pearson. On questions raised by the combination of tests based on discontinuous distributions. Biometrika, 37(3/4):383–398, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332389. See corri-

genda [Pea51c].

Pearson:1950:SNU

[Pea50b] E. S. Pearson. Some notes on the use of range. Biometrika, 37(1/2): 88–92, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332150.

Pearson:1951:BRBa

[Pea51a] E. S. Pearson. Book review: Contributions to Mathematical Statistics by R. A. Fisher. Biometrika, 38(1/2):257–259, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332332.

Pearson:1951:BRBb

[Pea51b] E. S. Pearson. Book review: First Course in Probability and Statistics by J. Neyman. Biometrika, 38(1/2):261–262, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332335.

Pearson:1951:CQR

[Pea51c] E. S. Pearson. Corrigenda: “On Questions Raised by the Combination of Tests Based on Discontinuous Distributions”. Biometrika, 38(1/2): 265, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332338. See

[Pea50a].

Pearson:1952:BRB

[Pea52a] E. S. Pearson. Book review: Applied Statistics, by Leonard H. C. Tippett. Biometrika, 39(3/4):440–441, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2334043.

Pearson:1952:CTA

[Pea52b] E. S. Pearson. Comparison of two approximations to the distribution of the range in small samples from normal populations. Biometrika, 39(1/2):130–136, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332471.

Pearson:1954:PID

[Pea54] E. S. Pearson. Publications of U.S. Department of Commerce, National Bureau of Standards. Biometrika, 41(1/2):285–286, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333038.

Pearce:1956:BRB

[Pea56] S. C. Pearce. Book review: Experimental Design and its Statistical Basis, by D. J. Finney; Experimental Design. Theory and Application. by W. T. Federer. Biometrika, 43(3/4):491–492, December 1956. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332939.

Pearce:1957:EOB

[Pea57a] S. C. Pearce. Experimenting with organisms as blocks. Biometrika, 44(1/2):141–149, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333246.

Pearson:1957:JW

[Pea57b] E. S. Pearson. John Wishart, 1898–1956. Biometrika, 44(1/2):1–8, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333236.

Pearson:1959:BRB

[Pea59a] E. S. Pearson. Book review: The Advanced Theory of Statistics. Volume I, Distribution Theory, by Maurice G. Kendall; Alan Stuart. Biometrika, 46(3/4):491–492, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333551.

Pearson:1959:NAD

[Pea59b] E. S. Pearson. Note on an approximation to the distribution of non-central χ^{2}. Biometrika, 46(3/4):364, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333533.

Pearson:1959:NMQ

[Pea59c] E. S. Pearson. Note on Mr. Quenouille’s Edgeworth Type A transformation. Biometrika, 46(1/2):203–204, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332821.

Penfold:1959:BRB

[Pen59] A. P. Penfold. Book review: Principles of Statistical Techniques. A First Course from the Beginnings for Schools and Universities, by P. G. Moore. Biometrika, 46(1/2):276, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332854.

Pearson:1951:CPF

[PH51] E. S. Pearson and H. O. Hartley. Charts of the power function for analysis of variance tests, derived from the non-central F -distribution. Biometrika, 38(1/2):112–130, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332321.

Phillips:1959:EPS

[Phi59] A. W. Phillips. The estimation of parameters in systems of stochastic differential equations. Biometrika, 46(1/2):67–76, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332809.

Pillai:1952:DSR

[Pil52] K. C. S. Pillai. On the distribution of ‘Studentized’ range. Biometrika, 39(1/2):194–195, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332480.

Pillai:1956:DLS

[Pil56] K. C. S. Pillai. On the distribution of the largest or the smallest root of a matrix in multivariate analysis. Biometrika, 43(1/2):122– 127, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333585.

Pillai:1959:UPP

[Pil59] K. C. S. Pillai. Upper percentage points of the extreme Studentized deviate from the sample mean. Biometrika, 46(3/4):473–474, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333545.

Patterson:1959:IHM

[PL59] H. D. Patterson and S. Lipton. An investigation of Hartley’s method for fitting an exponential curve. Biometrika, 46(3/4):281–292, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333525.

Plackett:1950:STL

[Pla50] R. L. Plackett. Some theorems in least squares. Biometrika, 37(1/2): 149–157, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332158.

Plackett:1954:RFN

[Pla54] R. L. Plackett. A reduction formula for normal multivariate integrals. Biometrika, 41(3/4):351–360, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332716.

Plackett:1958:SHP

[Pla58] R. L. Plackett. Studies in the history of probability and statistics: VII. The principle of the arithmetic mean. Biometrika, 45(1/2):130– 135, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333051.

Pearson:1951:TPP

[PM51] E. S. Pearson and Maxine Merrington. Tables of the 5% and 0 · 5% points of Pearson curves (with argument β_{1} and β_{2}) expressed in standard measure. Biometrika, 38(1/2):4–10, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332312.

Powell:1955:SFG

[Pow55] E. O. Powell. Some features of the generation times of individual bacteria. Biometrika, 42(1/2):16–44, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333420.

Power:1958:BRB

[Pow58] E. A. Power. Book review: Vector Analysis, by L. Brand. Biometrika, 45(3/4):598, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333229.

Preston:1953:GMA

[Pre53] Eric J. Preston. A graphical method for the analysis of statistical distributions into two normal components. Biometrika, 40(3/4):460– 464, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333366.

Pillai:1959:HG

[PS59] K. C. Sreedharan Pillai and Pablo Samson, Jr. On Hotelling’s generalization of T^{2}. Biometrika, 46(1/2):160–168, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332818.

Pillai:1959:DES

[PT59] K. C. S. Pillai and Benjamin P. Tienzo. On the distribution of the extreme Studentized deviate from the sample mean. Biometrika, 46(3/4):467–472, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333544.

Quenouille:1949:ATC

[Que49] M. H. Quenouille. Approximate tests of correlation in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 11 (1):68–84, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983696. See extension in [Que56, Tuk58],

and reviews [Mil74, Efr79].

Quenouille:1953:MVD

[Que53] M. H. Quenouille. Modifications to the variate-difference method. Biometrika, 40(3/4):383–408, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333356.

Quenouille:1956:NBE

[Que56] M. H. Quenouille. Notes on bias in estimation. Biometrika, 43 (3–4):353–360, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332914. This extends earlier work [Que49, Tuk58] on what is now known as the Quenouille–Tukey jacknife technique. See reviews [Mil74, Efr79].

Quenouille:1957:ETV

[Que57] M. H. Quenouille. The effect of transformations of variables upon their correlation coefficients. Biometrika, 44(1/2):272–273, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333267.

Quenouille:1959:TRO

[Que59] M. H. Quenouille. Tables of random observations from standard distributions. Biometrika, 46(1/2):178–202, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332820.

Ramasubban:1958:MDM

[Ram58] T. A. Ramasubban. The mean difference and the mean deviation of some discontinuous distributions. Biometrika, 45(3/4):549–556, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333200.

Ramasubban:1959:GMD

[Ram59] T. A. Ramasubban. The generalized mean differences of the binomial and Poisson distributions. Biometrika, 46(1/2):223–229, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332825.

Rao:1953:SMD

[Rao53] K. S. Rao. A simple method of deriving best critical regions similar to the sample space in tests of an important class of composite hypotheses. Biometrika, 40(1/2):231–233, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333118.

Rao:1959:SPI

[Rao59] C. Radhakrishna Rao. Some problems involving linear hypotheses in multivariate analysis. Biometrika, 46(1/2):49–58, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332807.

Ray:1956:SAA

[Ray56] W. D. Ray. Sequential analysis applied to certain experimental designs in the analysis of variance. Biometrika, 43(3/4):388–403, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332917.

Russell:1958:OWV

[RB58] Thomas S. Russell and Ralph Allan Bradley. One-way variances in a two-way classification. Biometrika, 45(1/2):111–129, June 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333050.

Reiersol:1956:NSG

[Rei56] Olav Reiersøl. A note on the signs of Gross correlation coefficients and partial correlation coefficients. Biometrika, 43(3/4):480–482, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332932.

Roy:1957:FCM

[RG57] S. N. Roy and R. Gnanadesikan. Further contributions to multivariate confidence bounds. Biometrika, 44(3/4):399–410, December 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332872. See corrigenda [RG61].

Roy:1958:NFC

[RG58] S. N. Roy and R. Gnanadesikan. A note on ‘further contributions to multivariate confidence bounds’. Biometrika, 45(3/4):581, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333209.

Roy:1961:CFC

[RG61] S. N. Roy and R. Gnanadesikan. Corrigenda: “Further Contributions to Multivariate Confidence Bounds”. Biometrika, 48(3/4):474, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332775. See [RG57].

Rao:1950:GSL

[RK50] K. S. Rao and David G. Kendall. On the generalized second limittheorem in the calculus of probabilities. Biometrika, 37(3/4):224–230, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332375.

Rushton:1955:DMS

[RM55] S. Rushton and A. J. Mautner. The deterministic model of a simple epidemic for more than one community. Biometrika, 42(1/2):126–132, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333429.

Roy:1956:ISN

[RM56] S. N. Roy and S. K. Mitra. An introduction to some non-parametric generalizations of analysis of variance and multivariate analysis. Biometrika, 43(3/4):361–376, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332915.

Royston:1956:SHP

[Roy56] Erica Royston. Studies in the history of probability and statistics: III. A note on the history of the graphical presentation of data. Biometrika, 43(3/4):241–247, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332903.

Roy:1956:ICP

[RS56] S. N. Roy and A. E. Sarhan. On inverting a class of patterned matrices. Biometrika, 43(1/2):227–231, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333601.

Ruben:1954:MOS

[Rub54] H. Ruben. On the moments of order statistics in samples from normal populations. Biometrika, 41(1/2):200–227, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333017. See corrigenda [Ano54d].

Ruben:1956:MRP

[Rub56a] H. Ruben. On the moments of the range and product moments of extreme order statistics in normal samples. Biometrika, 43(3/4):458– 460, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332925.

Ruben:1956:SSN

[Rub56b] H. Ruben. On the sum of squares of normal scores. Biometrika, 43(3/4):456–458, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332924. See corrigenda [Rub65].

Ruben:1965:CSS

[Rub65] H. Ruben. Corrigenda: “On the sum of squares of normal scores”. Biometrika, 52(3/4):669, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333726. See [Rub56b].

Rudra:1952:DTS

[Rud52] A. Rudra. Discrimination in time-series analysis. Biometrika, 39 (3/4):434–439, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334040.

Rushton:1950:ST

[Rus50] S. Rushton. On a sequential t-test. Biometrika, 37(3/4):326–333, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332385.

Rushton:1952:TSS

[Rus52] S. Rushton. On a two-sided sequential t-test. Biometrika, 39(3/4):302– 308, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334026. See

corrigenda [Ano54c].

Sampford:1955:TNB

[Sam55] M. R. Sampford. The truncated negative binomial distribution. Biometrika, 42(1/2):58–69, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333422.

Sankaran:1958:NTC

[San58] Munuswamy Sankaran. On Nair’s transformation of the correlation coefficient. Biometrika, 45(3/4):567–571, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333205.

Sankaran:1959:NCC

[San59] Munuswamy Sankaran. On the non-central chi-square distribution. Biometrika, 46(1/2):235–237, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332828.

Saw:1958:CMS

[Saw58a] J. G. Saw. Corrigenda: “Moments of Sample Moments of Censored Samples from a Normal Population”. Biometrika, 45(3/4):587–588, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333213. See

[Saw58b].

Saw:1958:MSM

[Saw58b] J. G. Saw. Moments of sample moments of censored samples from a normal population. Biometrika, 45(1/2):211–221, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333057. See corrigenda [Saw58a, Saw79].

Saw:1959:ENP

[Saw59] J. G. Saw. Estimation of the normal population parameters given a singly censored sample. Biometrika, 46(1/2):150–159, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332817.

Saw:1979:ACM

[Saw79] J. G. Saw. Amendments and corrections: “Moments of Sample Moments of Censored Samples from a Normal Population”. Biometrika, 66 (1):194, April 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335267. See

[Saw58b].

Scheffe:1953:MJA

[Sch53] Henry Scheffé. A method for judging all contrasts in the analysis of variance. Biometrika, 40(1/2):87–104, June 1953. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333100. See correction [Sch69].

Scheffe:1969:CMJ

[Sch69] Henry Scheffe. Corrections: “A Method for Judging All Contrasts in the Analysis of Variance”. Biometrika, 56(1):229, March 1969. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334713. See [Sch53].

Shenton:1950:MLE

[She50] L. R. Shenton. Maximum likelihood and the efficiency of the method of moments. Biometrika, 37(1/2):111–116, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332153.

Shenton:1951:EMM

[She51] L. R. Shenton. Efficiency of the method of moments and the Gram-Charlier Type A distribution. Biometrika, 38(1/2):58–73, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332317.

Shenton:1954:INI

[She54] L. R. Shenton. Inequalities for the normal integral including a new continued fraction. Biometrika, 41(1/2):177–189, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333015.

Shenton:1958:MEM

[She58] L. R. Shenton. Moment estimators and maximum likelihood. Biometrika, 45(3/4):411–420, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2333188. See corrigenda [She59a, She61].

Shenton:1959:CME

[She59a] L. R. Shenton. Corrigenda: “Moment Estimators and Maximum Likelihood”. Biometrika, 46(3/4):502, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333569. See [She58].

Shenton:1959:DME

[She59b] L. R. Shenton. The distribution of moment estimators. Biometrika, 46(3/4):296–305, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333527.

Shenton:1961:CME

[She61] L. R. Shenton. Corrigenda: “Moment Estimators and Maximum Likelihood”. Biometrika, 48(3/4):474, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332776. See [She58].

Silverstone:1950:NCK

[Sil50] H. Silverstone. A note on the cumulants of Kendall’s S-distribution. Biometrika, 37(3/4):231–235, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332376.

Sillitto:1951:IBC

[Sil51] G. P. Sillitto. Interrelations between certain linear systematic statistics of samples from any continuous population. Biometrika, 38(3/4):377– 382, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332583.

Sillitto:1957:EPC

[Sil57] G. P. Sillitto. An extension property of a class of balanced incomplete block designs. Biometrika, 44(1/2):278–279, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333270.

Simon:1955:CSD

[Sim55] Herbert A. Simon. On a class of skew distribution functions. Biometrika, 42(3–4):425–440, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333389.

Sathe:1957:ADS

[SK57] Y. S. Sathe and A. R. Kamat. Approximations to the distributions of some measures of dispersion based on successive differences. Biometrika, 44(3/4):349–359, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332867.

Skellam:1951:RDT

[Ske51] J. G. Skellam. Random dispersal in theoretical populations. Biometrika, 38(1/2):196–218, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332328.

Skellam:1952:SSE

[Ske52] J. G. Skellam. Studies in statistical ecology: I. Spatial pattern. Biometrika, 39(3/4):346–362, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334030.

Skellam:1958:DAN

[Ske58] J. G. Skellam. On the derivation and applicability of Neyman’s type a distribution. Biometrika, 45(1/2):32–36, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333043.

Smith:1951:BRB

[Smi51] Cedric A. B. Smith. Book review: Sampling Methods for Censuses and Surveys by F. Yates. Biometrika, 38(1/2):262–263, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332336.

Smith:1955:BRBa

[Smi55a] Cedric A. B. Smith. Book review: Outline of Biometry and Statistics, by C. I. Bliss; D. W. Calhoun. Biometrika, 42(3/4):541, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333411.

Smith:1955:BRBb

[Smi55b] Cedric A. B. Smith. Book review: Statistics and Mathematics in Biology, by O. Kempthorne; T. A. Bancroft; J. W. Gowen; J. L. Lush. Biometrika, 42(3/4):542, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333412.

Smith:1956:BRB

[Smi56] Cedric A. B. Smith. Book review: Population Genetics, by Ching Chun Li. Biometrika, 43(3/4):492, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332940.

Smith:1959:BRB

[Smi59a] Cedric A. B. Smith. Book review: Contribuito allo Studio delle Tavole di Nuzialita, by G. Panizzon. Biometrika, 46(3/4):495, December 1959.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333557.

Smith:1959:CCR

[Smi59b] W. L. Smith. Corrigenda: “On the Cumulants of Renewal Processes”. Biometrika, 46(3/4):502, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333570. See [Smi59c].

Smith:1959:CRP

[Smi59c] Walter L. Smith. On the cumulants of renewal processes. Biometrika, 46(1/2):1–29, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332804. See corrigenda [Smi59b].

Somerville:1954:SPO

[Som54] Paul N. Somerville. Some problems of optimum sampling. Biometrika, 41(3/4):420–429, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332722.

Somers:1959:RAP

[Som59] Robert H. Somers. The rank analogue of product-moment partial correlation and regression, with application to manifold, ordered contingency tables. Biometrika, 46(1/2):241–246, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332830.

Srivastava:1958:ENN

[Sri58] A. B. L. Srivastava. Effect of non-normality on the power function of t-test. Biometrika, 45(3/4):421–430, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333189.

Srivastava:1959:ENN

[Sri59] A. B. L. Srivastava. Effect of non-normality on the power of the analysis of variance test. Biometrika, 46(1/2):114–122, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332813.

Stevens:1950:FLP

[Ste50] W. L. Stevens. Fiducial limits of the parameter of a discontinuous distribution. Biometrika, 37(1/2):117–129, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332154.

Stevens:1951:MVE

[Ste51a] W. L. Stevens. Mean and variance of an entry in a contingency table. Biometrika, 38(3/4):468–470, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332593.

Steyn:1951:WDD

[Ste51b] H. S. Steyn. The Wishart distribution derived by solving simultaneous linear differential equations. Biometrika, 38(3/4):470–472, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332594.

Stevens:1952:SSN

[Ste52] W. L. Stevens. Samples with the same number in each stratum. Biometrika, 39(3/4):414–417, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334033.

Stevens:1953:TAT

[Ste53] W. L. Stevens. Tables of the angular transformation. Biometrika, 40(1/2):70–73, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333098.

Sterne:1954:SRC

[Ste54] Theodore E. Sterne. Some remarks on confidence or fiducial limits. Biometrika, 41(1/2):275–278, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333026.

Stevens:1957:SIP

[Ste57] W. L. Stevens. Shorter intervals for the parameter of the binomial and Poisson distributions. Biometrika, 44(3/4):436–440, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332874.

Stuart:1950:CFN

[Stu50] A. Stuart. The cumulants of the first n natural numbers. Biometrika, 37(3/4):446, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332395.

Stuart:1951:ADR

[Stu51] A. Stuart. An application of the distribution of the ranking concordance coefficient. Biometrika, 38(1/2):33–42, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332315.

Stuart:1952:BRB

[Stu52] A. Stuart. Book review: Statistical Methodology Reviews, 1941–1950, by Oscar Krisen Buros. Biometrika, 39(1/2):215, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332489.

Stuart:1953:ECS

[Stu53] A. Stuart. The estimation and comparison of strengths of association in contingency tables. Biometrika, 40(1/2):105–110, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333101.

Stuart:1954:BRB

[Stu54] Alan Stuart. Book review: Sampling Methods for Censuses and Surveys by F. Yates. Biometrika, 41(1/2):284, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333035.

Stuart:1955:PSE

[Stu55a] Alan Stuart. A paradox in statistical estimation. Biometrika, 42 (3/4):527–529, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333402.

Stuart:1955:THM

[Stu55b] Alan Stuart. A test for homogeneity of the marginal distributions in a two-way classification. Biometrika, 42(3/4):412–416, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333387.

Stuart:1956:BRBb

[Stu56a] A. Stuart. Book review: Measuring Business Changes; a Handbook of Significant Business Indicators, by Richard M. Snyder. Biometrika, 43(3/4):493, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332943.

Stuart:1956:BRBa

[Stu56b] A. Stuart. Book review: The Art of Investment, by A. G. Ellinger. Biometrika, 43(3/4):493, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332942.

Stuart:1956:BVK

[Stu56c] Alan Stuart. Bounds for the variance of Kendall’s rank correlation statistic. Biometrika, 43(3/4):474–477, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332930.

Stuart:1957:SEB

[Stu57] Alan Stuart. A singularity in the estimation of binomial variance. Biometrika, 44(1/2):262–264, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333262.

Stuart:1958:BRB

[Stu58] A. Stuart. Book review: Non-Parametric Methods in Statistics, by D. A. S. Fraser. Biometrika, 45(3/4):591–592, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333218.

Sukhatme:1958:TSD

[Suk58] Balkrishna V. Sukhatme. A two sample distribution free test for comparing variances. Biometrika, 45(3/4):544–548, December 1958. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333199.

Sundrum:1953:MSS

[Sun53a] R. M. Sundrum. A method of systematic sampling based on order properties. Biometrika, 40(3/4):452–456, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333363.

Sundrum:1953:MRC

[Sun53b] R. M. Sundrum. Moments of the rank correlation coefficient τ in the general case. Biometrika, 40(3/4):409–420, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333357.

Sundrum:1954:RBE

[Sun54] R. M. Sundrum. On the relation between estimating efficiency and the power of tests. Biometrika, 41(3/4):542–544, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332733.

Tanner:1951:DPC

[Tan51] J. C. Tanner. The delay to pedestrians crossing a road. Biometrika, 38(3/4):383–392, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332584.

Tanner:1953:PIB

[Tan53] J. C. Tanner. A problem of interference between two queues. Biometrika, 40(1/2):58–69, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333097.

Tanner:1958:PCA

[Tan58] J. C. Tanner. A problem in the combination of accident frequencies. Biometrika, 45(3/4):331–342, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333182.

Tate:1955:TCB

[Tat55] Robert F. Tate. The theory of correlation between two continuous variables when one is dichotomized. Biometrika, 42(1/2):205–216, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333437.

Taylor:1950:CPT

[Tay50] J. Taylor. The comparison of pairs of treatments in split-plot experiments. Biometrika, 37(3/4):443–444, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332393.

Taylor:1956:ELS

[Tay56] J. Taylor. Exact linear sequential tests for the mean of a normal distribution. Biometrika, 43(3/4):452–455, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332923.

Thatcher:1957:SHP

[Tha57] A. R. Thatcher. Studies in the history of probability and statistics. VI. A note on the early solutions of the problem of the duration of play. Biometrika, 44(3/4):515–518, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332881.

Thomas:1951:STR

[Tho51a] Marjorie Thomas. Some tests for randomness in plant populations. Biometrika, 38(1/2):102–111, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332320.

Thompson:1951:TLD

[Tho51b] H. R. Thompson. Truncated lognormal distributions: I. Solution by moments. Biometrika, 38(3/4):414–422, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332587.

Thomson:1953:SFD

[Tho53] George Wm. Thomson. Sacle factors and degrees of freedom for small sample sizes for χ-approximation to the range. Biometrika, 40(3/4):449–450, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333361.

Thompson:1954:NCD

[Tho54] H. R. Thompson. A note on contagious distributions. Biometrika, 41(1/2):268–271, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333024.

Thompson:1955:SPP

[Tho55a] H. R. Thompson. Spatial point processes, with applications to ecology. Biometrika, 42(1/2):102–115, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333426.

Thomson:1955:BRR

[Tho55b] George W. Thomson. Bounds for the ratio of range to standard deviation. Biometrika, 42(1/2):268–269, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333446.

Tocher:1950:ENP

[Toc50] K. D. Tocher. Extension of the Neyman-Pearson theory of tests to discontinuous variates. Biometrika, 37(1/2):130–144, June 1950. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332156.

Tocher:1952:NDP

[Toc52a] K. D. Tocher. A note on the design problem. Biometrika, 39(1/2): 189, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332477.

Tocher:1952:CSR

[Toc52b] K. D. Tocher. On the concurrence of a set of regression lines. Biometrika, 39(1/2):109–117, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332468.

Tukey:1955:IAR

[Tuk55] John W. Tukey. Interpolations and approximations related to the normal range. Biometrika, 42(3/4):480–485, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333393.

Tukey:1957:AUP

[Tuk57] John W. Tukey. Approximations to the upper 5% points of Fisher’s B distribution and non-central χ^{2}. Biometrika, 44(3/4):528–530, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332885.

Tukey:1958:BCQ

[Tuk58] John W. Tukey. Bias and confidence in not-quite large samples (abstract). Annals of Mathematical Statistics, 29(2):614, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706647. This

is the other half of the Quenouille–Tukey jackknife technique; see [Que49, Que56, Mil74, Efr79].

Trickett:1954:CTM

[TW54] W. H. Trickett and B. L. Welch. On the comparison of two means: Further discussion of iterative methods for calculating tables. Biometrika, 41(3/4):361–374, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332717.

Trickett:1956:FCV

[TWJ56] W. H. Trickett, B. L. Welch, and G. S. James. Further critical values for the two-means problem. Biometrika, 43(1/2):203–205, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333593.

Vajda:1959:BRBb

[Vaj59a] S. Vajda. Book review: Linear Programming: Methods and Applications, by Saul I. Gass. Biometrika, 46(3/4):494, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333554.

Vajda:1959:BRBa

[Vaj59b] S. Vajda. Book review: Scientific Programming in Business and Industry, by Andrew Vazsonyi. Biometrika, 46(1/2):274, June 1959. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332852.

vanderHeiden:1952:CTM

[vdH52] J. A. van der Heiden. On a correction term in the method of paired comparisons. Biometrika, 39(1/2):211–212, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332485.

VanElteren:1959:AET

[VN59] Ph. Van Elteren and G. E. Noether. The asymptotic efficiency of the χ_{r}^{2}-test for a balanced incomplete block design. Biometrika, 46 (3/4):475–477, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333546.

Walker:1950:ACB

[Wal50] Sir Gilbert Walker. Apparent correlation between independent series of autocorrelated observations. Biometrika, 37(1/2):184–185, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332164.

Walker:1956:GFT

[Wal56] A. M. Walker. A goodness of fit test for spectral distribution functions of stationary time series with normal residuals. Biometrika, 43(3/4):257–275, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332905.

Watson:1955:SCR

[Wat55] G. S. Watson. Serial correlation in regression analysis. I. Biometrika, 42(3/4):327–341, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333382.

Watson:1956:NCM

[Wat56a] G. S. Watson. A note on the circular multivariate distribution. Biometrika, 43(3/4):467, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332928.

Watson:1956:JDC

[Wat56b] G. S. Watson. On the joint distribution of the circular serial correlation coefficients. Biometrika, 43(1/2):161–168, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333589.

Watson:1957:BRB

[Wat57a] G. L. Watson. Book review: Irrationalzahlen by O. Perron. Biometrika, 44(1/2):299, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333284.

Watson:1957:GFT

[Wat57b] G. S. Watson. The χ^{2} goodness-of-fit test for normal distributions. Biometrika, 44(3/4):336–348, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332866.

Waugh:1955:ADB

[Wau55] W. A. O’N. Waugh. An age-dependent birth and death process. Biometrika, 42(3/4):291–306, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333379.

Waugh:1958:CMP

[Wau58] W. A. O’N. Waugh. Conditioned Markov processes. Biometrika, 45(1/2):241–249, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333060. See corrigenda [Wau59].

Waugh:1959:CCM

[Wau59] W. A. O’N. Waugh. Corrigenda: “Conditioned Markoff Processes”. Biometrika, 46(1/2):279, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332858. See [Wau58].

Welch:1951:CSM

[Wel51] B. L. Welch. On the comparison of several mean values: An alternative approach. Biometrika, 38(3/4):330–336, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332579.

Welch:1954:BRB

[Wel54] B. L. Welch. Book review: Cambridge Elementary Statistical Tables by D. V. Lindley; J. C. P. Miller. Biometrika, 41(1/2):285, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333037.

Welch:1958:BRB

[Wel58] B. L. Welch. Book review: Tables of the Non-Central t-Distribution, by G. J. Resnikoff; G. J. Lieberman. Biometrika, 45(3/4):596, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333225.

Watson:1956:SCR

[WH56] G. S. Watson and E. J. Hannan. Serial correlation in regression analysis. II. Biometrika, 43(3/4):436–448, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332921.

Whittle:1951:BRB

[Whi51] P. Whittle. Book review: Symposium on Stochastic Processes: Stochastic Processes and Statistical Physics by J. E. Moyal; Symposium on Stochastic Processes: Some Evolutionary Stochastic Processes by M. S. Bartlett; Symposium on Stochastic Processes: Stochastic Processes and Population Growth by D. G. Kendall. Biometrika, 38(1/2):263– 265, June 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332337.

Whittle:1952:TFT

[Whi52] P. Whittle. Tests of fit in time series. Biometrika, 39(3/4):309–318, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334027.

Whittle:1954:SPP

[Whi54] P. Whittle. On stationary processes in the plane. Biometrika, 41 (3/4):434–449, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332724. See corrigenda [Ano55f].

Whittle:1955:OSE

[Whi55] P. Whittle. The outcome of a stochastic epidemic — a note on Bailey’s paper. Biometrika, 42(1/2):116–122, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333427.

Whittle:1956:VYV

[Whi56] P. Whittle. On the variation of yield variance with plot size. Biometrika, 43(3/4):337–343, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332912.

Williams:1952:IIF

[Wil52a] E. J. Williams. The interpretation of interactions in factorial experiments. Biometrika, 39(1/2):65–81, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332464.

Williams:1952:SET

[Wil52b] E. J. Williams. Some exact tests in multivariate analysis. Biometrika, 39(1/2):17–31, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332460.

Williams:1952:USA

[Wil52c] E. J. Williams. Use of scores for the analysis of association in contingency tables. Biometrika, 39(3/4):274–289, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2334024.

Williams:1952:EDS

[Wil52d] R. M. Williams. Experimental designs for serially correlated observations. Biometrika, 39(1/2):151–167, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332474.

Williams:1953:TSC

[Wil53] E. J. Williams. Tests of significance for concurrent regression lines. Biometrika, 40(3/4):297–305, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333348.

Wilkie:1954:BRB

[Wil54] D. R. Wilkie. Book review: A Statistical Primer by F. N. David. Biometrika, 41(1/2):281–282, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333030.

Wilk:1955:RAG

[Wil55a] M. B. Wilk. The randomization analysis of a generalized randomized block design. Biometrika, 42(1/2):70–79, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333423. See corrigenda [Ano56d].

Williams:1955:STD

[Wil55b] E. J. Williams. Significance tests for discriminant functions and linear functional relationships. Biometrika, 42(3/4):360–381, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333384.

Williams:1956:SHP

[Wil56a] C. B. Williams. Studies in the history of probability and statistics: IV. A note on an early statistical study of literary style. Biometrika, 43(3/4):248–256, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332904.

Williams:1956:VMS

[Wil56b] R. M. Williams. The variance of the mean of systematic samples. Biometrika, 43(1/2):137–148, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333587.

Wilkinson:1957:APC

[Wil57] John W. Wilkinson. An analysis of paired comparison designs with incomplete repetitions. Biometrika, 44(1/2):97–113, June 1957. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333243.

Wilkie:1958:BRB

[Wil58a] D. R. Wilkie. Book review: Individual Differences in Night-Vision Efficiency, by M. H. Pirenne; F. H. C. Marriott; E. F. O’Doherty. Biometrika, 45(1/2):288–289, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333082.

Williams:1958:SRE

[Wil58b] E. J. Williams. Simultaneous regression equations in experimentation. Biometrika, 45(1/2):96–110, June 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333049.

Wintner:1957:SDR

[Win57] Aurel Wintner. Student’s distribution and Riemann’s elliptic geometry. Biometrika, 44(1/2):264–265, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333263.

Wise:1950:IBF

[Wis50] M. E. Wise. The incomplete beta function as a contour integral and a quickly converging series for its inverse. Biometrika, 37(3/4):208–218, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332373.

Wishart:1952:BRB

[Wis52a] John Wishart. Book review: Introduction to Statistical Analysis, by Wilfred J. Dixon; Frank J. Massey. Biometrika, 39(3/4):441–442, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334044.

Wishart:1952:MCS

[Wis52b] John Wishart. Moment coefficients of the k-statistics in samples from a finite population. Biometrika, 39(1/2):1–13, April 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2332458.

Wise:1954:QCE

[Wis54a] M. E. Wise. A quickly convergent expansion for cumulative hypergeometric probabilities, direct and inverse. Biometrika, 41(3/4):317–329, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332712. See cor-

rigenda [Ano55e].

Wishart:1954:BRB

[Wis54b] J. Wishart. Book review: Statistical Tables, by Research Association of Statistical Sciences. Biometrika, 41(3/4):566, December 1954. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332743.

Wishart:1954:FMD

[Wis54c] John Wishart. The factorial moments of the distribution of joins between line segments. Biometrika, 41(3/4):555–556, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332737.

Wise:1955:AFS

[Wis55] J. Wise. The autocorrelation function and the spectral density function. Biometrika, 42(1/2):151–159, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333432.

Wise:1956:SCS

[Wis56a] J. Wise. Stationarity conditions for stochastic processes of the autoregressive and moving-average type. Biometrika, 43(1/2):215–219, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333598.

Wishart:1956:PLN

[Wis56b] John Wishart. χ^{2} probabilities for large numbers of degrees of freedom. Biometrika, 43(1/2):92–95, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333581.

Wishart:1959:BRBb

[Wis59a] D. M. G. Wishart. Book review: Finite Queueing Tables, by L. G. Peck; R. N. Hazelwood. Biometrika, 46(3/4):496, December 1959.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333560.

Wishart:1959:BRBa

[Wis59b] D. M. G. Wishart. Book review: Queues, Inventories and Maintenance, by Philip M. Morse. Biometrika, 46(3/4):496, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333559.

Wishart:1953:OPF

[WM53] John Wishart and Theocharis Metakides. Orthogonal polynomial fitting. Biometrika, 40(3/4):361–369, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2333354.

Wooding:1956:MDC

[Woo56] R. A. Wooding. The multivariate distribution of complex normal variables. Biometrika, 43(1/2):212–215, June 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333597.

Watson:1956:CST

[WW56] G. S. Watson and E. J. Williams. On the construction of significance tests on the circle and the sphere. Biometrika, 43(3/4):344–352, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332913.

Yates:1955:NAC

[Yat55a] F. Yates. A note on the application of the combination of probabilities test to a set of 2 × 2 tables. Biometrika, 42(3/4):404–411, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333386.

Yates:1955:UTM

[Yat55b] F. Yates. The use of transformations and maximum likelihood in the analysis of quantal experiments involving two treatments. Biometrika, 42(3/4):382–403, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333385.

Zinger:1958:CBA

[ZSP58] A. Zinger and J. St-Pierre. On the choice of the best amongst three normal populations with known variances. Biometrika, 45(3/4):436– REFERENCES 146

446, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333191.